### v1

# Discrete equations from Bäcklund transformations of the fifth Painlevé equation

Abstract
In this paper discrete equations are derived from Bäcklund transformations of the fifth Painlevé equation, including a new discrete equation which has ternary symmetry. There are two classes of rational solutions of the fifth Painlevé equation, one expressed in terms of the generalised Laguerre polynomials and the other in terms of the generalised Umemura polynomials, both of which can be expressed as Wronskians of Laguerre polynomials. Hierarchies of rational solutions of the discrete equations are derived in terms of the generalised Laguerre and generalised Umemura polynomials. It is known that there is nonuniqueness of some rational solutions of the fifth Painlevé equation. Pairs of nonunique rational solutions are used to derive distinct hierarchies of rational solutions which satisfy the same discrete equation.
Mathematics Subject Classification (2020): 33E17 34M55 37J70 39A10
1 Introduction
In this paper we are concerned with discrete Painlevé equations associated with the fifth Painlevé equation
$\frac{\mathrm{d}^{2}{w}}{{\mathrm{d}{z}}^{2}}=\left(\frac{1}{2w}+\frac{1}{w-1}\right)\left(\frac{\mathrm{d}{w}}{{\mathrm{d}{z}}}\right)^{\!\!2}-\frac{1}{z}\frac{\mathrm{d}{w}}{{\mathrm{d}{z}}}+\frac{(w-1)^{2}(\alpha w^{2}+{\beta})}{z^{2}w}+\frac{\gamma w}{z}+\frac{\delta w(w+1)}{w-1},$ (1.1)
with $\alpha$, $\beta$, $\gamma$ and $\delta$ constants. We consider the generic case of (1.1) when $\delta\not=0$, so we set $\delta=-\tfrac{1}{2}$, without loss of generality (by rescaling $z$ if necessary) and obtain
$\frac{\mathrm{d}^{2}{w}}{{\mathrm{d}{z}}^{2}}=\left(\frac{1}{2w}+\frac{1}{w-1}\right)\left(\frac{\mathrm{d}{w}}{{\mathrm{d}{z}}}\right)^{\!\!2}-\frac{1}{z}\frac{\mathrm{d}{w}}{{\mathrm{d}{z}}}+\frac{(w-1)^{2}(\alpha w^{2}+{\beta})}{z^{2}w}+\frac{\gamma w}{z}-\frac{w(w+1)}{2(w-1)},$ (1.2)
which we will refer to as PV.
The six Painlevé equations were discovered by Painlevé, Gambier and their colleagues in the late 19th and early 20th centuries whilst studying second order ordinary differential equations of the form
$\frac{\mathrm{d}^{2}{w}}{{\mathrm{d}{z}}^{2}}=F\left(z,w,\frac{\mathrm{d}{w}}{{\mathrm{d}{z}}}\right),$ (1.3)
where $F$ is rational in ${\mathrm{d}}w/{\mathrm{d}}z$ and $w$ and analytic in $z$. The Painlevé transcendents, i.e. the solutions of the Painlevé equations, can be thought of as nonlinear analogues of the classical special functions. The Painlevé equations are regarded as being “completely integrable” and solvable using Riemann-Hilbert techniques [21]. The general solutions of the Painlevé equations are transcendental in the sense that they cannot be expressed in terms of known elementary functions and so require the introduction of a new transcendental function to describe their solution. However, it is well known that all the Painlevé equations, except PI, possess rational solutions, algebraic solutions and solutions expressed in terms of the classical special functions for special values of the parameters, see, e.g. [11, 33] and the references therein. These hierarchies are usually generated from “seed solutions” using the associated Bäcklund transformations and frequently can be expressed in the form of determinants.
Discrete Painlevé equations also have a long history. The discrete Painlevé I (dPI) equation is
$x_{n+1}+x_{n}+x_{n-1}=1+\frac{\lambda n+\rho}{x_{n}},$ (1.4)
with $\lambda$ and $\rho$ constants. Equation (1.4) was studied in 1939 by Shohat [66], developing ideas due to Laguerre [41] dating from 1885, in the study of orthogonal polynomials associated with the quartic weight $\exp(-\tfrac{1}{4}x^{4})$ on the real line. Subsequently this problem was studied by Freud [24], see also [8], though it was not recognised at the time as a discrete Painlevé equation. Equation (1.4) was then rediscovered in the 1980s in models of gauge field theory by Bessis, Itzykson and Zuber [7] and in quantum gravity by Brézin and Kazakov [9], who showed that its continuum limit is the first Painlevé equation (PI)
$\frac{\mathrm{d}^{2}{w}}{{\mathrm{d}{z}}^{2}}=6w^{2}+z,$ (1.5)
see also [20].
The discrete Painlevé II (dPII) equation is
$x_{n+1}+x_{n-1}=\frac{\zeta[(n+\lambda)x_{n}+\rho]}{1-x_{n}^{2}},$ (1.6)
with $\zeta$, $\lambda$ and $\rho$ constants, which was derived by Periwal and Shevitz [60] in a study of unitary matrix models in a double scaling limit using orthogonal polynomials on the unit circle. Around the same time, Nijhoff and Papageorgiou [47] independently derived dPII (1.6) and showed that its continuum limit is the second Painlevé equation (PII)
$\frac{\mathrm{d}^{2}{w}}{{\mathrm{d}{z}}^{2}}=2w^{3}+zw+\alpha,$ (1.7)
with $\alpha$ a constant; Periwal and Shevitz [60] gave a continuum limit to PII (1.7) with $\alpha=0$.
One method for deriving discrete Painlevé equations is using Bäcklund transformations of (continuous) Painlevé equations. For example, it is well known that PII (1.7) has the Bäcklund transformations
$w(z;\alpha+1)$ $=-w(z;\alpha)-\frac{2\alpha+1}{2w^{2}(z;\alpha)+2w^{\prime}(z;\alpha)+z},$ (1.8a)
$w(z;\alpha-1)$ $=-w(z;\alpha)-\frac{2\alpha-1}{2w^{2}(z;\alpha)-2w^{\prime}(z;\alpha)+z}.$ (1.8b)
with $^{\prime}\equiv\mathrm{d}/\mathrm{d}{z}$ [18, 25]. Hence letting $x_{n}=w(z;n+\lambda-\tfrac{1}{2})$, with $\alpha=n+\lambda-\tfrac{1}{2}$, then by eliminating $w^{\prime}(z;\alpha)$ we obtain another discrete equation, commonly known as alt-dPI
$\frac{n+\lambda}{x_{n+1}+x_{n}}+\frac{n+\lambda-1}{x_{n}+x_{n-1}}+2x_{n}^{2}+z=0,$ (1.9)
which has PI (1.5) as its continuum limit [19]. The alt-dPI equation (1.9) appeared in 1981 in the work of Jimbo and Miwa [35, equation (5.3)], though it was not recognised to be a discrete Painlevé equation until the work of Fokas, Grammaticos and Ramani [19].
In some respects the discrete Painlevé equations are much richer than the Painlevé equations. For example, they can have solutions which have no continuum limit and other properties which are lost in the continuous limit. Perhaps the most fundamental difference is that, whilst the Painlevé equations are unique up to a Möbius transformation, this is not true for the discrete Painlevé equations. For each discrete Painlevé equation there exists a number of different forms, as illustrated by dPI (1.4) and alt-dPI (1.9) which both have PI (1.5) as a continuum limit yet are not equivalent as there is no transformation relating the solutions (1.4) and (1.9). Further their exact solutions are quite different since dPI (1.4) arises from Bäcklund transformations of the fourth Painlevé equation (PIV)
$\frac{\mathrm{d}^{2}{w}}{{\mathrm{d}{z}}^{2}}=\frac{1}{2w}\left(\frac{\mathrm{d}{w}}{{\mathrm{d}{z}}}\right)^{\!2}+\tfrac{3}{2}w^{3}+4zw^{2}+2(z^{2}-\alpha)w+\frac{\beta}{w},$ (1.10)
with $\alpha$ and $\beta$ constants [27], whereas alt-dPI (1.9) arises from Bäcklund transformations of PII (1.7), as shown above.
The nomenclature for discrete Painlevé equations can be misleading. The original naming of discrete Painlevé equations came from their continuum limits but this can be ambiguous. For example, there are generalisations of dPI (1.4) and dPII (1.6), see equations (3.31) and (3.7), respectively, which have continuum limits to other Painlevé equations, as discussed below. Another approach is to classify the discrete Painlevé equations by the type of affine Weyl group and associated affine Dynkin diagram they are derived from, which is now commonly used. Sakai [64], extending the earlier work of Okamoto [55] for the (continuous) Painlevé equations, approached the discrete Painlevé equations from the point of view of algebraic geometry. Sakai clarified the algebraic nature of discrete Painlevé equations and gave a classification scheme. A comprehensive review of the geometric aspects of the continuous and discrete Painlevé equations is given by Kajiwara, Noumi and Yamada [38]. There is a recent discussion of the symmetry group of dPII (1.6) in [16]. Discrete Painlevé equations arise in various applications such as random matrix theory and orthogonal polynomials, cf. [71, 72].
There are three distinct types of discrete Painlevé equations: (i), those for which the discrete variable is additive, which is the case for the discrete equations we discuss; (ii), the $q$-discrete Painlevé equations for which the discrete variable is multiplicative; and (iii), elliptic discrete Painlevé equations whose coefficients are elliptic functions of the independent variable $n$.
From the works of Okamoto [56, 57, 58, 59], it is known that the parameter spaces of PIIPVI all admit the action of an extended affine Weyl group; the group acts as a group of Bäcklund transformations. In a series of papers, Noumi and Yamada [48, 49, 51, 53] have implemented this idea to derive a hierarchy of dynamical systems associated to the affine Weyl group of type $\widetilde{A}_{N}^{(1)}$, often known as “symmetric forms of the Painlevé equations”. The behaviour of each dynamical system varies depending on whether $N$ is even or odd. The first member of the $\widetilde{A}_{2n+1}^{(1)}$ hierarchy, i.e. $\widetilde{A}_{3}^{(1)}$, usually known as symmetric-PV (sPV), is equivalent to PV (1.2) and given by
$z\frac{\mathrm{d}{f_{0}}}{{\mathrm{d}{z}}}$ $=f_{0}f_{2}(f_{1}-f_{3})+(\tfrac{1}{2}-\vartheta_{2})f_{0}+\vartheta_{0}f_{2},$ (1.11a)
$z\frac{\mathrm{d}{f_{1}}}{{\mathrm{d}{z}}}$ $=f_{1}f_{3}(f_{2}-f_{0})+(\tfrac{1}{2}-\vartheta_{3})f_{1}+\vartheta_{1}f_{3},$ (1.11b)
$z\frac{\mathrm{d}{f_{2}}}{{\mathrm{d}{z}}}$ $=f_{2}f_{0}(f_{3}-f_{1})+(\tfrac{1}{2}-\vartheta_{0})f_{2}+\vartheta_{2}f_{0},$ (1.11c)
$z\frac{\mathrm{d}{f_{3}}}{{\mathrm{d}{z}}}$ $=f_{3}f_{1}(f_{0}-f_{2})+(\tfrac{1}{2}-\vartheta_{1})f_{3}+\vartheta_{3}f_{1},$ (1.11d)
with the normalisations
$f_{0}(z)+f_{2}(z)=\sqrt{z},\qquad f_{1}(z)+f_{3}(z)=\sqrt{z},$ (1.11e)
and $\vartheta_{0}$, $\vartheta_{1}$, $\vartheta_{2}$ and $\vartheta_{3}$ are constants such that
$\vartheta_{0}+\vartheta_{1}+\vartheta_{2}+\vartheta_{3}=1.$ (1.11f)
The symmetric system sPV (1.11) was found by Adler [1] in the context of periodic chains of Bäcklund transformations and has applications in random matrix theory, cf. [23].
The outline of this paper is as follows. In §2 we describe Bäcklund transformations of PV (1.2). In §3 we derive some discrete Painlevé equations from these Bäcklund transformations, including a new discrete equation which has a ternary symmetry. There have been earlier studies deriving discrete Painlevé equations from Bäcklund transformations of PV (1.2), see [19, 69]. Our study explicitly gives the Bäcklund transformations which generate the discrete Painlevé equations. These then can be used to obtain exact solutions of the discrete equations in terms of known solutions of PV (1.2). In §4 we describe the rational solutions of PV (1.2) which are expressed in terms of Wronskians of Laguerre polynomials. In §5 we derive hierarchies of rational solutions of the discrete Painlevé equations derived in §3 in terms of the rational solutions of PV (1.2) given in §4. Further pairs of rational solutions which although different, which satisfy PV (1.2) for the same parameters, are used to derive distinct hierarchies of rational solutions that satisfy the same discrete equation. Finally in §6 we discuss our results.
2 Bäcklund transformations of the fifth Painlevé equation
In this section we describe the Bäcklund transformations of PV (1.2).
Definition 2.1.
A Bäcklund transformation maps a solution of a given Painlevé equation to another solution of the same Painlevé equation, with different values of the parameters, or another equation.
Definition 2.2.
Let $w=w(z;\alpha,\beta,\gamma)$ be a solution of PV (1.2) such that
$\Phi(w;z)=z\frac{\mathrm{d}{w}}{{\mathrm{d}{z}}}-\varepsilon_{2}aw^{2}+(\varepsilon_{2}a-\varepsilon_{3}b+\varepsilon_{1}z)w+\varepsilon_{3}b\not=0,$ (2.1)
and
$\Phi(w;z)-2\varepsilon_{1}zw\not=0,$ (2.2)
where $a=\sqrt{2\alpha}$ and $b=\sqrt{-2\beta}$ and $\gamma=c$. Then the Bäcklund transformation $\mathcal{T}_{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}}$ is defined by
$\mathcal{T}_{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}}:\quad W(z;\alpha,\beta,\gamma)=1-\frac{2\varepsilon_{1}zw}{\Phi(w;z)}=1-\frac{2\varepsilon_{1}zw}{zw^{\prime}-\varepsilon_{2}aw^{2}+(\varepsilon_{2}a-\varepsilon_{3}b+\varepsilon_{1}z)w+\varepsilon_{3}b},$ (2.3)
with $^{\prime}\equiv\mathrm{d}/\mathrm{d}{z}$, $\varepsilon_{j}=\pm 1$, $j=1,2,3$, independently, and for the parameters
$\mathcal{T}_{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}}(a,b,c)=\Big(\tfrac{1}{2}\big[c+\varepsilon_{1}(1-\varepsilon_{3}b-\varepsilon_{2}a)\big],\tfrac{1}{2}\big(c-\varepsilon_{1}(1-\varepsilon_{3}b-\varepsilon_{2}a)\big),\varepsilon_{1}(\varepsilon_{3}b-\varepsilon_{2}a)\Big),$ (2.4)
cf. [29, 30, 31], [33, §39] or [54, §32.7(v)].
Note that we are using the same notation as in [54, §32.7(v)] which slightly differs from that used in [29, 30, 31] or [33, §39].
Remarks 2.3.
  1. 1.
    Since $\varepsilon_{j}=\pm 1$, $j=1,2,3$, independently, there are eight distinct transformations $\mathcal{T}_{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}}$. The parameter $\varepsilon_{1}$ reflects that the transformation $(z,\gamma)\to(-z,-\gamma)$ is a discrete symmetry of PV. The parameters $\varepsilon_{2}$ and $\varepsilon_{3}$ reflect that $\alpha=\tfrac{1}{2}a^{2}$ and $\beta=-\tfrac{1}{2}b^{2}$, respectively, so $a$ and $b$ can take either sign.
  2. 2.
    The Bäcklund transformation $\mathcal{T}_{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}}$ is usually given for the parameters $\alpha$, $\beta$ and $\gamma$, in terms of $a$, $b$ and $c$. However in order to apply a sequence of Bäcklund transformations, it’s preferable to define the action of Bäcklund transformation on the parameters $a$, $b$ and $c$, rather than $\alpha$, $\beta$ and $\gamma$, to avoid any ambiguity in taking a square root.
  3. 3.
    We note that the Hamiltonian for PV (1.2) in [35] is written in terms of the monodromy parameters $\theta_{0}$, $\theta_{1}$ and $\theta_{\infty}$, which are related to the parameters $\alpha$, $\beta$ and $\gamma$ through
    $\alpha=\tfrac{1}{8}(\theta_{0}-\theta_{1}+\theta_{\infty})^{2},\qquad\beta=-\tfrac{1}{8}(\theta_{0}-\theta_{1}-\theta_{\infty})^{2},\qquad\gamma=1-\theta_{0}-\theta_{1},$
    see also [57], as are the Schlesinger transformations, which relate two solutions of the associated monodromy problem, for PV (1.2) [22, 46]. Further the discrete symmetries of PV (1.2), corresponding to the extended affine Weyl group of type $A_{3}^{(1)}$, act naturally on the root variables, see, for example, [13].
  4. 4.
    When applying the Bäcklund transformation $\mathcal{T}_{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}}$ followed by the Bäcklund transformation $\mathcal{T}_{\sigma_{1},\sigma_{2},\sigma_{3}}$ we use the notation $\mathcal{T}_{\sigma_{1},\sigma_{2},\sigma_{3}}\circ\mathcal{T}_{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}}$, i.e.
    $\mathcal{T}_{\sigma_{1},\sigma_{2},\sigma_{3}}\big(\mathcal{T}_{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}}(w;a,b,c)\big)=\mathcal{T}_{\sigma_{1},\sigma_{2},\sigma_{3}}\circ\mathcal{T}_{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}}(w;a,b,c).$
3 Deriving discrete Painlevé equations
In this section we derive four discrete Painlevé equations which are associated with PV (1.2) using the Bäcklund transformations described in §2.
To derive a discrete Painlevé equation from Bäcklund transformations of PV (1.2), we use a Bäcklund transformation $\mathcal{R}_{+}$, which relates a solution $w(z;a,b,c)$ to another solution $w_{+}(z;a_{+},b_{+},c_{+})$ and its inverse transformation $\mathcal{R}_{-}=\mathcal{R}_{+}^{-1}$, which relates $w(z;a,b,c)$ to a third solution $w_{-}(z;a_{-},b_{-},c_{-})$. Eliminating the derivative between the two Bäcklund transformations gives an algebraic relation between $w(z;a,b,c)$, $w_{+}(z;a_{+},b_{+},c_{+})$ and $w_{-}(z;a_{-},b_{-},c_{-})$, which is the discrete equation. Then setting $w(z;a,b,c)=w_{n}$, ${w_{\pm}}(z;a_{\pm},b_{\pm},c_{\pm})=w_{n\pm 1}$, with $(a,b,c)=(a_{n},b_{n},c_{n})$ and $(a_{\pm},b_{\pm},c_{\pm})=(a_{n\pm 1},b_{n\pm 1},c_{n\pm 1})$, gives the chain
$\{w_{n-1},{a}_{n-1},{b}_{n-1},{c}_{n-1}\}\enskip\begin{array}[]{c}{\scriptstyle{\displaystyle{\mathcal{R}_{+}}}}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\\[-8.0pt] \longleftarrow\!\!\!-\!\!\!-\!\!\!-\\[-7.0pt] {\scriptstyle{\displaystyle\mathcal{R}_{-}}}\end{array}\enskip\{w_{n},a_{n},b_{n},c_{n}\}\enskip\begin{array}[]{c}{\scriptstyle{\displaystyle{\mathcal{R}_{+}}}}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\\[-8.0pt] \longleftarrow\!\!\!-\!\!\!-\!\!\!-\\[-7.0pt] {\scriptstyle{\displaystyle{\mathcal{R}_{-}}}}\end{array}\enskip\{w_{n+1},{a}_{n+1},{b}_{n+1},{c}_{n+1}\}.$
The parameters satisfy a linear discrete system derived from (2.4). See, for example, [14, 19] for more details of this procedure.
Remarks 3.1.
  1. 1.
    It is essential that $\mathcal{R}_{-}$ is the inverse transformation of $\mathcal{R}_{+}$. It is possible to create an algebraic relation between three solutions of PV (1.2) using two Bäcklund transformations $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ when $\mathcal{R}_{2}\not=\mathcal{R}_{1}^{-1}$, but this won’t necessarily result in a discrete Painlevé equation. For example, in [33, §39], see also [30, 31], there are algebraic relations of three (and four) solutions of PV, but these are not discrete Painlevé equations.
  2. 2.
    The Bessel function $J_{n}(z)$ satisfies the differential equation
    $z^{2}\frac{\mathrm{d}^{2}{J_{n}}}{{\mathrm{d}{z}}^{2}}+z\frac{\mathrm{d}{J_{n}}}{{\mathrm{d}{z}}}+(z^{2}-n^{2})J_{n}=0,$ (3.1)
    and the recurrence relations
    $z\frac{\mathrm{d}{J_{n}}}{{\mathrm{d}{z}}}(z)=zJ_{n-1}(z)-nJ_{n}(z),$ (3.2a)
    $z\frac{\mathrm{d}{J_{n}}}{{\mathrm{d}{z}}}(z)=nJ_{n}(z)-zJ_{n+1}(z),$ (3.2b)
    $zJ_{n+1}(z)-2nJ_{n}(z)+zJ_{n-1}(z)=0,$ (3.2c)
    cf. [54, §10]. Eliminating $\frac{\mathrm{d}{J_{n}}}{{\mathrm{d}{z}}}(z)$ between (3.2a) and (3.2b) yields the recurrence relation (3.2c). Letting $n\to n+1$ in (3.2a) and eliminating $J_{n+1}(z)$ using (3.2b) gives the differential equation (3.1). Thus the Bessel function $J_{n}(z)$ satisfies both the differential equation (3.1) and the difference equation (3.2c). An analogous situation occurs for the Painlevé equations, except PI (1.5).
In the next three subsections, we derive discrete equations using the Bäcklund transformations
[…]
as well as a discrete equation which has a ternary symmetry using the Bäcklund transformation
$\mathcal{R}_{4}=\mathcal{R}_{3}^{2}=\mathcal{T}_{-1,1,-1}\circ\mathcal{T}_{1,-1,-1}^{2}\circ\mathcal{T}_{1,-1,1}.$
3.1 First discrete equation: asymmetric discrete Painlevé II
Consider the Bäcklund transformation $\mathcal{R}_{1}=\mathcal{T}_{1,-1,1}$ which has an inverse $\mathcal{R}_{1}^{-1}=\mathcal{T}_{-1,1,-1}$, then
$w_{n+1}$ $=\mathcal{R}_{1}(w_{n})=1-\frac{2zw_{n}}{\displaystyle zw_{n}^{\prime}+a_{n}w_{n}^{2}-(a_{n}+b_{n}-z)w_{n}+b_{n}},$ (3.3a)
$w_{n-1}$ $=\mathcal{R}_{1}^{-1}(w_{n})=1+\frac{2zw_{n}}{\displaystyle zw_{n}^{\prime}-a_{n}w_{n}^{2}+(a_{n}+b_{n}-z)w_{n}-b_{n}},$ (3.3b)
with $^{\prime}\equiv\mathrm{d}/\mathrm{d}{z}$, where $a_{n}$, $b_{n}$ and $c_{n}$ satisfy
$(a_{n+1},b_{n+1},c_{n+1})$ $=\mathcal{R}_{1}(a_{n},b_{n},c_{n})=\left(\tfrac{1}{2}\big(a_{n}-b_{n}+c_{n}+1\big),\tfrac{1}{2}\big(-a_{n}+b_{n}+c_{n}-1\big),a_{n}+b_{n}\right),$
$(a_{n-1},b_{n-1},c_{n-1})$ $=\mathcal{R}_{1}^{-1}(a_{n},b_{n},c_{n})=\left(\tfrac{1}{2}\big(a_{n}-b_{n}+c_{n}-1\big),\tfrac{1}{2}\big(-a_{n}+b_{n}+c_{n}+1\big),a_{n}+b_{n}\right),$
which have solution
$a_{n}$ $=\tfrac{1}{2}\big[n+\lambda+\rho+(-1)^{n}\varphi\big],\qquad b_{n}=-\tfrac{1}{2}\big[n+\lambda-\rho-(-1)^{n}\varphi\big],\qquad c_{n}=\rho-(-1)^{n}\varphi,$ (3.4)
with $\lambda$, $\rho$ and $\varphi$ arbitrary constants. From (3.3) we obtain
$\frac{1}{w_{n+1}-1}$ $+\frac{1}{2w_{n}}\frac{\mathrm{d}{w_{n}}}{{\mathrm{d}{z}}}+\frac{a_{n}w_{n}^{2}-(a_{n}+b_{n}-z)w_{n}+b_{n}}{2zw_{n}}=0,$ (3.5a)
$\frac{1}{w_{n-1}-1}$ $-\frac{1}{2w_{n}}\frac{\mathrm{d}{w_{n}}}{{\mathrm{d}{z}}}+\frac{a_{n}w_{n}^{2}-(a_{n}+b_{n}-z)w_{n}+b_{n}}{2zw_{n}}=0,$ (3.5b)
so adding gives
$\frac{1}{w_{n+1}-1}+\frac{1}{w_{n-1}-1}+\frac{a_{n}w_{n}^{2}-(a_{n}+b_{n}-z)w_{n}+b_{n}}{zw_{n}}=0,$
which is equation (3.4) in [32].
Setting $w_{n}=(q_{n}+1)/(q_{n}-1)$ and $w_{n\pm 1}=(q_{n\pm 1}+1)/(q_{n\pm 1}-1)$ in (3.3) gives
$q_{n+1}$ $+\frac{2}{1-q_{n}^{2}}\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}-\frac{2}{z}\,\frac{(n+\lambda)q_{n}+\rho+(-1)^{n}\varphi}{1-q_{n}^{2}}=0,$ (3.6a)
$q_{n-1}$ $-\frac{2}{1-q_{n}^{2}}\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}-\frac{2}{z}\,\frac{(n+\lambda)q_{n}+\rho+(-1)^{n}\varphi}{1-q_{n}^{2}}=0.$ (3.6b)
Hence by adding we obtain
$q_{n+1}+q_{n-1}=\frac{4}{z}\frac{(n+\lambda)q_{n}+\rho+(-1)^{n}\varphi}{1-q_{n}^{2}},$ (3.7)
which is known as the asymmetric dPII equation [65] and is a generalisation of dPII (1.6). Further $q_{n}$ satisfies the second-order equation
$\frac{\mathrm{d}^{2}{q_{n}}}{{\mathrm{d}{z}}^{2}}=\frac{q_{n}}{q_{n}^{2}-1}\left(\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}\right)^{\!2}-\frac{1}{z}\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}-\frac{a_{n}^{2}(q_{n}+1)^{2}-b_{n}^{2}(q_{n}-1)^{2}}{z^{2}(q_{n}^{2}-1)}-\frac{c_{n}(q_{n}^{2}-1)}{2z}+\tfrac{1}{4}q_{n}(q_{n}^{2}-1),$ (3.8)
with $a_{n}$, $b_{n}$ and $c_{n}$ given by (3.4).
In order to consider the alternating term in (3.7), i.e. the binary symmetry, we set $q_{2n}=x_{n}$ and $q_{2n+1}=y_{n}$ in (3.7) and obtain the discrete system
$x_{n+1}+x_{n}$ $=\frac{4}{z}\frac{(2n+\lambda+1)y_{n}+\rho-\varphi}{1-y_{n}^{2}},$ (3.9a)
$y_{n}+y_{n-1}$ $=\frac{4}{z}\frac{(2n+\lambda)x_{n}+\rho+\varphi}{1-x_{n}^{2}},$ (3.9b)
which is regarded as a discrete PIII equation (dPIII) since it admits a continuous limit to the third Painlevé equation (PIII)
$\frac{\mathrm{d}^{2}{w}}{{\mathrm{d}{z}}^{2}}=\frac{1}{w}\left(\frac{\mathrm{d}{w}}{{\mathrm{d}{z}}}\right)^{\!2}-\frac{1}{z}\frac{\mathrm{d}{w}}{{\mathrm{d}{z}}}+\frac{\alpha w^{2}+\beta}{w}+\gamma w^{3}+\frac{\delta}{w},$ (3.10)
with $\alpha$, $\beta$, $\gamma$ and $\delta$ constants [26, 63].
We note that
$\mathcal{R}_{1}^{2}(a,b,c)=(a+1,b-1,c),$
and so
$\mathcal{R}_{1}^{2n}(a,b,c)=(n+a,b-n,c).$
Therefore since
$(a_{0},b_{0},c_{0})$ $=\big(\tfrac{1}{2}(\lambda+\rho+\varphi),-\tfrac{1}{2}(\lambda-\rho-\varphi),\rho-\varphi\big),$
$(a_{1},b_{1},c_{1})$ $=\big(\tfrac{1}{2}(\lambda+\rho-\varphi+1),-\tfrac{1}{2}(\lambda-\rho+\varphi+1),\rho+\varphi\big),$
then
$(a_{2n},b_{2n},c_{2n})$ $=\Big(\tfrac{1}{2}\big(2n+\lambda+\rho+\varphi\big),-\tfrac{1}{2}\big(2n+\lambda-\rho-\varphi\big),\rho-\varphi\Big),$
$(a_{2n+1},b_{2n+1},c_{2n+1})$ $=\Big(\tfrac{1}{2}\big(2n+1+\lambda+\rho-\varphi\big),-\tfrac{1}{2}\big(2n+1+\lambda-\rho+\varphi\big),\rho+\varphi\Big).$
If $w_{0}(z)$ satisfies PV (1.2) with $\alpha=\tfrac{1}{8}(\lambda+\rho+\varphi)^{2}$, $\beta=-\tfrac{1}{8}(\lambda-\rho-\varphi)^{2}$ and $\gamma=\rho-\varphi$ then
$x_{n}=q_{2n}=\frac{\mathcal{R}_{1}^{2n}(w_{0})+1}{\mathcal{R}_{1}^{2n}(w_{0})-1},\qquad y_{n}=q_{2n+1}=\frac{\mathcal{R}_{1}^{2n+1}(w_{0})+1}{\mathcal{R}_{1}^{2n+1}(w_{0})-1},$ (3.11)
satisfy (3.9).
3.2 Second discrete equation
Consider the Bäcklund transformation $\mathcal{R}_{2}=\mathcal{T}_{-1,-1,-1}\circ\mathcal{T}_{-1,1,-1}$ which has the inverse $\mathcal{R}_{2}^{-1}=\mathcal{T}_{1,1,1}\circ\mathcal{T}_{1,-1,1}$, then
$w_{n+1}=\mathcal{R}_{2}(w_{n})$ $=\frac{1}{w_{n}}\left\{1+\frac{(a_{n}-b_{n}-c_{n}-1)(w_{n}-1)^{2}}{\displaystyle zw_{n}^{\prime}-a_{n}w_{n}^{2}+(2a_{n}-c_{n}-1+z)w_{n}-(a_{n}-c_{n}-1)}\right\},$ (3.12a)
$w_{n-1}=\mathcal{R}_{2}^{-1}(w_{n})$ $=\frac{1}{w_{n}}\left\{1-\frac{(a_{n}-b_{n}-c_{n}+1)(w_{n}-1)^{2}}{\displaystyle zw_{n}^{\prime}+a_{n}w_{n}^{2}-(2a_{n}-c_{n}+1+z)w_{n}+(a_{n}-c_{n}+1)}\right\}.$ (3.12b)
with $^{\prime}\equiv\mathrm{d}/\mathrm{d}{z}$, where $a_{n}$, $b_{n}$ and $c_{n}$ satisfy
$(a_{n+1},b_{n+1},c_{n+1})$ $=\mathcal{R}_{2}(a_{n},b_{n},c_{n})=\Big(\tfrac{1}{2}\big(a_{n}+b_{n}-c_{n}-1\big),\tfrac{1}{2}\big(a_{n}+b_{n}+c_{n}+1\big),-a_{n}+b_{n}+1\Big),$
$(a_{n-1},b_{n-1},c_{n-1})$ $=\mathcal{R}_{2}^{-1}(a_{n},b_{n},c_{n})=\Big(\tfrac{1}{2}\big(a_{n}+b_{n}-c_{n}+1\big),\tfrac{1}{2}\big(a_{n}+b_{n}+c_{n}-1\big),-a_{n}+b_{n}-1\Big),$
which have solution
$a_{n}$ $=-\tfrac{1}{2}\big[n+\lambda+\rho+(-1)^{n}\varphi\big],\qquad b_{n}=\tfrac{1}{2}\big[n+\lambda-\rho+(-1)^{n}\varphi\big],\qquad c_{n}=n+\lambda-(-1)^{n}\varphi,$ (3.13)
with $\lambda$, $\rho$ and $\varphi$ arbitrary constants. From (3.12) we obtain
$\frac{a_{n}-b_{n}-c_{n}-1}{w_{n}w_{n+1}-1}$ $-\frac{z}{(w_{n}-1)^{2}}\frac{\mathrm{d}{w_{n}}}{{\mathrm{d}{z}}}+a_{n}-\frac{z-c_{n}-1}{w_{n}-1}-\frac{z}{\left(w_{n}-1\right)^{2}}=0,$ (3.14a)
$\frac{a_{n}-b_{n}-c_{n}+1}{w_{n}w_{n-1}-1}$ $+\frac{z}{(w_{n}-1)^{2}}\frac{\mathrm{d}{w_{n}}}{{\mathrm{d}{z}}}+a_{n}-\frac{z-c_{n}+1}{w_{n}-1}-\frac{z}{\left(w_{n}-1\right)^{2}}=0,$ (3.14b)
Hence adding and using (3.13) gives the difference equation
$\frac{2n+2\lambda+1}{w_{n}w_{n+1}-1}+\frac{2n+2\lambda-1}{w_{n}w_{n-1}-1}+n+\lambda+\rho+(-1)^{n}\varphi-\frac{2\big[n+\lambda-(-1)^{n}\varphi\big]}{w_{n}-1}+\frac{2zw_{n}}{(w_{n}-1)^{2}}=0,$ (3.15)
which is equivalent to equation (2.8) in [62].
Setting $w_{n}=(q_{n}+1)/(q_{n}-1)$ and $w_{n\pm 1}=(q_{n\pm 1}+1)/(q_{n\pm 1}-1)$ in (3.12) gives
$\frac{2n+2\lambda+1}{q_{n+1}+q_{n}}$ $-\frac{z}{1-q_{n}^{2}}\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}-\tfrac{1}{2}z+\frac{\big[n+\lambda+(-1)^{n}\varphi\big]q_{n}+\rho}{1-q_{n}^{2}}=0,$ (3.16a)
$\frac{2n+2\lambda-1}{q_{n}+q_{n-1}}$ $+\frac{z}{1-q_{n}^{2}}\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}-\tfrac{1}{2}z+\frac{\big[n+\lambda+(-1)^{n}\varphi\big]q_{n}+\rho}{1-q_{n}^{2}}=0.$ (3.16b)
Hence by adding we obtain
$\frac{2n+2\lambda+1}{q_{n+1}+q_{n}}+\frac{2n+2\lambda-1}{q_{n}+q_{n-1}}=z-\frac{2\big[n+\lambda+(-1)^{n}\varphi\big]q_{n}+2\rho}{1-q_{n}^{2}},$ (3.17)
and then letting $\xi_{n}=n+\lambda$ gives
$\frac{\xi_{n+1}+\xi_{n}}{q_{n+1}+q_{n}}+\frac{\xi_{n}+\xi_{n-1}}{q_{n}+q_{n-1}}=z-\frac{2\big[\xi_{n}+(-1)^{n}\varphi\big]q_{n}+2\rho}{1-q_{n}^{2}}.$ (3.18)
This is equation (2.9) in [62], equation (3.25) in [69] and equation (13) in [28]. Following [62], we introduce a new variable $p_{n}$ through
$p_{n}=1-\frac{2}{z}\,\frac{\xi_{n+1}+\xi_{n}}{q_{n+1}+q_{n}},$
and then (3.18) becomes
$p_{n}+p_{n-1}$ $=\frac{4}{z}\,\frac{[\xi_{n}+(-1)^{n}\varphi]q_{n}+\rho}{1-q_{n}^{2}},$ (3.19a)
$q_{n+1}+q_{n}$ $=\frac{2}{z}\,\frac{\xi_{n+1}+\xi_{n}}{1-p_{n}},$ (3.19b)
which is system (2.10) in [62]. Ramani, Ohta and Grammaticos [62] state that “the system (3.19) describes a motion in the same parameter space as asymmetric dPII which is associated with the $A_{3}$ affine Weyl group. While asymmetric dPII describes a motion along the axes, the geometry of (3.19) is more complicated: the associated motion is a ‘staircase’ one”. To our knowledge [^2] This was confirmed in an email communication with Basil Grammaticos., there is no discussion in the literature as to which (continuous) Painlevé equation arises in the continuum limit for either (3.18) or (3.19).
In order to consider the alternating term in (3.17), i.e. the binary symmetry, setting $q_{2n}=x_{n}$ and $q_{2n+1}=y_{n}$ yields the discrete system
$\frac{4n+2\lambda+3}{x_{n+1}+y_{n}}+\frac{4n+2\lambda+1}{x_{n}+y_{n}}$ $=z-\frac{2(2n+\lambda-\varphi+1)y_{n}+2\rho}{1-y_{n}^{2}},$ (3.20a)
$\frac{4n+2\lambda+1}{x_{n}+y_{n}}+\frac{4n+2\lambda-1}{x_{n}+y_{n-1}}$ $=z-\frac{2(2n+\lambda+\varphi)x_{n}+2\rho}{1-x_{n}^{2}},$ (3.20b)
and so if $\xi_{n}=n+\lambda$ then
$\frac{\xi_{2n+2}+\xi_{2n+1}}{x_{n+1}+y_{n}}+\frac{\xi_{2n+1}+\xi_{2n}}{x_{n}+y_{n}}$ $=z-\frac{2(\xi_{2n+1}-\varphi)y_{n}+2\rho}{1-y_{n}^{2}},$
$\frac{\xi_{2n+1}+\xi_{2n}}{x_{n}+y_{n}}+\frac{\xi_{2n}+\xi_{2n-1}}{x_{n}+y_{n-1}}$ $=z-\frac{2(\xi_{2n}+\varphi)x_{n}+2\rho}{1-x_{n}^{2}}.$
Further $q_{n}$ satisfies the second-order equation
$\frac{\mathrm{d}^{2}{q_{n}}}{{\mathrm{d}{z}}^{2}}=\frac{q_{n}}{q_{n}^{2}-1}\left(\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}\right)^{\!2}-\frac{1}{z}\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}-\frac{a_{n}^{2}(q_{n}+1)^{2}-b_{n}^{2}(q_{n}-1)^{2}}{z^{2}(q_{n}^{2}-1)}-\frac{c_{n}(q_{n}^{2}-1)}{2z}+\tfrac{1}{4}q_{n}(q_{n}^{2}-1),$ (3.21)
with $a_{n}$, $b_{n}$ and $c_{n}$ given by (3.13).
Remark 3.2.
Since the Bäcklund transformation $\mathcal{R}_{2}=\mathcal{T}_{-1,-1,-1}\circ\mathcal{T}_{-1,1,-1}$ involves a sequence of two Bäcklund transformations then it might be expected that the resulting transformation is quadratic in $\mathrm{d}{w}/\mathrm{d}{z}$. Generally if one applies the Bäcklund transformation $\mathcal{T}_{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}}$ followed by the Bäcklund transformation $\mathcal{T}_{\sigma_{1},\sigma_{2},\sigma_{3}}$ then the resulting transformation is quadratic in $\mathrm{d}{w}/\mathrm{d}{z}$, but in some cases there is a simplification which happens in this case.
We note that
$\mathcal{R}_{2}^{2}(a,b,c)=(a-1,b+1,c+2),$
and so
$\mathcal{R}_{2}^{2n}(a,b,c)=(a-n,b+n,c+2n).$
Therefore since
$(a_{0},b_{0},c_{0})$ $=\big(-\tfrac{1}{2}(\lambda+\rho+\varphi),\tfrac{1}{2}(\lambda-\rho+\varphi),\lambda-\varphi\big),$
$(a_{1},b_{1},c_{1})$ $=\big(-\tfrac{1}{2}(\lambda+\rho-\varphi+1),\tfrac{1}{2}(\lambda-\rho-\varphi+1),\lambda+\varphi+1\big).$
then
$(a_{2n},b_{2n},c_{2n})$ $=\big(-\tfrac{1}{2}(2n+\lambda+\rho+\varphi),\tfrac{1}{2}(2n+\lambda-\rho+\varphi),2n+\lambda-\varphi\big),$
$(a_{2n+1},b_{2n+1},c_{2n+1})$ $=\big(-\tfrac{1}{2}(2n+\lambda+\rho-\varphi+1),\tfrac{1}{2}(2n+\lambda-\rho-\varphi+1),2n+\lambda+\varphi+1\big).$
If $w_{0}(z)$ satisfies PV (1.2) with $\alpha=\tfrac{1}{8}(\lambda+\rho+\varphi)^{2}$, $\beta=-\tfrac{1}{8}(\lambda-\rho+\varphi)^{2}$ and $\gamma=\lambda-\varphi$ then
$x_{n}=q_{2n}=\frac{\mathcal{R}_{2}^{2n}(w_{0})+1}{\mathcal{R}_{2}^{2n}(w_{0})-1},\qquad y_{n}=q_{2n+1}=\frac{\mathcal{R}_{2}^{2n+1}(w_{0})+1}{\mathcal{R}_{2}^{2n+1}(w_{0})-1},$ (3.22)
satisfy (3.20).
3.3 Third discrete equation: ternary discrete PI
Consider the Bäcklund transformation $\mathcal{R}_{3}=\mathcal{T}_{-1,1,-1}\circ\mathcal{T}_{1,-1,-1}\circ\mathcal{T}_{1,-1,1}$, which has an inverse $\mathcal{R}_{3}^{-1}=\mathcal{T}_{-1,1,1}$, then
$w_{n+1}$ $=\mathcal{R}_{3}(w_{n})=1-\frac{2zw_{n}}{\displaystyle zw_{n}^{\prime}+a_{n}w_{n}^{2}-(a_{n}+b_{n}-z)w_{n}+b_{n}},$ (3.23a)
$w_{n-1}$ $=\mathcal{R}_{3}^{-1}(w_{n})=1+\frac{2zw_{n}}{\displaystyle zw_{n}^{\prime}-a_{n}w_{n}^{2}+(a_{n}-b_{n}-z)w_{n}+b_{n}},$ (3.23b)
with $^{\prime}\equiv\mathrm{d}/\mathrm{d}{z}$, where $a_{n}$, $b_{n}$ and $c_{n}$ satisfy
$(a_{n+1},b_{n+1},c_{n+1})$ $=\mathcal{R}_{3}(a_{n},b_{n},c_{n})=\left(\tfrac{1}{2}\big(a_{n}-b_{n}+c_{n}+1\big),\tfrac{1}{2}\big(a_{n}-b_{n}-c_{n}+1\big),a_{n}+b_{n}\right),$ (3.24a)
$(a_{n-1},b_{n-1},c_{n-1})$ $=\mathcal{R}_{3}^{-1}(a_{n},b_{n},c_{n})=\left(\tfrac{1}{2}\big(a_{n}+b_{n}+c_{n}-1\big),-\tfrac{1}{2}\big(a_{n}+b_{n}-c_{n}-1\big),a_{n}-b_{n}\right).$ (3.24b)
These have solution
$a_{n}$ $=\tfrac{1}{3}n+\rho+\lambda\cos\big(\tfrac{2}{3}\pi n\big)+\tfrac{1}{3}\sqrt{3}\,\varphi\sin\big(\tfrac{2}{3}\pi n\big),$ (3.25a)
$b_{n}$ $=\tfrac{1}{3}+\sqrt{3}\,\lambda\sin\big(\tfrac{2}{3}\pi n\big)-\varphi\cos\big(\tfrac{2}{3}\pi n\big),$ (3.25b)
$c_{n}$ $=\tfrac{1}{3}n+\rho-2\lambda\cos\big(\tfrac{2}{3}\pi n\big)-\tfrac{2}{3}\sqrt{3}\,\varphi\sin\big(\tfrac{2}{3}\pi n\big),$ (3.25c)
with $\lambda$, $\rho$ and $\varphi$ arbitrary constants (the factor of $\sqrt{3}$ is for convenience), i.e.
$\begin{split}a_{3n}&=n+\lambda+\rho,\\ a_{3n+1}&=n-\tfrac{1}{2}\lambda+\tfrac{1}{2}\varphi+\rho+\tfrac{1}{3},\\ a_{3n+2}&=n-\tfrac{1}{2}\lambda-\tfrac{1}{2}\varphi+\rho+\tfrac{2}{3},\end{split}\qquad\begin{split}b_{3n}&=-\varphi+\tfrac{1}{3},\\ b_{3n+1}&=\tfrac{3}{2}\lambda+\tfrac{1}{2}\varphi+\tfrac{1}{3},\\ b_{3n+2}&=-\tfrac{3}{2}\lambda+\tfrac{1}{2}\varphi+\tfrac{1}{3},\end{split}\qquad\begin{split}c_{3n}&=n-2\lambda+\rho,\\ c_{3n+1}&=n+\lambda-\varphi+\rho+\tfrac{1}{3},\\ c_{3n+2}&=n+\lambda+\varphi+\rho+\tfrac{2}{3}.\end{split}$
which have a ternary symmetry in contrast to the binary symmetry for the parameters discussed in §3.1 and §3.2 above.
Setting $w_{n}=1+1/x_{n}$ and $w_{n\pm 1}=1+1/x_{n\pm 1}$ in (3.23) gives
$x_{n+1}$ $-\frac{1}{2x_{n}(x_{n}+1)}\frac{\mathrm{d}{x_{n}}}{{\mathrm{d}{z}}}+\frac{a_{n}}{2zx_{n}}-\frac{b_{n}}{2z(x_{n}+1)}+\frac{1}{2}=0,$ (3.26a)
$x_{n-1}$ $+\frac{1}{2x_{n}(x_{n}+1)}\frac{\mathrm{d}{x_{n}}}{{\mathrm{d}{z}}}+\frac{a_{n}}{2zx_{n}}+\frac{b_{n}}{2z(x_{n}+1)}+\frac{1}{2}=0.$ (3.26b)
Hence by adding we obtain
$x_{n}(x_{n+1}+x_{n-1}+1)+\frac{a_{n}}{z}=0,$ (3.27)
which is known as the ternary dPI equation [27, 69]. We remark that this derivation of (3.27) was given by two of us (Clarkson and Mitchell, in collaboration with Dzhamay and Hone), see [13].
Since equation (3.27) has a ternary symmetry, then it can be written as a system of three equations. Setting $x_{3n}=X_{n}$, $x_{3n+1}=Y_{n}$ and $x_{3n+2}=Z_{n}$ in (3.27) yields the discrete system
$X_{n}(Y_{n}+Z_{n-1}+1)+\frac{a_{3n}}{z}$ $=0,$ (3.28a)
$Y_{n}(Z_{n}+X_{n}+1)+\frac{a_{3n+1}}{z}$ $=0,$ (3.28b)
$Z_{n}(X_{n+1}+Y_{n}+1)+\frac{a_{3n+2}}{z}$ $=0.$ (3.28c)
Further $x_{n}$ satisfies the second-order equation
$\frac{\mathrm{d}^{2}{x_{n}}}{{\mathrm{d}{z}}^{2}}=\frac{2x_{n}+1}{2x_{n}(x_{n}+1)}\left(\frac{\mathrm{d}{x_{n}}}{{\mathrm{d}{z}}}\right)^{\!2}-\frac{1}{z}\frac{\mathrm{d}{x_{n}}}{{\mathrm{d}{z}}}-\frac{a_{n}^{2}(x_{n}+1)^{2}-b_{n}^{2}x_{n}^{2}}{2z^{2}x_{n}(x_{n}+1)}-\frac{c_{n}x_{n}(x_{n}+1)}{z}+\frac{x_{n}(x_{n}+1)(2x_{n}+1)}{2},$ (3.29)
with $a_{n}$, $b_{n}$ and $c_{n}$ given by (3.25).
We note that
$\mathcal{R}_{3}^{3}(a,b,c)=(a+1,b,c+1),$
and so
$\mathcal{R}_{3}^{3n}(a,b,c)=(a+n,b,c+n).$
Therefore since
$(a_{0},b_{0},c_{0})$ $=\big(\lambda+\rho,-\varphi+\tfrac{1}{3},-2\lambda+\rho\big),$
$(a_{1},b_{1},c_{1})$ $=\big(-\tfrac{1}{2}\lambda+\tfrac{1}{2}\varphi+\rho+\tfrac{1}{3},\tfrac{3}{2}\lambda+\tfrac{1}{2}\varphi+\tfrac{1}{3},\lambda-\varphi+\rho+\tfrac{1}{3}\big),$
$(a_{2},b_{2},c_{2})$ $=\big(-\tfrac{1}{2}\lambda-\tfrac{1}{2}\varphi+\rho+\tfrac{2}{3},-\tfrac{3}{2}\lambda+\tfrac{1}{2}\varphi+\tfrac{1}{3},\lambda+\varphi+\rho+\tfrac{2}{3}\big),$
then
$(a_{3n},b_{3n},c_{3n})$ $=\big(n+\lambda+\rho,-\varphi+\tfrac{1}{3},n-2\lambda+\rho\big),$
$(a_{3n+1},b_{3n+1},c_{3n+1})$ $=\big(n-\tfrac{1}{2}\lambda+\tfrac{1}{2}\varphi+\rho+\tfrac{1}{3},\tfrac{3}{2}\lambda+\tfrac{1}{2}\varphi+\tfrac{1}{3},n+\lambda-\varphi+\rho+\tfrac{1}{3}\big),$
$(a_{3n+2},b_{3n+2},c_{3n+2})$ $=\big(n-\tfrac{1}{2}\lambda-\tfrac{1}{2}\varphi+\rho+\tfrac{2}{3},-\tfrac{3}{2}\lambda+\tfrac{1}{2}\varphi+\tfrac{1}{3},n+\lambda+\varphi+\rho+\tfrac{2}{3}\big).$
If $w_{0}$ satisfies PV (1.2) with $\alpha=\tfrac{1}{2}(\lambda+\rho)^{2}$, $\beta=-\tfrac{1}{2}(\varphi-\tfrac{1}{3})^{2}$ and $\gamma=-2\lambda+\rho$ then
$X_{n}=x_{3n}=\frac{1}{\mathcal{R}_{3}^{3n}(w_{0})-1},\qquad Y_{n}=x_{3n+1}=\frac{1}{\mathcal{R}_{3}^{3n+1}(w_{0})-1},\qquad Z_{n}=x_{3n+2}=\frac{1}{\mathcal{R}_{3}^{3n+2}(w_{0})-1},$
satisfy (3.28).
Remarks 3.3.
  1. 1.
    In terms of $w$, the transformation (3.23a) is the same as $\mathcal{T}_{1,-1,1}(w)$. However, the parameters $(a,b,c)$ map differently since
    $\mathcal{T}_{1,-1,1}(a,b,c)=\big(\tfrac{1}{2}(a-b+c-1),\tfrac{1}{2}(-a+b+c-1),a+b\big).$
    It is straightforward to show that $\mathcal{T}_{1,-1,1}\circ\mathcal{T}_{-1,1,1}(w)=w$, whereas $\mathcal{T}_{-1,1,1}\circ\mathcal{T}_{1,-1,1}(w)\not=w$, so $\mathcal{T}_{1,-1,1}$ is not the inverse of $\mathcal{T}_{-1,1,1}$.
  2. 2.
    It is interesting to note that given a solution $w_{0}(z)$ of PV (1.2) with $(\alpha_{0},\beta_{0},\gamma_{0})=\big(\tfrac{1}{2}a_{0}^{2},-\tfrac{1}{2}b_{0}^{2},c_{0}\big)$, the transformation $\mathcal{R}_{1}$ (3.3a) maps it to the solution $w_{1}(z)$ with $(\alpha_{1},\beta_{1},\gamma_{1})=\big(\tfrac{1}{2}a_{1}^{2},-\tfrac{1}{2}b_{1}^{2},c_{1}\big)$ and the transformation $\mathcal{R}_{3}$ (3.23a) maps it to the solution $w_{3}(z)$ with $(\alpha_{3},\beta_{3},\gamma_{3})=\big(\tfrac{1}{2}a_{3}^{2},-\tfrac{1}{2}b_{3}^{2},c_{3}\big)$ where
    $w_{1}(z)=w_{3}(z),\qquad a_{1}=a_{3},\qquad b_{1}=-b_{3},\qquad c_{1}=c_{3},$
    i.e. both transformations map $w_{0}(z)$ to the same solution of PV. However applying the transformation a second time gives different solutions since
    $\mathcal{R}_{1}^{2}(a_{0},b_{0},c_{0})$ $=(a_{0}+1,b_{0}-1,c_{0}),$
    $\mathcal{R}_{3}^{2}(a_{0},b_{0},c_{0})$ $=\big(\tfrac{1}{2}(a_{0}+b_{0}+c_{0}+1),-\tfrac{1}{2}(a_{0}+b_{0}-c_{0}-1),a_{0}-b_{0}+1\big),$
    so the transformations $\mathcal{R}_{1}$ (3.3) and $\mathcal{R}_{3}$ (3.23) will generate different hierarchies of solutions.
  3. 3.
    The symmetric case of equation (3.27), i.e. when $\lambda=\varphi=0$,
    $x_{n}(x_{n+1}+x_{n-1}+1)+\frac{\rho+\tfrac{1}{3}n}{z}=0,$ (3.30)
    is a discrete PI [61, equation (3.6)], whilst the full equation, which was derived from a discrete dressing transformation [27], is a discrete form of PIV (1.10).
  4. 4.
    The ternary dPI (3.27) is associated with PV (1.2), though the asymmetric generalisation of dPI (1.4), i.e.
    $x_{n+1}+x_{n}+x_{n-1}=1+\frac{\lambda n+\rho+(-1)^{n}\varphi}{x_{n}},$ (3.31)
    with $\lambda$, $\rho$ and $\varphi$ constants, is associated with PIV (1.10) cf. [20, 42], and can be derived from Bäcklund transformations of PIV [14, 19]. We remark that the continuum limit of (3.31) with $\varphi\not=0$ is PII (1.7) rather than PI (1.5), as shown in [15].
  5. 5.
    Recently Clarkson et al. [13] studied the special case of the ternary dPI equation given by
    $v_{n}(v_{n+1}+v_{n-1}+1)=(n+1)\varepsilon,\qquad\varepsilon>0,$ (3.32)
    with $v_{-1}=0$, which arises in quantum minimal surfaces [5, 34] and found a unique solution. The expressions for $v_{3n}$, $v_{3n+1}$ and $v_{3n+2}$ were written explicitly as ratios of Wronskian determinants in terms of the modified Bessel functions $K_{1/6}(z)$ and $K_{5/6}(z)$ and were all different showing the three-fold structure of the solutions.
  6. 6.
    Felder and Hoppe [17] show that ternary dPI equation (3.32) also arises in connection with orthogonal polynomials in the complex plane with respect to the weight
    $w(z;t)=\exp\left\{-t|z|^{2}+\tfrac{1}{3}\mathrm{i}(z^{3}+\overline{z}^{3})\right\},\qquad t>0,$
    with $z=x+\mathrm{i}y\in\mathbb{C}$, and has the same unique solution as in [13]. Teodorescu et al. [68], see also [67, equation (26)], show that equation (3.32) arises in the theory of random normal matrices [74], which motivated the study in [17].
  7. 7.
    A study of the asymptotics of solutions of (3.32) is given by Aptekarev and Novokshenov [2].
3.4 Fourth discrete equation: new equation with ternary symmetry
It is possible to derive a discrete equation relating the three solutions $x_{n}$, $x_{n+2}$ and $x_{n-2}$ of the ternary dPI equation (3.27) as illustrated in the following Lemma.
Lemma 3.4.
Suppose that $x_{n}$, $x_{n+2}$ and $x_{n-2}$ are solutions of the ternary dPI equation (3.27), then they satisfy the discrete equation
$\frac{a_{n+1}}{x_{n+2}+x_{n}+1}+\frac{a_{n-1}}{x_{n}+x_{n-2}+1}=z+\frac{a_{n}}{x_{n}},$ (3.33)
where $a_{n}$ is given by (3.25a), i.e.
$a_{n}=\tfrac{1}{3}n+\rho+\lambda\cos\big(\tfrac{2}{3}\pi n\big)+\tfrac{1}{3}\sqrt{3}\,\varphi\sin\big(\tfrac{2}{3}\pi n\big),$ (3.34)
with $\lambda$, $\rho$ and $\varphi$ arbitrary constants.
Proof.
Letting $n\to n-1$ in equation (3.27) gives
$x_{n-1}(x_{n}+x_{n-2}+1)+\frac{a_{n-1}}{z}=0.$
Solving this equation and (3.27) for $x_{n+1}$ and $x_{n-1}$ gives
$x_{n+1}=\frac{a_{n-1}}{z(x_{n}+x_{n-2}+1)}-1-\frac{a_{n}}{zx_{n}},\qquad x_{n-1}=-\frac{a_{n-1}}{z(x_{n}+x_{n-2}+1)}.$
Letting $n\to n+2$ in the expression for $x_{n-1}$ gives
$x_{n+1}=-\frac{a_{n+1}}{z(x_{n+2}+x_{n}+1)}.$
Equating the two expressions for $x_{n+1}$ gives the result. ∎
An alternative proof is to consider the Bäcklund transformation $\mathcal{R}_{4}=\mathcal{R}_{3}^{2}$, with $\mathcal{R}_{4}^{-1}=\mathcal{T}_{-1,1,1}^{2}$, then
$w_{n+2}=\mathcal{R}_{4}(w_{n})$ $=\frac{1}{w_{n}}\left\{1+\frac{(a_{n}-b_{n}+c_{n}+1)(w_{n}-1)^{2}}{\displaystyle zw_{n}^{\prime}+a_{n}w_{n}^{2}-(2a_{n}+c_{n}+1-z)w_{n}+a_{n}+c_{n}+1}\right\},$ (3.35a)
$w_{n-2}=\mathcal{R}_{4}^{-1}(w_{n})$ $=\frac{1}{w_{n}}\left\{1+\frac{(a_{n}+b_{n}+c_{n}-1)(w_{n}-1)^{2}}{\displaystyle zw_{n}^{\prime}-a_{n}w_{n}^{2}+(2a_{n}+c_{n}-1-z)w_{n}-a_{n}-c_{n}+1}\right\},$ (3.35b)
with $^{\prime}\equiv\mathrm{d}/\mathrm{d}z$ and $a_{n}$, $b_{n}$ and $c_{n}$ are given by (3.25). Letting $w_{n}=1+1/x_{n}$ and $w_{n\pm 2}=1+1/x_{n\pm 2}$ gives
$\frac{a_{n}-b_{n}+c_{n}+1}{x_{n+2}+x_{n}+1}$ $=z-\frac{z}{x_{n}(x_{n}+1)}\frac{\mathrm{d}{x_{n}}}{{\mathrm{d}{z}}}+\frac{(a_{n}-b_{n})x_{n}+a_{n}}{x_{n}(x_{n}+1)},$ (3.36a)
$\frac{a_{n}+b_{n}+c_{n}-1}{x_{n}+x_{n-2}+1}$ $=z+\frac{z}{x_{n}(x_{n}+1)}\frac{\mathrm{d}{x_{n}}}{{\mathrm{d}{z}}}+\frac{(a_{n}+b_{n})x_{n}+a_{n}}{x_{n}(x_{n}+1)},$ (3.36b)
and so by adding we obtain
$\frac{a_{n+1}}{x_{n+2}+x_{n}+1}+\frac{a_{n-1}}{x_{n}+x_{n-2}+1}=z+\frac{a_{n}}{x_{n}},$
since from (3.24), $a_{n+1}=\tfrac{1}{2}(a_{n}-b_{n}+c_{n}+1)$ and $a_{n-1}=\tfrac{1}{2}(a_{n}+b_{n}+c_{n}-1)$. The parameters $\widetilde{a}_{n}$, $\widetilde{b}_{n}$ and $\widetilde{c}_{n}$ satisfy
$\begin{split}\widetilde{a}_{n+1}&=\tfrac{1}{2}(\widetilde{a}_{n}+\widetilde{b}_{n}+\widetilde{c}_{n}+1),\\ \widetilde{b}_{n+1}&=-\tfrac{1}{2}(\widetilde{a}_{n}+\widetilde{b}_{n}-\widetilde{c}_{n}-1),\\ \widetilde{c}_{n+1}&=\widetilde{a}_{n}-\widetilde{b}_{n}+1,\end{split}\qquad\quad\begin{split}\widetilde{a}_{n-1}&=\tfrac{1}{2}(\widetilde{a}_{n}-\widetilde{b}_{n}+\widetilde{c}_{n}-1),\\ \widetilde{b}_{n-1}&=\tfrac{1}{2}(\widetilde{a}_{n}-\widetilde{b}_{n}-\widetilde{c}_{n}+1),\\ \widetilde{c}_{n-1}&=\widetilde{a}_{n}+\widetilde{b}_{n}-1,\end{split}$
which have solution
$\widetilde{a}_{n}$ $=\tfrac{2}{3}n+\rho+\lambda\cos\big(\tfrac{2}{3}\pi n\big)+\tfrac{1}{3}\sqrt{3}\,\varphi\sin\big(\tfrac{2}{3}\pi n\big),$
$\widetilde{b}_{n}$ $=\tfrac{1}{3}-\sqrt{3}\,\lambda\sin\big(\tfrac{2}{3}\pi n\big)+\varphi\cos\big(\tfrac{2}{3}\pi n\big),$
$\widetilde{c}_{n}$ $=\tfrac{2}{3}n+\rho-2\lambda\cos\big(\tfrac{2}{3}\pi n\big)-\tfrac{2}{3}\sqrt{3}\,\varphi\sin\big(\tfrac{2}{3}\pi n\big),$
with $\lambda$, $\rho$ and $\varphi$ arbitrary constants. We remark that the Bäcklund transformation (3.35) has similar structure to the Bäcklund transformation $\mathcal{R}_{2}$ (3.12), except the parameters have a ternary symmetry rather than a binary symmetry. We note that since $\mathcal{R}_{3}=\mathcal{T}_{-1,1,-1}\circ\mathcal{T}_{1,-1,-1}\circ\mathcal{T}_{1,-1,1}$, then
$\mathcal{R}_{4}=\mathcal{R}_{3}^{2}$ $=\mathcal{T}_{-1,1,-1}\circ\mathcal{T}_{1,-1,-1}\circ\mathcal{T}_{1,-1,1}\circ\mathcal{T}_{-1,1,-1}\circ\mathcal{T}_{1,-1,-1}\circ\mathcal{T}_{1,-1,1}$
$=\mathcal{T}_{-1,1,-1}\circ\mathcal{T}_{1,-1,-1}\circ\mathcal{T}_{1,-1,-1}\circ\mathcal{T}_{1,-1,1}$
as $\mathcal{T}_{1,-1,1}\circ\mathcal{T}_{-1,1,-1}=\mathcal{I}$, the identity.
Corollary 3.5.
If $x_{n}$ satisfies ternary dPI equation (3.27) and $a_{n}$ is given by (3.34), then $q_{n}=x_{2n}$ and $p_{n}=x_{2n+1}$ satisfy
$\frac{a_{2n+1}}{q_{n+1}+q_{n}+1}+\frac{a_{2n-1}}{q_{n}+q_{n-1}+1}$ $=z+\frac{a_{2n}}{q_{n}},$ (3.37a)
$\frac{a_{2n+2}}{p_{n+1}+p_{n}+1}+\frac{a_{2n}}{p_{n}+p_{n-1}+1}$ $=z+\frac{a_{2n+1}}{p_{n}},$ (3.37b)
respectively. Further $X_{n}=x_{3n}$, $Y_{n}=x_{3n+1}$ and $Z_{n}=x_{3n+2}$ satisfy the system
$\frac{a_{3n+1}}{X_{n}+Z_{n}+1}+\frac{a_{3n-1}}{X_{n}+Y_{n-1}+1}$ $=z+\frac{a_{3n}}{X_{n}},$ (3.38a)
$\frac{a_{3n+2}}{X_{n+1}+Y_{n}+1}+\frac{a_{3n}}{Y_{n}+Z_{n-1}+1}$ $=z+\frac{a_{3n+1}}{Y_{n}},$ (3.38b)
$\frac{a_{3n+3}}{Y_{n+1}+Z_{n}+1}+\frac{a_{3n+1}}{X_{n}+Z_{n}+1}$ $=z+\frac{a_{3n+2}}{Z_{n}}.$ (3.38c)
Proof.
Letting $n\to 2n$ and $n\to 2n+1$ in (3.33) and then setting $q_{n}=x_{2n}$ and $p_{n}=x_{2n+1}$, respectively, gives (3.37). Similarly for the system (3.38), which is equivalent to the system (3.28). ∎
3.5 Differential-difference equations
In the previous subsections, we have obtained discrete equations by eliminating the derivative between two equations, i.e. by adding the equations in (3.5), (3.14) and (3.26). If instead we subtract the equations then we obtain differential-difference equations.
  1. (i)
    Subtracting the equations in (3.6) gives the modified Volterra lattice [73]
    $\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}=\tfrac{1}{4}(q_{n}^{2}-1)(q_{n+1}-q_{n-1}).$
    which is a special case of equation (3.39) below.
  2. (ii)
    Subtracting the equations in (3.16) gives
    $z\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}=\tfrac{1}{2}(1-q_{n}^{2})\left\{\frac{2n+2\lambda+1}{q_{n+1}+q_{n}}-\frac{2n+2\lambda-1}{q_{n}+q_{n-1}}\right\}.$
    which is a non-autonomous special case of equation (3.40) below.
  3. (iii)
    Subtracting the equations in (3.26) gives
    $z\frac{\mathrm{d}{x_{n}}}{{\mathrm{d}{z}}}=zx_{n}(x_{n}+1)(x_{n+1}-x_{n-1})-{b_{n}x_{n}},$
    with $b_{n}$ given by (3.25b).
  4. (iv)
    Subtracting the equations in (3.36) gives
    $z\frac{\mathrm{d}{x_{n}}}{{\mathrm{d}{z}}}=-x_{n}(x_{n}+1)\left\{\frac{a_{n+1}}{x_{n+2}+x_{n}+1}-\frac{a_{n-1}}{x_{n}+x_{n-2}+1}-\right\}-b_{n}x_{n},$
    with $a_{n}$ and $b_{n}$ given by (3.25). Consequently $q_{n}=\tfrac{1}{2}(x_{2n}+1)$ and $p_{n}=\tfrac{1}{2}(x_{2n+1}+1)$ satisfy
    $z\frac{\mathrm{d}{q_{n}}}{{\mathrm{d}{z}}}$ $=(1-q_{n}^{2})\left\{\frac{a_{2n+1}}{q_{n+1}+q_{n}}-\frac{a_{2n-1}}{q_{n}+q_{n-1}}\right\}-b_{2n}(q_{n}-1),$
    $z\frac{\mathrm{d}{p_{n}}}{{\mathrm{d}{z}}}$ $=(1-p_{n}^{2})\left\{\frac{a_{2n+2}}{p_{n+1}+p_{n}}-\frac{a_{2n}}{p_{n}+p_{n-1}}\right\}-b_{2n+1}(p_{n}-1),$
    which are non-autonomous special cases of equation (3.40) below.
Remark 3.6.
Yamilov [76] gives a classification of (autonomous) Volterra-type equations. The first two are:
$\frac{\mathrm{d}{u_{n}}}{{\mathrm{d}{z}}}$ $=P(u_{n})(u_{n+1}-u_{n-1}),$ (3.39)
$\frac{\mathrm{d}{u_{n}}}{{\mathrm{d}{z}}}$ $=P(u_{n}^{2})\left(\frac{1}{u_{n+1}+u_{n}}-\frac{1}{u_{n}+u_{n-1}}\right),$ (3.40)
where $P(u)=au^{2}+bu+c$, with $a$, $b$ and $c$ constants.
4 Rational solutions of PV
In this section we describe the rational solutions of PV (1.2), which are classified in the following theorem.
Theorem 4.1.
Equation (1.2) has a rational solution if and only if one of the following holds:
  1. (i)
    $\alpha=\tfrac{1}{2}m^{2}$, $\beta=-\tfrac{1}{2}(m+2n+1+\kappa)^{2}$, $\gamma=\kappa$, for $m\geq 1$;
  2. (ii)
    $\alpha=\tfrac{1}{2}(m+\kappa)^{2}$, $\beta=-\tfrac{1}{2}(n+\varepsilon\kappa)^{2}$, $\gamma=m+\varepsilon n$, with $\varepsilon=\pm 1$, provided that $m\not=0$ or $n\not=0$;
  3. (iii)
    $\alpha=\tfrac{1}{2}(m+\tfrac{1}{2})^{2}$, $\beta=-\tfrac{1}{2}(n+\tfrac{1}{2})^{2}$, $\gamma=\kappa$, provided that $m\not=0$ or $n\not=0$,
where $m,n\in\mathbb{Z}$ and $\kappa$ is an arbitrary constant, together with the solutions obtained through the symmetries
$\mathcal{S}_{1}:\qquad$ ${w_{1}}({z};{\alpha_{1}},{\beta_{1}},{\gamma_{1}},-\tfrac{1}{2})={w(-z;\alpha,\beta,\gamma,-\tfrac{1}{2})},$ $({\alpha_{1}},{\beta_{1}},{\gamma_{1}},-\tfrac{1}{2})=(\alpha,\beta,-\gamma,-\tfrac{1}{2}),$ (4.1)
$\mathcal{S}_{2}:\qquad$ ${w_{2}}({z};{\alpha_{2}},{\beta_{2}},{\gamma_{3}},-\tfrac{1}{2})=\frac{1}{w(z;\alpha,\beta,\gamma,-\tfrac{1}{2})},$ $({\alpha_{2}},{\beta_{2}},{\gamma_{3}},-\tfrac{1}{2})=(-\beta,-\alpha,-\gamma,-\tfrac{1}{2}),$ (4.2)
where $w(z;\alpha,\beta,\gamma,-\tfrac{1}{2})$ is a solution of PV (1.2).
Proof.
See Kitaev, Law and McLeod [40, Theorem 1.1]; also [33, Theorem 40.3].∎
Remarks 4.2.
  1. 1.
    Kitaev, Law and McLeod [40, Theorem 1.1] give four cases, though their cases (I) and (II) are related by the symmetry (4.2). Further Kitaev, Law and McLeod [40, Theorem 1.2] prove that if $\gamma\not\in\mathbb{Z}$ in case (i), then the rational solution is unique and if $\gamma\in\mathbb{Z}$ then there are at most two rational solutions, as illustrated in [12]; see also [3, 4].
  2. 2.
    Rational solutions in case (i) are expressed in terms of generalised Laguerre polynomials, which are discussed in §5.1, and in cases (ii) and (iii) in terms of generalised Umemura polynomials, discussed in §5.2.
4.1 Generalised Laguerre polynomials
Rational solutions in case (i) of Theorem 4.1 are expressed in terms of generalised Laguerre polynomials, which are written in terms of a determinant of Laguerre polynomials, or equivalently as a Wronskian, see [12, §3]. These rational solutions are special cases of the solutions of PV (1.2) expressible in terms of the Kummer functions $M(a,b,z)$ and $U(a,b,z)$ discussed in [43, 57] since
$U(-n,\alpha+1,z)=(-1)^{n}(\alpha+1)_{n}M(-n,\alpha+1,z)=(-1)^{n}n!\,L_{n}^{(\alpha)}(z),$ (4.3)
with $L_{n}^{(\alpha)}(z)$ the Laguerre polynomial
$L_{n}^{(\alpha)}(z)=\frac{z^{-\alpha}\,\mathrm{e}^{z}}{n!}\frac{\mathrm{d}^{n}{}}{{\mathrm{d}{z}}^{n}}\left(z^{n+\alpha}\,\mathrm{e}^{-z}\right),\qquad n\geq 0,$ (4.4)
see [54, equation (13.6.19)]. First we define the generalised Laguerre polynomial $T_{m,n}^{(\mu)}(z)$; properties of it are discussed in [6, 12].
Definition 4.3.
The generalised Laguerre polynomial $T_{m,n}^{(\mu)}(z)$, which is a polynomial of degree $(m+1)n$, is defined by
$T_{m,n}^{(\mu)}(z)=\det\left[\frac{\mathrm{d}^{j+k}{}}{{\mathrm{d}{z}}^{j+k}}L_{m+n}^{(\mu+1)}(z)\right]_{j,k=0}^{n-1},\qquad m\geq 0,\quad n\geq 1,$ (4.5)
where $L_{n}^{(\mu)}(z)$ is the Laguerre polynomial (4.4).
For case (i) in Theorem 4.1, Kitaev, Law and McLeod [40] state that rational solutions arise when
$\alpha=\tfrac{1}{2}\ell^{2},\qquad\beta=-\tfrac{1}{2}(k+\varepsilon\gamma)^{2},\qquad\varepsilon=\pm 1,$
where $\ell>0$ and $k+\ell$ is odd, with $k,\ell\in\mathbb{Z}$, and $\beta\not=0$ when $|k|<\ell$. Since the sign of $\gamma$ can be changed by letting $z\to-z$, we set $\varepsilon=1$, without loss of generality. There are three cases (a), $k>\ell$, (b), $|k|\leq\ell$ and (c), $k>\ell$, which are given in the following theorem.
Theorem 4.4.
Suppose that $T_{m,n}^{(\mu)}(z)$ is the generalised Laguerre polynomial given by (4.5) and $L_{n}^{(\alpha)}(z)$ is the Laguerre polynomial, then PV (1.2) has the following three sets of rational solutions.
  1. (a)
    If $\ell=m$ and $k=m+2n+1$, with $m\geq 1$, so $k+\ell=2m+2n+1$, then for $n\geq 1$
    ${w}_{m,n}^{(\mu)}(z)=\frac{T_{m-1,n}^{(\mu)}(z)\,T_{m-1,n+1}^{(\mu)}(z)}{T_{m,n}^{(\mu)}(z)\,T_{m-2,n+1}^{(\mu)}(z)},\qquad(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}m^{2},-\tfrac{1}{2}(m+2n+1+\mu)^{2},\mu\big).$ (4.6a)
    In the case when $n=0$, then for $m\geq 1$
    ${w}_{m,0}^{(\mu)}(z)=\frac{L_{m}^{(\mu+1)}(z)}{L_{m-1}^{(\mu+1)}(z)},\qquad(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}m^{2},-\tfrac{1}{2}(m+1+\mu)^{2},\mu\big).$ (4.6b)
  2. (b)
    If $\ell=m+n+1$ and k=mnk=m-n, with $m\geq 0$, so $k+\ell=2m+1$, then for $m\geq 1$
    ${v}_{m,n}^{(\mu)}(z)=\frac{m-n+\mu}{m+n+1}\,\frac{T_{m,n}^{(\mu-2n)}(z)\,T_{m-1,n+1}^{(\mu-2n-2)}(z)}{T_{m-1,n}^{(\mu-2n)}(z)\,T_{m,n+1}^{(\mu-2n-2)}(z)},\quad(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(m+n+1)^{2},-\tfrac{1}{2}(m-n+\mu)^{2},\mu\big).$ (4.7a)
    In the case when $m=0$ and $n\geq 1$ then
    ${v}_{0,n}^{(\mu)}(z)$ $={\frac{n-\mu}{n+1}\,\frac{L_{n}^{(-\mu-1)}(-z)}{L_{n+1}^{(-\mu-1)}(-z)}},$ $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(n+1)^{2},-\tfrac{1}{2}(n-\mu)^{2},\mu\big),$ (4.7b)
    $\widetilde{v}_{0,n}^{(\mu)}(z)$ $=\frac{L_{n+1}^{(-1-\mu)}(-z)}{L_{n}^{(-1-\mu)}(-z)},$ $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(n+1)^{2},-\tfrac{1}{2}(n-\mu)^{2},\mu+2\big).$ (4.7c)
    In the case when $n=0$ and $m\geq 0$ then
    ${v}_{m,0}^{(\mu)}(z)$ $=\frac{m+\mu}{m+1}\,\frac{L^{(\mu-1)}_{m}(z)}{L^{(\mu-1)}_{m+1}(z)},$ $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(m+1)^{2},-\tfrac{1}{2}(m+\mu)^{2},\mu\big),$ (4.7d)
    $\widetilde{v}_{m,0}^{(\mu)}(z)$ $=\frac{L^{(\mu-1)}_{m+1}(z)}{L^{(\mu-1)}_{m}(z)},$ $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(m+1)^{2},-\tfrac{1}{2}(m+\mu)^{2},\mu-2\big).$ (4.7e)
  3. (c)
    If $\ell=m$ and k=m2n1k=-m-2n-1, with $m\geq 1$ and $n\geq 0$, so $k+\ell=-2m-1$, then ${u}_{m,n}^{(\mu)}(z)={w}_{m,n}^{(-\mu)}(-z)$, i.e.
    ${u}_{m,n}^{(\mu)}(z)=\frac{T_{n-1,m}^{(\mu-2m-2n)}(z)\,T_{n,m}^{(\mu-2m-2n-2)}(z)}{T_{n,m-1}^{(\mu-2m-2n)}(z)\,T_{n-1,m+1}^{(\mu-2m-2n-2)}(z)},\qquad(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}m^{2},-\tfrac{1}{2}(m+2n+1-\mu)^{2},\mu\big).$ (4.8a)
    In the case when $n=0$, then for $m\geq 1$
    ${u}_{m,0}^{(\mu)}(z)=\frac{L_{m}^{(1-\mu)}(-z)}{L_{m-1}^{(1-\mu)}(-z)},\qquad(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}m^{2},-\tfrac{1}{2}(m+1-\mu)^{2},\mu\big).$ (4.8b)
Proof.
See Clarkson and Dunning [12]. The solutions when either $m=0$ or $n=0$ follow from recurrence relations for the Laguerre polynomial, cf. [54, §18.9]. ∎
Remark 4.5.
The rational solutions in Theorem 4.4 can also be written as logarithmic derivatives
$w_{m,n}^{(\mu)}(z)$ $=\frac{z}{m}\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{m-2,n+1}^{(\mu)}(z)}{T_{m,n}^{(\mu)}(z)}-\frac{z-m-2n-1-\mu}{m},$ (4.9a)
$v_{m,n}^{(\mu)}(z)$ $=\frac{z}{m+n+1}\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{m-1,n}^{(\mu-2n)}(z)}{T_{m,n+1}^{(\mu-2n-2)}(z)}+1,$ (4.9b)
$u_{m,n}^{(\mu)}(z)$ $=\frac{z}{m}\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{n,m-1}^{(\mu-2m-2n)}}{T_{n-1,m+1}^{(\mu-2m-2n-2)}}+\frac{z+m+2n+1-\mu}{m}.$ (4.9c)
Recall that solutions of the discrete equations (3.7) and (3.17) also satisfy the second-order equation
$\frac{\mathrm{d}^{2}{q}}{{\mathrm{d}{z}}^{2}}=\frac{q}{q^{2}-1}\left(\frac{\mathrm{d}{q}}{{\mathrm{d}{z}}}\right)^{\!2}-\frac{1}{z}\frac{\mathrm{d}{q}}{{\mathrm{d}{z}}}-\frac{2\alpha(q+1)^{2}+2\beta(q-1)^{2}}{z^{2}(q^{2}-1)}-\frac{\gamma(q^{2}-1)}{2z}+\tfrac{1}{4}q(q^{2}-1),$ (4.10)
and solutions of the ternary dPI (3.27) also satisfy the second-order equation
$\frac{\mathrm{d}^{2}{x}}{{\mathrm{d}{z}}^{2}}=\frac{2x+1}{2x(x+1)}\left(\frac{\mathrm{d}{x}}{{\mathrm{d}{z}}}\right)^{\!2}-\frac{1}{z}\frac{\mathrm{d}{x}}{{\mathrm{d}{z}}}-\frac{\alpha(x+1)^{2}+\beta x^{2}}{z^{2}x(x+1)}-\frac{\gamma x(x+1)}{z}+\frac{x(x+1)(2x+1)}{2}.$ (4.11)
In the following Lemma we derive the solutions of equations (4.10) and (4.11) in terms of generalised Laguerre polynomials.
Lemma 4.6.
If $w_{m,n}^{(\mu)}(z)$, $v_{m,n}^{(\mu)}(z)$ and $u_{m,n}^{(\mu)}(z)$ are rational solutions of PV defined in Theorem 4.4 and $T_{m,n}^{(\mu)}(z)$ is the generalised Laguerre polynomial given by (4.5), then
$\mathcal{W}_{m,n}^{(\mu)}(z)=\frac{w_{m,n}^{(\mu)}(z)+1}{w_{m,n}^{(\mu)}(z)-1}$ $=1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{m-1,n}^{(\mu+1)}(z)}{T_{m-1,n+1}^{(\mu-1)}(z)},$ (4.12a)
$\mathcal{V}_{m,n}^{(\mu)}(z)=\frac{v_{m,n}^{(\mu)}(z)+1}{v_{m,n}^{(\mu)}(z)-1}$ $=-1+\frac{2\mu}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{m-1,n+1}^{(\mu-2n-1)}(z)}{T_{m,n}^{(\mu-2n-1)}(z)},$ (4.12b)
$\mathcal{U}_{m,n}^{(\mu)}(z)=\frac{u_{m,n}^{(\mu)}(z)+1}{u_{m,n}^{(\mu)}(z)-1}$ $=1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{n,m}^{(\mu-2m-2n-1)}}{T_{n-1,m}^{(\mu-2m-2n-1)}},$ (4.12c)
are solutions of equation (4.10) and
$\mathcal{X}_{m,n}^{(\mu)}(z)=\frac{1}{w_{m,n}^{(\mu)}(z)-1}$ $=\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{m-1,n}^{(\mu+1)}(z)}{T_{m-1,n+1}^{(\mu-1)}(z)},$ (4.13a)
$\mathcal{Y}_{m,n}^{(\mu)}(z)=\frac{1}{v_{m,n}^{(\mu)}(z)-1}$ $=-1+\frac{\mu}{z}+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{m-1,n+1}^{(\mu-2n-1)}(z)}{T_{m,n}^{(\mu-2n-1)}(z)},$ (4.13b)
$\mathcal{Z}_{m,n}^{(\mu)}(z)=\frac{1}{u_{m,n}^{(\mu)}(z)-1}$ $=\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{n,m}^{(\mu-2m-2n-1)}}{T_{n-1,m}^{(\mu-2m-2n-1)}},$ (4.13c)
are solutions of equation (4.11).
Proof.
Since
$\frac{w+1}{w-1}=1+\frac{2}{w-1},$ (4.14)
then we prove (4.13a) first. By definition
${w}_{m,n}^{(\mu)}=\frac{T_{m-1,n}^{(\mu)}\,T_{m-1,n+1}^{(\mu)}}{T_{m,n}^{(\mu)}\,T_{m-2,n+1}^{(\mu)}},$
and so
$\frac{1}{w_{m,n}^{(\mu)}-1}=\frac{T_{m,n}^{(\mu)}\,T_{m-2,n+1}^{(\mu)}}{T_{m-1,n}^{(\mu)}\,T_{m-1,n+1}^{(\mu)}-T_{m,n}^{(\mu)}\,T_{m-2,n+1}^{(\mu)}}.$
Also
$\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{m-1,n}^{(\mu+1)}}{T_{m-1,n+1}^{(\mu-1)}}=\frac{{\rm D}_{z}\big(T_{m-1,n}^{(\mu+1)}{\,\scriptscriptstyle\bullet\,}T_{m-1,n+1}^{(\mu-1)}\big)}{T_{m-1,n}^{(\mu+1)}\,T_{m-1,n+1}^{(\mu-1)}}=\frac{T_{m,n}^{(\mu)}\,T_{m-2,n+1}^{(\mu)}}{T_{m-1,n}^{(\mu)}\,T_{m-1,n+1}^{(\mu)}-T_{m,n}^{(\mu)}\,T_{m-2,n+1}^{(\mu)}},$
since from the identity (A.1c)
${\rm D}_{z}\big(T_{m-1,n}^{(\mu+1)}{\,\scriptscriptstyle\bullet\,}T_{m-1,n+1}^{(\mu-1)}\big)=T_{m,n}^{(\mu)}\,T_{m-2,n+1}^{(\mu)},$
where ${\rm D}_{z}(f{\,\scriptscriptstyle\bullet\,}g)$ is the Hirota operator defined by
${\rm D}_{z}(f{\,\scriptscriptstyle\bullet\,}g)=\frac{\mathrm{d}{f}}{{\mathrm{d}{z}}}g-f\frac{\mathrm{d}{g}}{{\mathrm{d}{z}}}$}, (4.15)
and from the identity (A.2c)
$T_{m-1,n}^{(\mu+1)}\,T_{m-1,n+1}^{(\mu-1)}=\ T_{m-1,n}^{(\mu)}\,T_{m-1,n+1}^{(\mu)}-T_{m,n}^{(\mu)}\,T_{m-2,n+1}^{(\mu)},$
which completes the proof of (4.13a). The identity (4.12a) follows from equation (4.14).
By definition
${v}_{m,n}^{(\mu+2n+1)}=\frac{m+n+\mu+1}{m+n+1}\,\frac{T_{m,n}^{(\mu+1)}\,T_{m-1,n+1}^{(\mu-1)}}{T_{m-1,n}^{(\mu+1)}\,T_{m,n+1}^{(\mu-1)}},$
and so
$\frac{1}{{v}_{m,n}^{(\mu+2n+1)}-1}+1$ $=\frac{(m+n+\mu+1)T_{m,n}^{(\mu+1)}\,T_{m-1,n+1}^{(\mu-1)}}{(m+n+\mu+1)T_{m,n}^{(\mu+1)}\,T_{m-1,n+1}^{(\mu-1)}-(m+n+1)T_{m-1,n}^{(\mu+1)}\,T_{m,n+1}^{(\mu-1)}}.$
From (4.7a) and (4.9b) it follows
$z{\rm D}_{z}\big(T_{m-1,n}^{(\mu+1)}{\,\scriptscriptstyle\bullet\,}T_{m,n+1}^{(\mu-1)}\big)=(m+n+\mu+1)T_{m,n}^{(\mu+1)}\,T_{m-1,n+1}^{(\mu-1)}-(m+n+1)T_{m,n+1}^{(\mu-1)}\,T_{m-1,n}^{(\mu+1)},$
and since from (A.1b)
${\rm D}_{z}\big(T_{m-1,n}^{(\mu+1)}{\,\scriptscriptstyle\bullet\,}T_{m,n+1}^{(\mu-1)}\big)=T_{m,n}^{(\mu)}\,T_{m-1,n+1}^{(\mu)},$
then we have
$zT_{m,n}^{(\mu)}\,T_{m-1,n+1}^{(\mu)}=(m+n+\mu+1)T_{m,n}^{(\mu+1)}\,T_{m-1,n+1}^{(\mu-1)}-(m+n+1)T_{m,n+1}^{(\mu-1)}\,T_{m-1,n}^{(\mu+1)}.$
Therefore
$\frac{1}{{v}_{m,n}^{(\mu+2n+1)}-1}+1=\frac{m+n+\mu+1}{z}\,\frac{T_{m,n}^{(\mu+1)}\,T_{m-1,n+1}^{(\mu-1)}}{T_{m-1,n+1}^{(\mu)}\,T_{m,n}^{(\mu)}},$
and
$\frac{m+n+\mu+1}{z}+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{m-1,n+1}^{(\mu)}}{T_{m,n}^{(\mu)}}=\frac{m+n+\mu+1}{z}\,\frac{T_{m,n}^{(\mu+1)}\,T_{m-1,n+1}^{(\mu-1)}}{T_{m-1,n+1}^{(\mu)}\,T_{m,n}^{(\mu)}}.$
Hence we obtain (4.13b). The identity (4.12b) follows from equation (4.14).
By definition
$u_{m,n}^{(\mu+2m+2n+1)}(z)=\frac{T_{n-1,m}^{(\mu+1)}\,T_{n,m}^{(\mu-1)}}{T_{n,m-1}^{(\mu+1)}\,T_{n-1,m+1}^{(\mu-1)}},$
so
$\frac{1}{u_{m,n}^{(\mu+2m+2n+1)}(z)-1}=\frac{T_{n,m-1}^{(\mu+1)}\,T_{n-1,m+1}^{(\mu-1)}}{T_{n-1,m}^{(\mu+1)}\,T_{n,m}^{(\mu-1)}-T_{n,m-1}^{(\mu+1)}\,T_{n-1,m+1}^{(\mu-1)}}.$
Also
$\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{n,m}^{(\mu)}}{T_{n-1,m}^{(\mu)}}=\frac{{\rm D}_{z}\big(T_{n,m}^{(\mu)}{\,\scriptscriptstyle\bullet\,}T_{n-1,m}^{(\mu)}\big)}{T_{n,m}^{(\mu)}\,T_{n-1,m}^{(\mu)}}=\frac{T_{n,m-1}^{(\mu+1)}\,T_{n-1,m+1}^{(\mu-1)}}{T_{n-1,m}^{(\mu+1)}\,T_{n,m}^{(\mu-1)}-T_{n,m-1}^{(\mu+1)}\,T_{n-1,m+1}^{(\mu-1)}},$
since from the identity (A.1f)
${\rm D}_{z}\big(T_{n,m}^{(\mu)}{\,\scriptscriptstyle\bullet\,}T_{n-1,m}^{(\mu)}\big)=T_{n,m-1}^{(\mu+1)}\,T_{n-1,m+1}^{(\mu-1)},$
and from the identity (A.2b)
$T_{n,m}^{(\mu)}\,T_{n-1,m}^{(\mu)}=T_{n-1,m}^{(\mu+1)}\,T_{n,m}^{(\mu-1)}-T_{n,m-1}^{(\mu+1)}\,T_{n-1,m+1}^{(\mu-1)},$
which completes the proof of (4.13c). The identity (4.12c) follows from equation (4.14). ∎
4.2 Generalised Umemura polynomials
Rational solutions in cases (ii) and (iii) of Theorem 4.1 are expressed in terms of generalised Umemura polynomials. Umemura [70] defined some polynomials through a differential-difference equation to describe rational solutions of PV (1.2); see also [10, 50, 75]. Subsequently these were generalised by Masuda, Ohta and Kajiwara [45], who defined the generalised Umemura polynomial $R_{m,n}^{(\kappa)}(z)$ through coupled differential-difference equations and gave a representation as the determinant
$R_{m,n}^{(\ell)}(z)=\left|\begin{matrix}q_{1}^{(\ell)}&\cdots&q_{-m+2}^{(\ell)}&q_{-m+1}^{(\ell)}&\cdots&q_{-m-n+2}^{(\ell)}\\[2.0pt] q_{3}^{(\ell)}&\cdots&q_{-m+4}^{(\ell)}&q_{-m+3}^{(\ell)}&\cdots&q_{-m-n+4}^{(\ell)}\\[2.0pt] \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\[2.0pt] q_{2m-1}^{(\ell)}&\cdots&q_{m}^{(\ell)}&q_{m-1}^{(\ell)}&\cdots&q_{m-n}^{(\ell)}\\[2.0pt] p_{n-m}^{(\ell)}&\cdots&p_{n+1}^{(\ell)}&p_{n}^{(\ell)}&\cdots&p_{2n-1}^{(\ell)}\\[2.0pt] p_{n-m-2}^{(\ell)}&\cdots&p_{n-1}^{(\ell)}&p_{n-2}^{(\ell)}&\cdots&p_{2n-3}^{(\ell)}\\[2.0pt] \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\[2.0pt] p_{-n-m+2}^{(\ell)}&\cdots&p_{-n+1}^{(\ell)}&p_{-n+2}^{(\ell)}&\cdots&p_{1}^{(\ell)}\end{matrix}\right|,$ (4.16)
where $p_{k}^{(\ell)}(z)=L_{k}^{(\ell-1)}(\tfrac{1}{2}z)$ and $q_{k}^{(\ell)}(z)=L_{k}^{(\ell-1)}(-\tfrac{1}{2}z)$, with $p_{k}^{(\ell)}=q_{k}^{(\ell)}=0$ for $k<0$, and $L_{k}^{(\alpha)}(z)$ is the Laguerre polynomial (4.4). Recently Masuda [44] studied properties of zeros of the polynomial $R_{m,n}^{(\ell)}(z)$.
Rather than use (4.16) for the generalised Umemura polynomial, we prefer to use an equivalent representation given in Definition 4.7, which we find simpler to compute and useful to prove algebraic and combinatorial properties of the polynomial, though we do not pursue this further here.
Definition 4.7.
The generalised Umemura polynomial $U_{m,n}^{(\kappa)}(z)$ is given by the Wronskian
$U_{m,n}^{(\kappa)}(z)=c_{m,n}\exp(\tfrac{1}{2}mz)\operatorname{Wr}\!\left({\varphi^{(\kappa)}_{1}},{\varphi^{(\kappa)}_{3}},\ldots,{\varphi^{(\kappa)}_{2m-1}};{\psi^{(\kappa)}_{1}},{\psi^{(\kappa)}_{3}},\ldots,{\psi^{(\kappa)}_{2n-1}}\right),\quad m,n\geq 0,$ (4.17)
where
$c_{m,n}=(-1)^{m}\,2^{(m+n)(m+n+1)/2}\prod_{j=1}^{m-1}\frac{(2j+1)!}{j!}\prod_{k=1}^{n-1}\frac{(2k+1)!}{k!},$ (4.18)
$\varphi^{(\kappa)}_{n}(z)=\mathrm{e}^{-z/2}\,L_{n}^{(\kappa)}(\tfrac{1}{2}z)$} and $\psi^{(\kappa)}_{n}(z)=L_{n}^{(\kappa)}(-\tfrac{1}{2}z)$}, with $\kappa$ a parameter, and $L_{n}^{(\kappa)}(z)$ is the Laguerre polynomial (4.4) and $U_{0,0}^{(\kappa)}(z)=1$.
Remarks 4.8.
  1. 1.
    The generalised Laguerre polynomial $T_{m,n}^{(\mu)}(z)$ is a Wronskian involving one sequence of Laguerre polynomials, whereas the generalised Umemura polynomial $U_{m,n}^{(\kappa)}(z)$ involves two sequences of Laguerre polynomials, unless either $m=0$ or $n=0$. This is analogous to the situation for rational solutions of PIV (1.10), where one set of rational solutions involves one sequence of Hermite polynomials and a second set involves two sequences of Hermite polynomials, cf. [11, 52].
  2. 2.
    The generalised Umemura polynomial $U_{m,n}^{(\kappa)}(z)$ given by (4.17) is a monic polynomial of degree $\tfrac{1}{2}m(m+1)+\tfrac{1}{2}n(n+1)$ and has the symmetry
    $U_{m,n}^{(\kappa)}(z)=U_{n,m}^{(-\kappa-2m-2n)}(z).$ (4.19)
  3. 3.
    The polynomials $R_{m,n}^{(\ell)}(z)$ and $U_{m,n}^{(\kappa)}(z)$ given by (4.16) and (4.17) respectively, are equivalent, which can be shown using matrix row and column operations. Further the two polynomials are related as follows
    $U_{m,n}^{(\kappa)}(z)=c_{m,n}(-1)^{m+n(n+1)/2}\left(\tfrac{1}{2}\right)^{(m+n)(m+n-1)/2}R_{m,n}^{(-\kappa-m-n)}(z),$
    with $c_{m,n}$ given by (4.18).
  4. 4.
    When $m=0$, the polynomial $U_{0,n}^{(\kappa)}(z)$ is equivalent to the polynomial defined by Umemura [70] through a differential-difference equation. Also $U_{0,n}^{(\kappa)}(z)$ is equivalent to the determinant $\tau_{m}^{(n)}(z)$ defined in Theorem 5.1; see also [50]. Due to the symmetry (4.19), the same is true for the polynomial $U_{m,0}^{(\kappa)}(z)$.
Theorem 4.9.
Suppose that $U_{m,n}^{(\kappa)}(z)$ is the generalised Umemura polynomial given by (4.17).
  1. (a)
    In case (ii) of Theorem 4.1, PV (1.2) has the rational solutions for $m\geq 1$ and $n\geq 1$
    $\widehat{w}_{m,n}^{(\kappa)}(z)$ $=-\frac{U_{m,n-1}^{(2\kappa)}(z)\,U_{m-1,n}^{(2\kappa+2)}(z)}{U_{m,n-1}^{(2\kappa+2)}(z)\,U_{m-1,n}^{(2\kappa)}(z)},$ $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(m+\kappa)^{2},-\tfrac{1}{2}(n+\kappa)^{2},m+n\big),$ (4.20a)
    $\widehat{v}_{m,n}^{(\kappa)}(z)$ $=-\frac{U_{m-1,n-1}^{(2\kappa-2n+3)}(z)\,U_{m,n}^{(2\kappa-2n-1)}(z)}{U_{m-1,n-1}^{(2\kappa-2n+1)}(z)\,U_{m,n}^{(2\kappa-2n+1)}(z)},$ $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(m+\kappa)^{2},-\tfrac{1}{2}(n-\kappa)^{2},m-n\big),$ (4.20b)
    In the case when $n=0$, then for $m\geq 1$
    $\widehat{w}_{m,0}^{(\kappa)}(z)=\widehat{v}_{m,0}^{(\kappa)}(z)=-\frac{U_{m,0}^{(2\kappa-1)}(z)\,U_{m-1,0}^{(2\kappa+2)}(z)}{U_{m,0}^{(2\kappa+1)}(z)\,U_{m-1,0}^{(2\kappa)}(z)},\qquad(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(m+\kappa)^{2},-\tfrac{1}{2}\kappa^{2},m\big),$ (4.20c)
    and in the case when $m=0$ then for $n\geq 1$
    $\widehat{w}_{0,n}^{(\kappa)}(z)=\widehat{v}_{0,n}^{(-\kappa)}(z)=-\frac{U_{0,n}^{(2\kappa+1)}(z)\,U_{0,n-1}^{(2\kappa)}(z)}{U_{0,n}^{(2\kappa-1)}(z)\,U_{0,n-1}^{(2\kappa+2)}(z)},\qquad(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}\kappa^{2},-\tfrac{1}{2}(n+\kappa)^{2},n\big).$ (4.20d)
  2. (b)
    In case (iii) of Theorem 4.1, PV (1.2) has the rational solution for $m\geq 1$ and $n\geq 1$
    $\widehat{u}_{m,n}^{(\kappa)}(z)=-\,\frac{U_{m,n-1}^{(\kappa+1)}(z)\,U_{m,n+1}^{(\kappa-1)}(z)}{U_{m-1,n}^{(\kappa+1)}(z)\,U_{m+1,n}^{(\kappa-1)}(z)},\qquad(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(m+\tfrac{1}{2})^{2},-\tfrac{1}{2}(n+\tfrac{1}{2})^{2},m+n+\kappa\big).$ (4.21a)
    In the case when $n=0$, then for $m\geq 1$
    $\widehat{u}_{m,0}^{(\kappa)}(z)=-\,\frac{U_{m,0}^{(\kappa)}(z)\,U_{m,1}^{(\kappa-1)}(z)}{U_{m-1,0}^{(\kappa+1)}(z)\,U_{m+1,0}^{(\kappa-1)}(z)},\qquad(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(m+\tfrac{1}{2})^{2},-\tfrac{1}{8},m+\kappa\big),$ (4.21b)
    and in the case when $m=0$ then for $n\geq 1$
    $\widehat{u}_{0,n}^{(\kappa)}(z)=-\,\frac{U_{0,n-1}^{(\kappa+1)}(z)\,U_{0,n+1}^{(\kappa-1)}(z)}{U_{0,n}^{(\kappa)}(z)\,U_{1,n}^{(\kappa-1)}(z)},\qquad(\alpha,\beta,\gamma)=\big(\tfrac{1}{8},-\tfrac{1}{2}(n+\tfrac{1}{2})^{2},n+\kappa\big).$ (4.21c)
Proof.
See Masuda, Ohta and Kajiwara [45, Theorem 1.1].∎
Remark 4.10.
The rational solutions in Theorem 4.9 can also be written as logarithmic derivatives
$\widehat{w}_{m,n}^{(\kappa)}(z)$ $=\frac{z}{m+\kappa}\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{m-1,n}^{(2\kappa)}(z)}{U_{m,n-1}^{(2\kappa+2)}(z)}-\frac{n+\kappa}{m+\kappa},$ (4.22a)
$\widehat{v}_{m,n}^{(\kappa+n)}(z)$ $=\frac{z}{m+n+\kappa}\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{m-1,n-1}^{(2\kappa+1)}(z)}{U_{m,n}^{(2\kappa+1)}(z)}-\frac{\kappa}{m+n+\kappa},$ (4.22b)
$\widehat{u}_{m,n}^{(\kappa)}(z)$ $=\frac{2z}{2m+1}\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{m-1,n}^{(\kappa+1)}(z)}{U_{m+1,n}^{(\kappa-1)}(z)}+1.$ (4.22c)
In the following Lemma we derive the solutions of equations (4.10) and (4.11) in terms of generalised Umemura polynomials.
Lemma 4.11.
If $\widehat{w}_{m,n}^{(\kappa)}(z)$, $\widehat{v}_{m,n}^{(\kappa)}(z)$ and $\widehat{u}_{m,n}^{(\kappa)}(z)$ are rational solutions of PV defined in Theorem 4.9 and $U_{m,n}^{(\mu)}(z)$ is the generalised Umemura polynomial given by (4.17), then
$\widehat{\mathcal{W}}_{m,n}^{(\kappa)}(z)=\frac{\widehat{w}_{m,n}^{(\kappa)}(z)+1}{\widehat{w}_{m,n}^{(\kappa)}(z)-1}$ $=2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{m,n}^{(2\kappa)}(z)}{U_{m-1,n-1}^{(2\kappa+2)}(z)},$ (4.23a)
$\widehat{\mathcal{V}}_{m,n}^{(\kappa)}(z)=\frac{\widehat{v}_{m,n}^{(\kappa)}(z)+1}{\widehat{v}_{m,n}^{(\kappa)}(z)-1}$ $=2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{m,n-1}^{(2\kappa-2n+1)}(z)}{U_{m-1,n}^{(2\kappa-2n+1)}(z)},$ (4.23b)
$\widehat{\mathcal{U}}_{m,n}^{(\kappa)}(z)=\frac{\widehat{u}_{m,n}^{(\kappa)}(z)+1}{\widehat{u}_{m,n}^{(\kappa)}(z)-1}$ $=\frac{2(m+n+\kappa)}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{m,n}^{(\kappa+1)}(z)}{U_{m,n}^{(\kappa-1)}(z)},$ (4.23c)
are solutions of (4.10) and
$\widehat{\mathcal{X}}_{m,n}^{(\kappa)}(z)=\frac{1}{\widehat{w}_{m,n}^{(\kappa)}(z)-1}$ $=-\frac{1}{2}+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{m,n}^{(2\kappa)}(z)}{U_{m-1,n-1}^{(2\kappa+2)}(z)},$ (4.24a)
$\widehat{\mathcal{Y}}_{m,n}^{(\kappa)}(z)=\frac{1}{\widehat{v}_{m,n}^{(\kappa)}(z)-1}$ $=-\frac{1}{2}+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{m,n-1}^{(2\kappa-2n+1)}(z)}{U_{m-1,n}^{(2\kappa-2n+1)}(z)},$ (4.24b)
$\widehat{\mathcal{Z}}_{m,n}^{(\kappa)}(z)=\frac{1}{\widehat{u}_{m,n}^{(\kappa)}(z)-1}$ $=-\frac{1}{2}+\frac{(m+n+\kappa)}{z}+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{m,n}^{(\kappa+1)}(z)}{U_{m,n}^{(\kappa-1)}(z)}$ (4.24c)
are solutions of (4.11).
Proof.
We note that (4.24) follow from (4.23) using (4.14). By definition
$\widehat{w}_{m,n}^{(\kappa)}(z)=-\frac{U_{m,n-1}^{(2\kappa)}(z)\,U_{m-1,n}^{(2\kappa+2)}(z)}{U_{m,n-1}^{(2\kappa+2)}(z)\,U_{m-1,n}^{(2\kappa)}(z)},$
so that we have
$\frac{\widehat{w}_{m,n}^{(\kappa)}(z)+1}{\widehat{w}_{m,n}^{(\kappa)}(z)-1}=\frac{U_{m,n-1}^{(2\kappa)}(z)\,U_{m-1,n}^{(2\kappa+2)}(z)-U_{m,n-1}^{(2\kappa+2)}(z)\,U_{m-1,n}^{(2\kappa)}(z)}{U_{m,n-1}^{(2\kappa)}(z)\,U_{m-1,n}^{(2\kappa+2)}(z)+U_{m,n-1}^{(2\kappa+2)}(z)\,U_{m-1,n}^{(2\kappa)}(z)},$
$\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{m,n}^{(2\kappa)}(z)}{U_{m-1,n-1}^{(2\kappa+2)}(z)}=\frac{{\rm D}_{z}\big(U_{m,n}^{(2\kappa)}(z){\,\scriptscriptstyle\bullet\,}U_{m-1,n-1}^{(2\kappa+2)}(z)\big)}{U_{m,n}^{(2\kappa)}(z)\,U_{m-1,n-1}^{(2\kappa+2)}(z)},$
where ${\rm D}_{z}(f{\,\scriptscriptstyle\bullet\,}g)$ is the Hirota operator (4.15). From the identities (A.9b) and (A.10b)
$4{\rm D}_{z}\big(U_{m,n}^{(2\kappa)}(z){\,\scriptscriptstyle\bullet\,}U_{m-1,n-1}^{(2\kappa+2)}(z)\big)$ $=U_{m,n-1}^{(2\kappa)}(z)\,U_{m-1,n}^{(2\kappa+2)}(z)-U_{m,n-1}^{(2\kappa+2)}(z)\,U_{m-1,n}^{(2\kappa)}(z),$
$2U_{m,n}^{(2\kappa)}(z)\,U_{m-1,n-1}^{(2\kappa+2)}(z)$ $=U_{m,n-1}^{(2\kappa)}(z)\,U_{m-1,n}^{(2\kappa+2)}(z)+U_{m,n-1}^{(2\kappa+2)}(z)\,U_{m-1,n}^{(2\kappa)}(z),$
respectively, therefore
$\frac{2{\rm D}_{z}\big(U_{m,n}^{(2\kappa)}(z){\,\scriptscriptstyle\bullet\,}U_{m-1,n-1}^{(2\kappa+2)}(z)\big)}{U_{m,n}^{(2\kappa)}(z)\,U_{m-1,n-1}^{(2\kappa+2)}(z)}=\frac{U_{m,n-1}^{(2\kappa)}(z)\,U_{m-1,n}^{(2\kappa+2)}(z)-U_{m,n-1}^{(2\kappa+2)}(z)\,U_{m-1,n}^{(2\kappa)}(z)}{U_{m,n-1}^{(2\kappa)}(z)\,U_{m-1,n}^{(2\kappa+2)}(z)+U_{m,n-1}^{(2\kappa+2)}(z)\,U_{m-1,n}^{(2\kappa)}(z)},$
which proves (4.23a).
By definition
$\widehat{v}_{m,n}^{(\kappa+n-1/2)}(z)=-\frac{U_{m-1,n-1}^{(2\kappa+2)}(z)\,U_{m,n}^{(2\kappa-2)}(z)}{U_{m-1,n-1}^{(2\kappa)}(z)\,U_{m,n}^{(2\kappa)}(z)},$
so that we have
$\frac{\widehat{v}_{m,n}^{(\kappa+n-1/2)}(z)+1}{\widehat{v}_{m,n}^{(\kappa+n-1/2)}(z)-1}=\frac{U_{m,n}^{(2\kappa-2)}\,U_{m-1,n-1}^{(2\kappa+2)}-U_{m,n}^{(2\kappa)}\,U_{m-1,n-1}^{(2\kappa)}}{U_{m,n}^{(2\kappa-2)}\,U_{m-1,n-1}^{(2\kappa+2)}+U_{m,n}^{(2\kappa)}\,U_{m-1,n-1}^{(2\kappa)}},$
$\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{m,n-1}^{(2\kappa)}}{U_{m-1,n}^{(2\kappa)}}=\frac{{\rm D}_{z}\big(U_{m,n-1}^{(2\kappa)}{\,\scriptscriptstyle\bullet\,}U_{m-1,n}^{(2\kappa)}\big)}{U_{m,n-1}^{(2\kappa)}\,U_{m-1,n}^{(2\kappa)}}.$
From the identities (A.9a) and (A.10a)
$4{\rm D}_{z}\big(U_{m,n-1}^{(2\kappa)}{\,\scriptscriptstyle\bullet\,}U_{m-1,n}^{(2\kappa)}\big)$ $=U_{m,n}^{(2\kappa-2)}\,U_{m-1,n-1}^{(2\kappa+2)}-U_{m,n}^{(2\kappa)}\,U_{m-1,n-1}^{(2\kappa)},$
$2U_{m,n-1}^{(2\kappa)}\,U_{m-1,n}^{(2\kappa)}$ $=U_{m,n}^{(2\kappa-2)}\,U_{m-1,n-1}^{(2\kappa+2)}+U_{m,n}^{(2\kappa)}\,U_{m-1,n-1}^{(2\kappa)},$
respectively, therefore
$\frac{2{\rm D}_{z}\big(U_{m,n-1}^{(2\kappa)}{\,\scriptscriptstyle\bullet\,}U_{m-1,n}^{(2\kappa)}\big)}{U_{m,n-1}^{(2\kappa)}\,U_{m-1,n}^{(2\kappa)}}=\frac{U_{m,n}^{(2\kappa-2)}\,U_{m-1,n-1}^{(2\kappa+2)}-U_{m,n}^{(2\kappa)}\,U_{m-1,n-1}^{(2\kappa)}}{U_{m,n}^{(2\kappa-2)}\,U_{m-1,n-1}^{(2\kappa+2)}+U_{m,n}^{(2\kappa)}\,U_{m-1,n-1}^{(2\kappa)}},$
which proves (4.23b).
By definition
$\widehat{u}_{m,n}^{(\kappa)}(z)=-\,\frac{U_{m,n-1}^{(\kappa+1)}(z)\,U_{m,n+1}^{(\kappa-1)}(z)}{U_{m-1,n}^{(\kappa+1)}(z)\,U_{m+1,n}^{(\kappa-1)}(z)},$
and so
$\frac{\widehat{u}_{m,n}^{(\kappa)}+1}{\widehat{u}_{m,n}^{(\kappa)}-1}=\frac{U_{m,n-1}^{(\kappa+1)}\,U_{m,n+1}^{(\kappa-1)}-U_{m-1,n}^{(\kappa+1)}\,U_{m+1,n}^{(\kappa-1)}}{U_{m,n-1}^{(\kappa+1)}\,U_{m,n+1}^{(\kappa-1)}+U_{m-1,n}^{(\kappa+1)}\,U_{m+1,n}^{(\kappa-1)}}.$
Also
$\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{m,n}^{(\kappa+1)}}{U_{m,n}^{(\kappa-1)}}+\frac{m+n+\kappa}{z}$ $=\frac{{\rm D}_{z}\big(U_{m,n}^{(\kappa+1)}{\,\scriptscriptstyle\bullet\,}U_{m,n}^{(\kappa-1)}\big)}{U_{m,n}^{(\kappa+1)}\,U_{m,n}^{(\kappa-1)}}+\frac{m+n+\kappa}{z}$
$=\frac{z{\rm D}_{z}\big(U_{m,n}^{(\kappa+1)}{\,\scriptscriptstyle\bullet\,}U_{m,n}^{(\kappa-1)}\big)+(m+n+\kappa)U_{m,n}^{(\kappa+1)}\,U_{m,n}^{(\kappa-1)}}{zU_{m,n}^{(\kappa+1)}\,U_{m,n}^{(\kappa-1)}}$
It can be shown that
$U_{m,n-1}^{(\kappa+1)}\,U_{m,n+1}^{(\kappa-1)}+U_{m-1,n}^{(\kappa+1)}\,U_{m+1,n}^{(\kappa-1)}$ $=2zU_{m,n}^{(\kappa+1)}\,U_{m,n}^{(\kappa-1)},$ (4.25)
$U_{m,n-1}^{(\kappa+1)}\,U_{m,n+1}^{(\kappa-1)}-U_{m-1,n}^{(\kappa+1)}\,U_{m+1,n}^{(\kappa-1)}$ $=4z{\rm D}_{z}\big(U_{m,n}^{(\kappa+1)}{\,\scriptscriptstyle\bullet\,}U_{m,n}^{(\kappa-1)}\big)+4(m+n+\kappa)U_{m,n}^{(\kappa+1)}\,U_{m,n}^{(\kappa-1)},$ (4.26)
which proves (4.23c). Equation (4.25) is the identity (A.13) and using computer algebra, we have verified equation (4.26) for several small values of $m$ and $n$. ∎
5 Rational solutions of the discrete equations
In this section we discuss rational solutions of the discrete equations derived in §3. It is known that there are rational solutions of the dPII equation (1.6), see [36, 37, 39, 50]. This is illustrated in the following theorem.
Theorem 5.1.
Define $\tau_{m}^{(n)}(z)$ to be the $m\times m$ determinant given by
[…]
where $p^{(n)}_{k}(z)=L^{(n)}_{k}(\tfrac{1}{2}z)$, with $L^{(n)}_{k}(z)$ the Laguerre polynomial, for $k\geq 0$ and $p^{(n)}_{k}(z)=0$ for $k<0$, then
$q_{n}(z;m)=\frac{\tau_{m+1}^{(n+1)}(z)\,\tau_{m}^{(n-1)}(z)}{\tau_{m+1}^{(n)}(z)\,\tau_{m}^{(n)}(z)}-1,$ (5.2)
satisfies
$q_{n+1}+q_{n-1}=\frac{4}{z}\frac{(n+1)q_{n}-m-1}{1-q_{n}^{2}},$ (5.3)
which is dPII (1.6) with $\zeta=4/z$, $\lambda=1$ and $\rho=-m-1$.
Proof.
See Kajiwara, Yamamoto and Ohta [39].∎
Remarks 5.2.
  1. 1.
    The determinant $\tau_{m}^{(n)}(z)$ can be written as the Wronskian
    $\tau_{m}^{(n)}(z)=(-1)^{m(m-1)/2}\operatorname{Wr}\big(p^{(n-m+1)}_{1},p^{(n-m+1)}_{3},\ldots,p^{(n-m+1)}_{2m-1}\big)$, (5.4)
    with $p^{(n)}_{k}(z)=L^{(n)}_{k}(\tfrac{1}{2}z)$, and is related to the generalised Umemura polynomial by
    $U_{m,0}^{(n-m+1)}(z)=U_{0,m}^{(-n-m-1)}(z)=c_{0,m}(-1)^{m(m+1)/2}\left(\tfrac{1}{2}\right)^{m(m-1)/2}\tau_{m}^{(n)}(z).$
  2. 2.
    The function $q_{n}(z;m)$ satisfies equation (4.10) with
    $(\alpha,\beta,\gamma)=\big(\tfrac{1}{8}(m-n)^{2},-\tfrac{1}{8}(m+n+2)^{2},-m-1\big).$
5.1 Rational solutions in terms of generalised Laguerre polynomials
Each of the transformations $\mathcal{R}_{1}$ (3.3), $\mathcal{R}_{2}$ (3.12) and $\mathcal{R}_{3}$ (3.23) is actually four transformations since there are two independent choices of sign for $a=\pm\sqrt{2\alpha}$ and $b=\pm\sqrt{2\beta}$. This can lead to four different solutions from applying the transformation to a given solution. This is illustrated in Table 5.1 which shows the effect on the parameters of applying the transformations $\mathcal{R}_{1}$ (3.3) and $\mathcal{R}_{1}^{2}$ on a solution $w(z)$ of PV (1.2) with $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}a^{2},-\tfrac{1}{2}b^{2},c\big)$; see also Example 5.3.
[…]
Table 5.1: The effect on the parameters of applying the transformations $\mathcal{R}_{1}$ (3.3) and $\mathcal{R}_{1}^{2}$ on a solution $w(z)$ of PV with $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}a^{2},-\tfrac{1}{2}b^{2},c\big)$.
5.1.1 First discrete equation: asymmetric discrete PII
For the Bäcklund transformation $\mathcal{R}_{1}=\mathcal{T}_{1,-1,1}$, we know that $\mathcal{R}_{1}^{2}(a,b,c)=(a+1,b-1,c)$, so that
$\mathcal{R}_{1}^{2}\big(w_{m,n}^{(\mu)};m,m+2n+1+\mu,\mu\big)$ $=\big(w_{m+1,n-1}^{(\mu)};m+1,m+2n+\mu,\mu\big),$ (5.5a)
$\mathcal{R}_{1}^{2}\big(w_{m,n}^{(\mu)};-m,m+2n+1+\mu,\mu\big)$ $=\big(w_{m-1,n}^{(\mu)};1-m,m+2n+\mu,\mu\big),$ (5.5b)
$\mathcal{R}_{1}^{2}\big(w_{m,n}^{(\mu)};m,-m-2n-1-\mu,\mu\big)$ $=\big(w_{m+1,n}^{(\mu)};m+1,-m-2n-2-\mu,\mu\big),$ (5.5c)
$\mathcal{R}_{1}^{2}\big(w_{m,n}^{(\mu)};-m,-m-2n-1-\mu,\mu\big)$ $=\big(w_{m-1,n+1}^{(\mu)};1-m,-m-2n-2-\mu,\mu\big).$ (5.5d)
From (5.5) we see that except in the case (5.5c), the Bäcklund transformation $\mathcal{R}_{1}$ (3.3) can only be applied a finite number of times before the sequence terminates as either $m$ or $n$ becomes zero, as illustrated in the following example.
Example 5.3.
Consider the rational solution
$w_{1,1}^{(\mu-3)}(z)$ $=-\frac{(z-\mu+1)\big[z^{2}-2\mu z+\mu(\mu+1)\big]}{z^{2}-2\mu z+\mu(\mu-1)},$
which satisfies PV (1.2) with $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2},-\tfrac{1}{2}(\mu+1)^{2},\mu-3\big)$, so that
$q_{0}(z)=\frac{w_{1,1}^{(\mu-3/2)}(z)+1}{w_{1,1}^{(\mu-3/2)}(z)-1}=1+\frac{2}{z-\mu}-\frac{4(z-\mu+1)}{z^{2}-2\mu z+\mu(\mu-1)}.$
Applying the Bäcklund transformation $\mathcal{R}_{1}$ (3.3), with $(a,b,c)=(1,\mu+1,\mu-3)$, gives the sequence of rational solutions
$q_{1}^{[a]}(z)$ $=1-\frac{2}{z-\mu+1}+\frac{4(z-\mu)}{z^{2}-2\mu z+\mu(\mu-1)},$
$q_{2}^{[a]}(z)$ $=1-\frac{4(z-\mu+1)}{z^{2}-2(\mu-1)z+(\mu-1)(\mu-2)},\qquad q_{3}^{[a]}(z)=1.$
Applying $\mathcal{R}_{1}$ (3.3), with $(a,b,c)=(-1,\mu+1,\mu-3)$, gives the sequence of rational solutions
$q_{1}^{[b]}(z)=-1+\frac{2\mu}{z}-\frac{2}{z-\mu+1},\qquad q_{2}^{[b]}(z)=1,$
and applying $\mathcal{R}_{1}$ (3.3), with $(a,b,c)=(-1,-\mu-1,\mu-3)$, gives the sequence of rational solutions
$q_{1}^{[d]}(z)=-1-\frac{4(z-\mu)}{z^{2}-2\mu z+\mu(\mu+1)},\qquad q_{2}^{[d]}(z)=1.$
In each case the Bäcklund transformation only produces a finite number of solutions.
Applying $\mathcal{R}_{1}$ (3.3), with $(a,b,c)=(1,-\mu-1,\mu-3)$, gives an infinite sequence of solutions. The next few are
$q_{1}(z)$ $=1-\frac{2\mu}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{2}-2\mu z+\mu(\mu-1)}{z^{2}-2\mu z+\mu(\mu+1)},$
$q_{2}(z)$ $=1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{2}-2(\mu+1)z+\mu(\mu+1)}{z^{4}-4\mu z^{3}+6\mu^{2}z^{2}-4\mu(\mu^{2}-1)z+\mu^{2}(\mu^{2}-1)},$
$q_{3}(z)$ $=1-\frac{2\mu}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{3}-3(\mu+1)z^{2}+3\mu(\mu+1)z-\mu(\mu^{2}-1)}{z^{4}-4(\mu+1)z^{3}+6(\mu+1)^{2}z^{2}-4\mu(\mu+1)(\mu+2)z+\mu(\mu+1)^{2}(\mu+2)}.$
Generally for $n\geq 0$
$q_{2n}(z)=1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{n,1}^{(\mu-2)}(z)}{T_{n,2}^{(\mu-4)}(z)},\qquad q_{2n+1}(z)=1-\frac{2\mu}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{n+1,1}^{(\mu-3)}(z)}{T_{n,2}^{(\mu-3)}(z)},$
satisfy
$q_{n+1}+q_{n-1}=\frac{4}{z}\,\frac{(n+\mu+2)q_{n}-\tfrac{3}{2}-(\mu-\tfrac{3}{2})(-1)^{n}}{1-q_{n}^{2}},$
which is asymmetric dPII (3.7) with $\lambda=\mu+2$, $\rho=-\tfrac{3}{2}$ and $\varphi=-\mu+\tfrac{3}{2}$, whilst $x_{n}=q_{2n}$ and $y_{n}=q_{2n+1}$ satisfy
$x_{n+1}+x_{n}$ $=\frac{4}{z}\,\frac{(2n+\mu+3)y_{n}+\mu-3}{1-y_{n}^{2}},\qquad y_{n}+y_{n-1}=\frac{4}{z}\,\frac{(2n+\mu+2)x_{n}-\mu}{1-x_{n}^{2}}.$
Also, applying the inverse transformation $\mathcal{R}_{1}^{-1}=\mathcal{T}_{-1,1,-1}$ we obtain
$q_{-1}(z)=1-\frac{2\mu}{z}+\frac{2}{z-\mu+1},\qquad q_{-2}(z)=1.$
We note that
$\mathcal{Q}_{1}^{3}\left(q_{1,1}^{(\mu)};1,\mu+1,\mu-3\right)=1,\qquad\mathcal{Q}_{1}^{2}\left(q_{1,1}^{(\mu)};-1,\mu+1,\mu-3\right)=1,\qquad\mathcal{Q}_{1}^{2}\left(q_{1,1}^{(\mu)};-1,-\mu-1,\mu-3\right)=1.$
where $\mathcal{Q}_{1}$ is the Bäcklund transformation
$\mathcal{Q}_{1}(q;a,b,c)=\left(\frac{2}{z(q^{2}-1)}\left\{{z}\frac{\mathrm{d}{q}}{{\mathrm{d}{z}}}-(a-b)q-a-b\right\};\tfrac{1}{2}(a-b+c+1),\tfrac{1}{2}(-a+b+c-1),a+b\right).$
Further if $\mathcal{R}_{1}$ (3.3) is the Bäcklund transformation given by (5.5c) then
$w_{1,1}^{(\mu-3)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\frac{1}{{v}_{1,1}^{(\mu)}(z)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip w_{2,1}^{(\mu-3)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\frac{1}{{v}_{2,1}^{(\mu)}(z)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip w_{3,1}^{(\mu-3)}(z)$
More generally, using the solution $w_{m,n}^{(\mu)}(z)$, for fixed $m$ and $n$, then we have the following Lemma.
Lemma 5.4.
Consider the rational functions
$Q_{2N}(z)=X_{N}(z)=\frac{w_{N+m,n}^{(\mu)}(z)+1}{w_{N+m,n}^{(\mu)}(z)-1},\qquad Q_{2N+1}(z)=Y_{N}(z)=\frac{1+v_{N+m,n}^{(\mu+2n+1)}(z)}{1-v_{N+m,n}^{(\mu+2n+1)}(z)},$
with $w_{m,n}^{(\mu)}(z)$ and $v_{m,n}^{(\mu)}(z)$ as given in Theorem 4.4, then $Q_{N}(z)$ satisfies
$Q_{N+1}+Q_{N-1}=\frac{4}{z}\frac{(N+2m+2n+\mu+1)Q_{N}-n-\tfrac{1}{2}-(n+\mu+\tfrac{1}{2})(-1)^{N}}{1-Q_{N}^{2}},$
which is asymmetric dPII (3.7) with $\lambda=2m+2n+\mu+1$, $\rho=-n-\tfrac{1}{2}$ and $\varphi=-n-\mu-\tfrac{1}{2}$, whilst $X_{N}(z)$ and $Y_{N}(z)$ satisfy the discrete system
$X_{N+1}+X_{N}$ $=\frac{4}{z}\frac{(2N+2m+2n+\mu+2)Y_{N}+\mu}{1-Y_{N}^{2}},$ (5.6a)
$Y_{N}+Y_{N-1}$ $=\frac{4}{z}\frac{(2N+2m+2n+\mu+1)X_{N}-2n-\mu-1}{1-X_{N}^{2}}.$ (5.6b)
Proof.
From the transformations
$\mathcal{R}_{1}\big(w_{m,n}^{(\mu)};m,-m-2n-1-\mu,\mu\big)$ $=\big({1/{v}_{m,n}^{(\mu+2n+1)}};m+n+1+\mu,-m-n-1,-\mu-2n-1\big),$
$\mathcal{R}_{1}\big(v_{m,n}^{(\mu)};m+n+1,-m+n-\mu,\mu\big)$ $=\big({1/{w}_{m+1,n}^{(\mu-2n-1)}};m+\mu+1,-m-1,-\mu+2n+1\big),$
we obtain the sequence of solutions
$w_{m,n}^{(\mu)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\frac{1}{{v}_{m,n}^{(\mu+2n+1)}}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip w_{m+1,n}^{(\mu)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\frac{1}{{v}_{m+1,n}^{(\mu+2n+1)}}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip w_{m+2,n}^{(\mu)}$
from which the result follows. We note that
$\mathcal{R}_{1}^{2}\Big(w_{m,n}^{(\mu)}\Big)=w_{m+1,n}^{(\mu)},\qquad\mathcal{R}_{1}^{2}\Big(1/{v}_{m,n}^{(\mu)}\Big)=1/{v}_{m+1,n}^{(\mu)}.$
Remarks 5.5.
  1. 1.
    An equivalent result can be obtained using the inverse transformation $\mathcal{R}_{1}^{-1}=\mathcal{T}_{-1,1,-1}$ since
    $\mathcal{R}_{1}^{-1}\big(w_{m,n}^{(\mu)};-m,m+2n+1+\mu,\mu\big)$ $=\big(v_{m,n}^{(\mu+2n+1)};-m-n-1,m+n+1+\mu,2n+1+\mu\big),$
    $\mathcal{R}_{1}^{-1}\big(v_{m,n}^{(\mu)};-m-n-1,m-n+\mu,\mu\big)$ $=\big(w_{m+1,n}^{(\mu-2n-1)};-m-1,m+1+\mu,-2n-1+\mu\big).$
  2. 2.
    We note that if $\mu=-n-\tfrac{1}{2}$ then
    $Q_{2N}(z)=\frac{w_{N+m,n}^{(-n-1/2)}(z)+1}{w_{N+m,n}^{(-n-1/2)}(z)-1},\qquad Q_{2N+1}(z)=\frac{v_{N+m,n}^{(n+1/2)}(z)+1}{v_{N+m,n}^{(n+1/2)}(z-1},$
    are solutions to
    $Q_{N+1}+Q_{N-1}=\frac{4}{z}\frac{(N+2m+n+\tfrac{1}{2})Q_{N}+n+\tfrac{1}{2}}{1-Q_{N}^{2}},$
    which is dPII (1.6) with $\zeta=4/z$, $\lambda=2m+n+\tfrac{1}{2}$ and $\rho=n+\tfrac{1}{2}$. Hence, although there is no alternating term, the structure of the even and odd terms is different.
Alternatively, using the solution $v_{m,n}^{(\mu)}(z)$, for fixed $m$ and $n$, then we have the following Lemma.
Lemma 5.6.
Consider the rational functions
$Q_{2N}(z)=X_{N}(z)=\frac{v_{m,N+n}^{(\mu)}+1}{v_{m,N+n}^{(\mu)}-1},\qquad Q_{2N+1}(z)=Y_{N}(z)=\frac{u_{N+n+1,m}^{(\mu+2m+1)}+1}{u_{N+n+1,m}^{(\mu+2m+1)}-1},$
with $v_{m,n}^{(\mu)}(z)$ and $u_{m,n}^{(\mu)}(z)$ as given in Theorem 4.4, then $Q_{N}(z)$ satisfies
$Q_{N+1}+Q_{N-1}=\frac{4}{z}\frac{(N+2n+1-\mu)Q_{N}+\mu+(m+\tfrac{1}{2})[1+(-1)^{N}]}{1-Q_{N}^{2}},$
which is asymmetric dPII (3.7) with $\lambda=2n+1-\mu$, $\rho=m+\tfrac{1}{2}+\mu$ and $\varphi=m+\tfrac{1}{2}$, whilst $X_{N}(z)$ and $Y_{N}(z)$ satisfy the discrete system
$X_{N+1}+X_{N}$ $=\frac{4}{z}\frac{(2N+2n+2-\mu)Y_{N}+\mu}{1-Y_{N}^{2}},$ (5.7a)
$Y_{N}+Y_{N-1}$ $=\frac{4}{z}\frac{(2N+2n+1-\mu)X_{N}+\mu+2m+1}{1-X_{N}^{2}}.$ (5.7b)
Proof.
From the transformations
$\mathcal{R}_{1}\big(v_{m,n}^{(\mu)};m+n+1,m-n+\mu,\mu\big)$ $=\big(u_{n+1,m}^{(\mu+2m+1)};n+1,-n-1-\mu,2m+1+\mu\big),$
$\mathcal{R}_{1}\big(u_{m,n}^{(\mu)};m,-m-2n-1+\mu,\mu\big)$ $=\big(v_{n,m}^{(\mu)};m+n+1,-m-n-1+\mu,-2n-1+\mu\big)$
we obtain the sequence of solutions
$v_{m,n}^{(\mu)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{n+1,m}^{(\mu+2m+1)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip v_{m,n+1}^{(\mu)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{n+2,m}^{(\mu+2m+1)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip v_{m,n+2}^{(\mu)}$
from which the result follows. We note that
$\mathcal{R}_{1}^{2}\Big(v_{m,n}^{(\mu)}\Big)=w_{m,n+1}^{(\mu)},\qquad\mathcal{R}_{1}^{2}\Big(u_{n,m}^{(\mu)}\Big)=u_{n+1,m}^{(\mu)}.$
5.1.2 Second discrete equation
The effect on the parameters of applying the transformations $\mathcal{R}_{2}$ (3.12) and $\mathcal{R}_{2}^{2}$ on a solution $w(z)$ of PV with $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}a^{2},-\tfrac{1}{2}b^{2},c\big)$ is given in Table 5.2.
[…]
Table 5.2: The effect on the parameters of applying the transformations $\mathcal{R}_{2}$ (3.12) and $\mathcal{R}_{2}^{2}$ on a solution $w(z)$ of PV with $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}a^{2},-\tfrac{1}{2}b^{2},c\big)$.
In the following example, we illustrate how using the Bäcklund transformation $\mathcal{R}_{2}$ (3.12) gives rise to a hierarchy of solutions of (3.17) in terms of generalised Laguerre polynomials.
Example 5.7.
Applying the Bäcklund transformation $\mathcal{R}_{2}$ (3.12), with $(a,b,c)=(-3,\mu,\mu)$, to the solution
$v_{1,1}^{(\mu)}(z)=1-\frac{z\big[z^{2}-2\mu z+\mu(\mu+1)\big]\big[z^{2}-2\mu z+\mu(\mu-1)\big]}{(z-\mu)\big[z^{4}-4\mu z^{3}+6\mu^{2}z^{2}-4\mu(\mu^{2}-1)z+\mu^{2}(\mu^{2}-1)\big]},$
we obtain the sequence of rational solutions of PV (1.2)
$v_{1,1}^{(\mu)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{2,1}^{(\mu+4)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip v_{1,2}^{(\mu+2)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{3,1}^{(\mu+6)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip v_{1,3}^{(\mu+4)}$
. Hence we obtain the solutions of equation (3.17)
$q_{0}(z)$ $=-1+\frac{2\mu}{z}+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{2}-2\mu+\mu(\mu+1)}{z^{2}-2\mu+\mu(\mu-1)},$
$q_{1}(z)$ $=1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{4}-4(\mu+1)z^{3}+6(\mu+1)^{2}z^{2}-4\mu(\mu+1)(\mu+2)z+\mu(\mu+1)^{2}(\mu+2)}{z^{2}-2\mu+\mu(\mu+1)}.$
Consequently, if we define the rational functions
$q_{2n}(z)$ $=x_{n}(z)=\frac{v_{1,n+1}^{(2n+\mu)}(z)+1}{v_{1,n+1}^{(2n+\mu)}(z)-1}=-1+\frac{2(\mu+2n)}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{0,n+2}^{(\mu-3)}}{T_{1,n+1}^{(\mu-3)}(z)},$
$q_{2n+1}(z)$ $=y_{n}(z)=\frac{u_{n+2,1}^{(2n+\mu+4)}(z)+1}{u_{n+2,1}^{(2n+\mu+4)}(z)-1}=1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{1,n+2}^{(\mu-3)}}{T_{0,n+2}^{(\mu-3)}(z)},$
then $q_{n}(z)$ satisfies
$\frac{n+\mu+2}{q_{n+1}+q_{n}}+\frac{n+\mu+1}{q_{n}+q_{n-1}}=\tfrac{1}{2}z-\frac{\big\{n+\mu+\tfrac{3}{2}[1+(-1)^{n}]\big\}q_{n}+3-\mu}{1-q_{n}^{2}},$
which is equation (3.17) with $\lambda=\mu+\tfrac{3}{2}$, $\rho=3-\mu$ and $\varphi=\tfrac{3}{2}$, whilst $x_{n}(z)$ and $y_{n}(z)$ satisfy the discrete system
$\frac{2n+\mu+3}{x_{n+1}+y_{n}}+\frac{2n+\mu+2}{x_{n}+y_{n}}$ $=\tfrac{1}{2}z-\frac{(2n+\mu+1)y_{n}+3-\mu}{1-y_{n}^{2}},$
$\frac{2n+\mu+2}{x_{n}+y_{n}}+\frac{2n+\mu+1}{x_{n}+y_{n-1}}$ $=\tfrac{1}{2}z-\frac{(2n+\mu+3)x_{n}+3-\mu}{1-x_{n}^{2}}.$
More generally, using the solution $v_{m,n}^{(\mu)}(z)$, for fixed $m$ and $n$, then we have the following Lemma.
Lemma 5.8.
Consider the rational functions
$Q_{2N}(z)=X_{N}(z)=\frac{v_{m,N+n}^{(2N+\mu)}(z)+1}{v_{m,N+n}^{(2N+\mu)}(z)-1},\qquad Q_{2N+1}(z)=Y_{N}(z)=\frac{u_{N+n+1,m}^{(2N+\mu+2m+2)}(z)+1}{u_{N+n+1,m}^{(2N+\mu+2m+2)}(z)-1},$
with $v_{m,n}^{(\mu)}(z)$ and $u_{m,n}^{(\mu)}(z)$ as given in Theorem 4.4, then $Q_{N}(z)$ satisfies
$\frac{N+m+\mu+1}{Q_{N+1}+Q_{N}}+\frac{N+m+\mu}{Q_{N}+Q_{N-1}}=\tfrac{1}{2}z-\frac{\big\{N+[1+(-1)^{N}](m+\tfrac{1}{2})+\mu\big\}Q_{N}+2n+1-\mu}{1-Q_{N}^{2}},$
which is equation (3.17) with $\lambda=m+\mu-\tfrac{1}{2}$, $\rho=2n-\mu+1$ and $\varphi=m+\tfrac{1}{2}$, whilst $X_{N}(z)$ and $Y_{N}(z)$ satisfy the discrete system
$\frac{2N+m+\mu+2}{X_{N+1}+Y_{N}}+\frac{2N+m+\mu+1}{X_{N}+Y_{N}}$ $=\tfrac{1}{2}z-\frac{(2N+\mu+1)Y_{N}+2n+1-\mu}{1-Y_{N}^{2}},$ (5.8a)
$\frac{2N+m+\mu+1}{X_{N}+Y_{N}}+\frac{2N+m+\mu}{X_{N}+Y_{N-1}}$ $=\tfrac{1}{2}z-\frac{(2N+2m+\mu+1)X_{N}+2n+1-\mu}{1-X_{N}^{2}}.$ (5.8b)
Proof.
From the transformations
$\mathcal{R}_{2}\big(v_{m,n}^{(\mu)};-m-n-1,m-n+\mu,\mu\big)$ $=\big(u_{n+1,m}^{(\mu+2m+2)};-n-1,-n+\mu,2m+2+\mu\big),$
$\mathcal{R}_{2}\big(u_{n,m}^{(\mu+2m+2)};-n,-n+1+\mu,2m+2+\mu\big)$ $=\big(v_{m,n}^{(\mu+2)};m+n+1,-m-n-1+\mu,2m-1+\mu\big),$
we obtain the sequence of solutions
$v_{m,n}^{(\mu)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{n+1,m}^{(\mu+2m+2)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip v_{m,n+1}^{(\mu+2)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{n+2,m}^{(\mu+2m+4)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip v_{m,n+2}^{(\mu+4)}$
from which the result follows. We note that
$\mathcal{R}_{1}^{2}\Big(v_{m,n}^{(\mu)}\Big)=v_{m,n+1}^{(\mu+2)},\qquad\mathcal{R}_{1}^{2}\Big(u_{n,m}^{(\mu)}\Big)=u_{n+1,m}^{(\mu+2)}.$
Alternatively, using the solution $w_{m,n}^{(\mu)}(z)$, for fixed $m$ and $n$, then we have the following Lemma.
Lemma 5.9.
Consider the rational functions
$Q_{2N}(z)=X_{N}(z)=\frac{w_{N+m,n}^{(\mu-2N)}(z)+1}{w_{N+m,n}^{(\mu-2N)}(z)-1},\qquad Q_{2N+1}(z)=Y_{N}(z)=\frac{v_{N+m,n}^{(\mu+2n-2N)}(z)+1}{v_{N+m,n}^{(\mu+2n-2N)}(z)-1},$
with $w_{m,n}^{(\mu)}(z)$ and $v_{m,n}^{(\mu)}(z)$ as given in Theorem 4.4, then $Q_{N}(z)$ satisfies
$\frac{N-n-\mu}{Q_{N+1}+Q_{N}}+\frac{N-n-\mu-1}{Q_{N}+Q_{N-1}}+\tfrac{1}{2}z+\frac{\big\{N-\mu-(n+\tfrac{1}{2})[1+(-1)^{N}]\big\}Q_{N}+2m+2n+1+\mu}{1-Q_{N}^{2}}=0,$
which is equation (3.17) with $\lambda=-n-\mu-\tfrac{1}{2}$, $\rho=2m+2n+\mu+1$ and $\varphi=-n-\tfrac{1}{2}$, whilst $X_{N}(z)$ and $Y_{N}(z)$ satisfy the discrete system
$\frac{2N-n-\mu+1}{X_{N+1}+Y_{N}}+\frac{2N-n-\mu}{X_{N}+Y_{N}}+\tfrac{1}{2}z+\frac{(2N-\mu+1)Y_{N}+2m+2n+1+\mu}{1-Y_{N}^{2}}=0,$ (5.9a)
$\frac{2N-n-\mu}{X_{N}+Y_{N}}+\frac{2N-n-\mu-1}{X_{N}+Y_{N-1}}+\tfrac{1}{2}z+\frac{(2N-2n-\mu-1)X_{N}+2m+2n+1+\mu}{1-X_{N}^{2}}=0.$ (5.9b)
Proof.
From the transformations
$\mathcal{R}_{2}^{-1}\big(w_{m,n}^{(\mu)};m,m+2n+1+\mu,\mu\big)$ $=\big(v_{m,n}^{(\mu+2n)};m+n+1,m+n+\mu,2n+\mu\big),$
$\mathcal{R}_{2}^{-1}\big(v_{m,n}^{(\mu+2n)};m+n+1,m+n+\mu,2n+\mu\big)$ $=\big(w_{m+1,n}^{(\mu-2)};m,m+2n+1+\mu,\mu\big)$
we obtain the sequence of solutions
$w_{m,n}^{(\mu)}\enskip\begin{array}[]{c}\mathcal{R}_{2}^{-1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip v_{m,n}^{(\mu+2n)}\enskip\begin{array}[]{c}\mathcal{R}_{2}^{-1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip w_{m+1,n}^{(\mu-2)}\enskip\begin{array}[]{c}\mathcal{R}_{2}^{-1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip v_{m,n}^{(\mu+2n-2)}\enskip\begin{array}[]{c}\mathcal{R}_{2}^{-1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip w_{m+2,n}^{(\mu-4)}$
from which the result follows. ∎
Remark 5.10.
From the transformations
$\mathcal{R}_{2}^{-1}\big(w_{m,n}^{(\mu)};m,-m-2n-1-\mu,\mu\big)$ $=\big(1/{u}_{n,m+1}^{(\mu+2m+2n+2n)};-n-k,-n-1,-2m-2n-1-\mu\big),$
$\mathcal{R}_{2}^{-1}\big(1/{u}_{n,m}^{(\mu)};2m+n+1-\mu,-n,\mu\big)$ $=\big(w_{m+1,n}^{(\mu-2m-2n-2)};m+1,m-\mu,-2m-2n-2+\mu\big),$
we obtain the sequence of solutions
$w_{m,n}^{(\mu)}\enskip\begin{array}[]{c}\mathcal{R}_{2}^{-1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\frac{1}{{u}_{n,m+1}^{(\mu+2m+2n+2)}}\enskip\begin{array}[]{c}\mathcal{R}_{2}^{-1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip w_{m+1,n+1}^{(\mu-2)}\enskip\begin{array}[]{c}\mathcal{R}_{2}^{-1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\frac{1}{{u}_{n+1,m+2}^{(\mu+2m+2n+4)}}\enskip\begin{array}[]{c}\mathcal{R}_{2}^{-1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip w_{m+2,n+2}^{(\mu-4)}.$
5.1.3 Third discrete equation: ternary discrete PI
The effect on the parameters of applying the transformations $\mathcal{R}_{3}$ (3.23), $\mathcal{R}_{3}^{2}$ and $\mathcal{R}_{3}^{3}$ on a solution $w(z)$ of PV with $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}a^{2},-\tfrac{1}{2}b^{2},c\big)$ is given in Table 5.3.
[…]
Table 5.3: The effect on the parameters of applying the transformation $\mathcal{R}_{3}$ (3.23), $\mathcal{R}_{3}^{2}$ and $\mathcal{R}_{3}^{3}$ on a solution $w(z)$ of PV with $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}a^{2},-\tfrac{1}{2}b^{2},c\big)$.
In the following example, we illustrate how using the Bäcklund transformation $\mathcal{R}_{3}$ (3.23) gives rise to a hierarchy of solutions of ternary dPI (3.27) in terms of generalised Laguerre polynomials.
Example 5.11.
Applying the Bäcklund transformation $\mathcal{R}_{3}$ (3.23), with $(a,b,c)=(\mu-1,-1,-\mu+3)$, to the solution
$\frac{1}{{w}_{1,1}^{(\mu-3)}(z)}=-\frac{z^{2}-2\mu z+\mu(\mu-1)}{(z-\mu+1)\big[z^{2}-2\mu z+\mu(\mu+1)\big]},$
we obtain the sequence of rational solutions of PV (1.2)
$\frac{1}{{w}_{1,1}^{(\mu-3)}}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip{v}_{1,1}^{(\mu)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{2,1}^{(\mu+3)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\frac{1}{{w}_{1,2}^{(\mu-4)}}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip{v}_{1,2}^{(\mu+1)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{3,1}^{(\mu+4)}$
so that the effect of applying $\mathcal{R}_{3}$ three times is
$\mathcal{R}_{3}^{3}\Big(1/{{w}_{1,1}^{(\mu-3)}}\Big)=1/{{w}_{1,1}^{(\mu-4)}},\qquad\mathcal{R}_{3}^{3}\Big({v}_{1,1}^{(\mu)}\Big)={v}_{1,2}^{(\mu+1)},\qquad\mathcal{R}_{3}^{3}\Big({u}_{2,1}^{(\mu+3)}\Big)={u}_{3,1}^{(\mu+4)}.$
Hence we obtain the rational solutions of the ternary dPI (3.27)
$x_{0}(z)$ $=\frac{w_{1,1}^{(\mu-3)}(z)}{1-w_{1,1}^{(\mu-3)}(z)}=-1+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{2}-2(\mu-1)z+\mu(\mu-1)}{z-\mu},$
$x_{1}(z)$ $=\frac{1}{v_{1,1}^{(\mu)}(z)-1}=-1+\frac{\mu}{z}+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{2}-2\mu z+\mu(\mu-1)}{z^{2}-2\mu z+\mu(\mu+1)},$
$x_{2}(z)$ $=\frac{1}{u_{2,1}^{(\mu+3)}(z)-1}=\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{4}-4\mu z^{3}+6\mu^{2}z^{2}-4\mu(\mu^{2}-1)z+\mu^{2}(\mu^{2}-1)z}{z^{2}-2(\mu+1)z+\mu(\mu-1)}.$
Consequently, if we define the rational functions
$x_{3n}(z)=X_{n}(z)$ $=\frac{w_{1,n+1}^{(\mu-n-3)}(z)}{1-w_{1,n+1}^{(\mu-n-3)}(z)}=-1+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{0,n+2}^{(\mu-n-4)}(z)}{T_{0,n+1}^{(\mu-n-2)}(z)},$
$x_{3n+1}(z)=Y_{n}(z)$ $=\frac{1}{v_{1,n+1}^{(\mu+n)}(z)-1}=-1+\frac{\mu+n}{z}+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{0,n+2}^{(\mu-n-3)}(z)}{T_{1,n+1}^{(\mu-n-3)}(z)},$
$x_{3n+2}(z)=Z_{n}(z)$ $=\frac{1}{u_{n+2,1}^{(\mu+n+3)}(z)-1}=\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{1,n+2}^{(\mu-n-4)}(z)}{T_{0,n+2}^{(\mu-n-4)}(z)},$
then $x_{n}(z)$ satisfies ternary dPI
$x_{n}(x_{n+1}+x_{n-1}+1)+\frac{a_{n}}{z}=0,$
with
$a_{n}=\tfrac{1}{3}(n+\mu+5)+\tfrac{2}{3}(\mu-1)\cos\!\big(\tfrac{2}{3}\pi n\big)+\tfrac{4}{9}\sqrt{3}\,\sin\!\big(\tfrac{2}{3}\pi n\big),$
whilst $X_{n}(z)$, $Y_{n}(z)$ and $Z_{n}(z)$ satisfy the discrete system
$X_{n}(Y_{n}+Z_{n-1}+1)+\frac{n+\mu+1}{z}=0,$
$Y_{n}(Z_{n}+X_{n}+1)+\frac{n+3}{z}=0,$
$Z_{n}(X_{n+1}+Y_{n}+1)+\frac{n+2}{z}=0.$
More generally, using the solution $w_{m,n}^{(\mu)}(z)$, for fixed $m$ and $n$, as the initial solution then we have the following Lemma.
Lemma 5.12.
Consider the rational functions
$X_{N}(z)=\frac{w_{m,N+n}^{(\mu-N)}(z)}{1-w_{m,N+n}^{(\mu-N)}(z)},\quad Y_{N}(z)=\frac{1}{v_{m,N+n}^{(\mu+N+2n+1)}(z)-1},\quad Z_{N}(z)=\frac{1}{u_{N+n+1,m}^{(\mu+N+2m+2n+2)}(z)-1},$ (5.10)
with $w_{m,n}^{(\mu)}(z)$, $v_{m,n}^{(\mu)}(z)$ and $u_{m,n}^{(\mu)}(z)$ as given in Theorem 4.4, then $X_{N}(z)$, $Y_{N}(z)$ and $Z_{N}(z)$ satisfy the discrete system
$X_{N}(Y_{N}+Z_{N-1}+1)+\frac{N+m+2n+\mu+1}{z}=0,$ (5.11a)
$Y_{N}(Z_{N}+X_{N}+1)+\frac{N+m+n+1}{z}=0,$ (5.11b)
$Z_{N}(X_{N+1}+Y_{N}+1)+\frac{N+n+1}{z}=0,$ (5.11c)
which is the system (3.28) with $a_{3N}=N+m+2n+\mu+1$, $a_{3N+1}=N+m+n+1$ and $a_{3N+2}=N+n+1$, i.e. (3.25a) with $\lambda=\tfrac{1}{3}(m+2n+2\mu+1)$, $\rho=\tfrac{1}{3}(2m+4n+\mu+2)$ and $\varphi=m+\tfrac{1}{3}$.
Proof.
From the transformations
$\mathcal{R}_{3}\big(1/{w}_{m,n}^{(\mu)};m+2n+1+\mu,-m,-\mu\big)=\big(v_{m,n}^{(\mu+2n+1)};m+n+1,m+n+\mu+1,2n+1+\mu\big),$
$\mathcal{R}_{3}\big(v_{m,n}^{(\mu+2n+1)};m+n+1,m+n+\mu+1,2n+1+\mu\big)=\big(u_{n+1,m}^{(\mu+2m+2n+2)};n+1,-n-\mu,2m+2n+2+\mu\big),$
$\mathcal{R}_{3}\big(u_{n+1,m}^{(\mu+2m+2n+2)};n+1,-n-\mu,2m+2n+2+\mu\big)=\big(1/{w}_{m,n+1}^{(\mu-1)};m+2n+\mu+2,-m,\mu-1\big),$
we obtain the sequence of solutions
$\frac{1}{{w}_{m,n}^{(\mu)}}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip v_{m,n}^{(\mu+2n+1)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{n+1,m}^{(\mu+2m+2n+2)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\frac{1}{{w}_{m,n+1}^{(\mu-1)}}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip v_{m,n+1}^{(\mu+2n+2)}$
from which the result follows. We note that
$\mathcal{R}_{3}^{3}\Big(1/w_{m,n}^{(\mu)}\Big)=1/w_{m,n+1}^{(\mu-1)},\qquad\mathcal{R}_{3}^{3}\Big({v}_{m,n}^{(\mu)}\Big)={v}_{m,n+1}^{(\mu+1)},\qquad\mathcal{R}_{3}^{3}\Big({u}_{n,m}^{(\mu)}\Big)={u}_{n+1,m}^{(\mu+1)}.$
Lemma 5.13.
Consider the rational functions
$X_{N}(z)=\frac{1}{v_{m,N+n}^{(N+\mu)}(z)-1},\quad Y_{N}(z)=\frac{1}{u_{N+n+1,m}^{(N+\mu+2m+1)}(z)-1},\quad Z_{N}(z)=\frac{w_{m,N+n+1}^{(-N+\mu-2n-2)}(z)}{1-w_{m,N+n+1}^{(-N+\mu-2n-2)}(z)},$ (5.12)
with $w_{m,n}^{(\mu)}(z)$, $v_{m,n}^{(\mu)}(z)$ and $u_{m,n}^{(\mu)}(z)$ as given in Theorem 4.4, then $X_{N}(z)$, $Y_{N}(z)$ and $Z_{N}(z)$ satisfy the discrete system
$X_{N}(Y_{N}+Z_{N-1}+1)+\frac{N+m+n+1}{z}=0,$ (5.13a)
$Y_{N}(Z_{N}+X_{N}+1)+\frac{N+n+1}{z}=0,$ (5.13b)
$Z_{N}(X_{N+1}+Y_{N}+1)+\frac{N+m+\mu+1}{z}=0,$ (5.13c)
which is the system (3.28) with $a_{3N}=N+m+n+1$, $a_{3N+1}=N+n+1$ and $a_{3N+2}=N+m+\mu+1$, i.e. (3.25a) with $\lambda=\tfrac{1}{3}(m+n-\mu+1)$, $\rho=\tfrac{1}{3}(2m+2n+\mu+2)$ and $\varphi=-m+n-\mu+\tfrac{1}{3}$.
Proof.
From the transformations
$\mathcal{R}_{3}\big(v_{m,n}^{(\mu)};m+n+1,m-n+\mu,\mu\big)$ $=\big(u_{n+1,m}^{(\mu+2m+1)};n+1,n+1-\mu,2m+1+\mu\big),$
$\mathcal{R}_{3}\big(u_{m,n}^{(\mu)};m,m+2n+1-\mu,\mu\big)$ $=\big(1/{w}_{n,m}^{(\mu-2m-2n-1)};-n+\mu,-n,2m+2n+1-\mu\big),$
$\mathcal{R}_{3}\big(1/{w}_{m,n}^{(\mu)};m+2n+1+\mu,-m,-\mu\big)$ $=\big(v_{m,n}^{(\mu+2n+1)};m+n+1,m+n+\mu+1,2n+1+\mu\big),$
we obtain the sequence of solutions
$v_{m,n}^{(\mu)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{n+1,m}^{(\mu+2m+1)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\frac{1}{{w}_{m,n+1}^{(\mu-2n-2)}}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip v_{m,n+1}^{(\mu+1)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{n+2,m}^{(\mu+2m+2)}$
from which the result follows. We note that
$\mathcal{R}_{3}^{3}\Big(1/w_{m,n}^{(\mu)}\Big)=1/w_{m,n+1}^{(\mu-1)},\qquad\mathcal{R}_{3}^{3}\Big({v}_{m,n}^{(\mu)}\Big)={v}_{m,n+1}^{(\mu+1)},\qquad\mathcal{R}_{3}^{3}\Big({u}_{n,m}^{(\mu)}\Big)={u}_{n+1,m}^{(\mu+1)}.$
Remark 5.14.
There are no solutions in terms of generalised Laguerre polynomials in the symmetric case when $\lambda=\varphi=0$.
5.1.4 Partial difference systems
In this subsection we derive some partial difference systems which have solutions in terms of generalised Laguerre polynomials.
Lemma 5.15.
If we define
$\mathcal{W}_{m,n}^{(\mu)}=\frac{{w}_{m,n}^{(\mu)}+1}{{w}_{m,n}^{(\mu)}-1},\qquad\mathcal{V}_{m,n}^{(\mu)}=\frac{{v}_{m,n}^{(\mu)}+1}{{v}_{m,n}^{(\mu)}-1},\qquad\mathcal{U}_{m,n}^{(\mu)}=\frac{{u}_{m,n}^{(\mu)}+1}{{u}_{m,n}^{(\mu)}-1},$ (5.14)
with $w_{m,n}^{(\mu)}(z)$, $v_{m,n}^{(\mu)}(z)$ and $u_{m,n}^{(\mu)}(z)$ as given in Theorem 4.4, then they satisfy the partial difference systems
$\mathcal{W}_{m+1,n}^{(\mu)}+\mathcal{W}_{m,n}^{(\mu)}$ $+\frac{4}{z}\frac{(2m+2n+\mu+2)\mathcal{V}_{m,n}^{(\mu+2n+1)}-\mu}{1-\big(\mathcal{V}_{m,n}^{(\mu+2n+1)}\big)^{\!2}}=0,$ (5.15a)
$\mathcal{V}_{m,n}^{(\mu+2n+1)}+\mathcal{V}_{m-1,n}^{(\mu+2n+1)}$ $+\frac{4}{z}\frac{(2m+2n+\mu+1)\mathcal{W}_{m,n}^{(\mu)}-2n-\mu-1}{1-\big(\mathcal{W}_{m,n}^{(\mu)}\big)^{\!2}}=0,$ (5.15b)
and
$\mathcal{V}_{m,n+1}^{(\mu)}+\mathcal{V}_{m,n}^{(\mu)}$ $=\frac{4}{z}\frac{(2n+2-\mu)\,\mathcal{U}_{n+1,m}^{(\mu+2m+1)}-\mu}{1-\big(\mathcal{U}_{n+1,m}^{(\mu+2m+1)}\big)^{\!2}},$ (5.16a)
$\mathcal{U}_{n+1,m}^{(\mu+2m+1)}+\mathcal{U}_{n,m}^{(\mu+2m+1)}$ $=\frac{4}{z}\frac{(2n+1-\mu)\mathcal{V}_{m,n}^{(\mu)}+2m+\mu+1}{1-\big(\mathcal{V}_{m,n}^{(\mu)}\big)^{\!2}}.$ (5.16b)
Proof.
These follow from setting $N=0$ in (5.6) and (5.7), respectively. ∎
Lemma 5.16.
If $\mathcal{W}_{m,n}^{(\mu)}$, $\mathcal{V}_{m,n}^{(\mu)}$ and $\mathcal{U}_{m,n}^{(\mu)}$ are given by (5.14), then they satisfy the partial difference systems
$\mathcal{W}_{m+1,n}^{(\mu)}+\mathcal{W}_{m,n}^{(\mu)}$ $+\frac{4}{z}\frac{(2n+\mu+1)\,\mathcal{V}_{m,n}^{(\mu+2n+1)}-\mu}{1-\big(\mathcal{V}_{m,n}^{(\mu+2n+1)}\big)^{\!2}}=0,$ (5.17a)
$\mathcal{V}_{m,n}^{(\mu+2n+1)}+\mathcal{V}_{m-1,n}^{(\mu+2n+1)}$ $+\frac{4}{z}\frac{(2m+2n+\mu+1)\mathcal{W}_{m,n}^{(\mu)}-(2n+\mu+1)}{1-\big(\mathcal{W}_{m,n}^{(\mu)}\big)^{\!2}}=0,$ (5.17b)
and
$\mathcal{W}_{m+1,n-1}^{(\mu)}+\mathcal{W}_{m,n}^{(\mu)}$ $+\frac{4}{z}\frac{(2n+\mu)\,\mathcal{U}_{n,m}^{(\mu+2m+2n+1)}-\mu}{1-\big(\mathcal{U}_{n,m}^{(\mu+2m+2n+1)}\big)^{\!2}}=0,$ (5.18a)
$\mathcal{U}_{n+1,m-1}^{(\mu+2m+2n+1)}+\mathcal{U}_{n,m}^{(\mu+2m+2n+1)}$ $+\frac{4}{z}\frac{(2n+\mu+1)\mathcal{W}_{m,n}^{(\mu)}-(2m+2n+\mu+1)}{1-\big(\mathcal{W}_{m,n}^{(\mu)}\big)^{\!2}}=0.$ (5.18b)
Proof.
From the transformations
$\mathcal{R}_{1}\big(w_{m,n}^{(\mu)};-m,m+2n+\mu+1,\mu\big)$ $=\big(v_{m-1,n}^{(\mu+2n+1)};-m-n,m+n+\mu,\mu+2n+1\big),$
$\mathcal{R}_{1}\big(v_{m,n}^{(\mu)};-m-n-1,m-n+\mu,\mu\big)$ $=\big(w_{m,n}^{(\mu-2n-1)};-m,m+\mu,\mu-2n-1\big),$
we obtain the sequence
$w_{m,n}^{(\mu)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip v_{m-1,n}^{(\mu+2n+1)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip w_{m-1,n}^{(\mu)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip v_{m-2,n}^{(\mu+2n+1)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip w_{m-2,n}^{(\mu)}$
from which the system (5.17) follows.
From the transformations
$\mathcal{R}_{1}\big(w_{m,n}^{(\mu)};m,m+2n+\mu+1,\mu\big)$ $=\big(u_{n,m}^{(\mu+2m+2n+1)};-n,n+\mu,\mu+2m+2n+1\big),$
$\mathcal{R}_{1}\big(u_{m,n}^{(\mu)};-m,-m-2n-1+\mu,\mu\big)$ $=\big(w_{n+1,m-1}^{(\mu-2m-2n-1)};n+1,-n-1+\mu,\mu-2m-2n-1\big),$
we obtain the sequence
$w_{m,n}^{(\mu)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{n,m}^{(\mu+2m+2n+1)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip w_{m+1,n-1}^{(\mu)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{n-1,m+1}^{(\mu+2m+2n+1)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip w_{m+2,n-2}^{(\mu)}$
from which the system (5.18) follows. ∎
Lemma 5.17.
If $\mathcal{W}_{m,n}^{(\mu)}$, $\mathcal{V}_{m,n}^{(\mu)}$ and $\mathcal{U}_{m,n}^{(\mu)}$ are given by (5.14), then they satisfy the partial difference systems
$\frac{m+\mu+1}{\mathcal{V}_{m,n}^{(\mu+1)}+\mathcal{U}_{n,m}^{(\mu+2m+1)}}+\frac{m+\mu}{\mathcal{V}_{m,n-1}^{(\mu-1)}+\mathcal{U}_{n,m}^{(\mu+2m+1)}}$ $=\tfrac{1}{2}z-\frac{\mu\,\mathcal{U}_{n,m}^{(\mu+2m+1)}+2n-\mu}{1-\big(\mathcal{U}_{n,m}^{(\mu+2m+1)}\big)^{\!2}},$ (5.19a)
$\frac{m+\mu}{\mathcal{V}_{m,n-1}^{(\mu-1)}+\mathcal{U}_{n,m}^{(\mu+2m+1)}}+\frac{m+\mu-1}{\mathcal{V}_{m,n-1}^{(\mu-1)}+\mathcal{U}_{n-1,m}^{(\mu+2m-1)}}$ $=\tfrac{1}{2}z-\frac{(2m+\mu)\mathcal{V}_{m,n-1}^{(\mu-1)}+2n-\mu}{1-\big(\mathcal{V}_{m,n-1}^{(\mu-1)}\big)^{\!2}},$ (5.19b)
and
$\frac{n+\mu-1}{\mathcal{W}_{m+1,n}^{(\mu-2)}+\mathcal{V}_{m,n}^{(\mu+2n)}}+\frac{n+\mu}{\mathcal{W}_{m,n}^{(\mu)}+\mathcal{V}_{m,n}^{(\mu+2n)}}$ $=\tfrac{1}{2}z-\frac{(\mu-1)\mathcal{V}_{m,n}^{(\mu+2n)}-(2m+2n+1+\mu)}{1-\big(\mathcal{V}_{m,n}^{(\mu+2n)}\big)^{\!2}},$ (5.20a)
$\frac{n+\mu}{\mathcal{W}_{m,n}^{(\mu)}+\mathcal{V}_{m,n}^{(\mu+2n)}}+\frac{n+\mu+1}{\mathcal{W}_{m,n}^{(\mu)}+\mathcal{V}_{m-1,n}^{(\mu+2n+2)}}$ $=\tfrac{1}{2}z-\frac{(\mu+2n+1)\mathcal{W}_{m,n}^{(\mu)}-(2m+2n+1+\mu)}{1-\big(\mathcal{W}_{m,n}^{(\mu)}\big)^{\!2}}.$ (5.20b)
Proof.
These follow from setting $N=0$ in (5.8) and (5.9), respectively. ∎
Lemma 5.18.
Consider the rational functions
$\mathcal{X}_{m,n}^{(\mu)}=\frac{w_{m,n}^{(\mu)}}{1-w_{m,n}^{(\mu)}},\qquad\mathcal{Y}_{m,n}^{(\mu)}=\frac{1}{v_{m,n}^{(\mu)}-1},\qquad\mathcal{Z}_{m,n}^{(\mu)}=\frac{1}{u_{m,n}^{(\mu)}-1},$
with $w_{m,n}^{(\mu)}(z)$, $v_{m,n}^{(\mu)}(z)$ and $u_{m,n}^{(\mu)}(z)$ as given in Theorem 4.4, then they satisfy the partial difference system
$\mathcal{X}_{m,n}^{(\mu)}\big(\mathcal{Y}_{m,n}^{(\mu+2n+1)}+\mathcal{Z}_{n,m}^{(\mu+2m+2n+1)}+1\big)+\frac{m+2n+\mu+1}{z}=0,$ (5.21a)
$\mathcal{Y}_{m,n}^{(\mu+2n+1)}\big(\mathcal{Z}_{n+1,m}^{(\mu+2m+2n+2)}+\mathcal{X}_{m,n}^{(\mu)}+1\big)+\frac{m+n+1}{z}=0,$ (5.21b)
$\mathcal{Z}_{n+1,m}^{(\mu+2m+2n+2)}\big(\mathcal{X}_{m,n+1}^{(\mu-1)}+\mathcal{Y}_{m,n}^{(\mu+2n+1)}+1\big)+\frac{n+1}{z}=0.$ (5.21c)
Proof.
This follows from setting $N=0$ in (5.11). ∎
5.2 Rational solutions in terms of generalised Umemura polynomials
5.2.1 First discrete equation: asymmetric discrete PII
In the following example, we illustrate how using the Bäcklund transformation $\mathcal{R}_{1}$ (3.3) gives rise to a hierarchy of solutions of asymmetric dPII (3.7) in terms of generalised Umemura polynomials.
Example 5.19.
Applying the Bäcklund transformation $\mathcal{R}_{1}$ with $(a,b,c)=(\kappa+\tfrac{1}{2},-\kappa+\tfrac{1}{2},2)$ to the solution
${w}_{2,1}^{(\kappa-3/2)}(z)=-\frac{\big[z^{2}-16\kappa(\kappa+1)\big]\big[z^{3}-6(2\kappa-1)z^{2}+48\kappa(\kappa-1)z-32\kappa(\kappa-1)(2\kappa-1)\big]}{\big[z^{2}-16\kappa(\kappa-1)\big]\big[z^{3}-6(2\kappa+1)z^{2}+48\kappa(\kappa+1)z-32\kappa(\kappa+1)(2\kappa+1)\big]},$
gives the sequence of rational solutions
${w}_{2,1}^{(\kappa-3/2)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip v_{2,1}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip w_{2,1}^{(\kappa-1/2)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip v_{2,1}^{(\kappa+1)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip w_{2,1}^{(\kappa+1/2)}.$
Hence we obtain solutions of asymmetric dPII (3.7)
$q_{0}(z)$ $=\frac{w_{2,1}^{(\kappa-3/2)}(z)+1}{w_{2,1}^{(\kappa-3/2)}(z)-1}=2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{4}-8\kappa z^{3}+128\kappa(\kappa^{2}-1)z-256\kappa^{2}(\kappa^{2}-1)}{z-4\kappa},$
$q_{1}(z)$ $=\frac{{v}_{2,1}^{(\kappa)}(z)+1}{{v}_{2,1}^{(\kappa)}(z)-1}=2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{2}-16\kappa(\kappa+1)}{z^{3}-6(2\kappa+1)z^{2}+48\kappa(\kappa+1)z-32\kappa(\kappa+1)(2\kappa+1)},$
$q_{2}(z)$ $=\frac{w_{2,1}^{(\kappa-1/2)}(z)+1}{w_{2,1}^{(\kappa-1/2)}(z)-1}=2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{4}-8(\kappa+1)+128\kappa(\kappa+1)(\kappa+2)-256\kappa(\kappa+1)^{2}(\kappa+2)}{z-4(\kappa+1)},$
$q_{3}(z)$ $=\frac{{v}_{2,1}^{(\kappa+1)}(z)+1}{{v}_{2,1}^{(\kappa+1)}(z)-1}=2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{2}-16(\kappa+1)(\kappa+2)}{z^{3}-6(2\kappa+3)z^{2}+48(\kappa+1)(\kappa+2)z-32(\kappa+1)(\kappa+2)(2\kappa+3)}.$
If we define the rational functions
$q_{2n}(z)$ $=x_{n}(z)=\frac{w_{2,1}^{(\kappa+n-3/2)}(z)+1}{w_{2,1}^{(\kappa+n-3/2)}(z)-1}=2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{2,1}^{(2\kappa+2n-3/2)}(z)}{U_{1,0}^{(2\kappa+2n+1/2)}(z)}$
$=2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{4}-8(n+\kappa)z^{3}+128(n+\kappa)\big[(n+\kappa)^{2}-1\big]z-256(n+\kappa)^{2}\big[(n+\kappa)^{2}-1\big]}{z-4(n+\kappa)},$
$q_{2n+1}(z)$ $=y_{n}(z)=\frac{{v}_{2,1}^{(\kappa+n)}(z)+1}{{v}_{2,1}^{(\kappa+n)}(z)-1}=2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{2,0}^{(2\kappa+2n+1/2)}(z)}{U_{1,1}^{(2\kappa+2n+1/2)}(z)}$
$=2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{3}-6(2n+2\kappa+1)z^{2}+48(n+\kappa)(n+\kappa+1)-32(n+\kappa)(n+\kappa+1)(2n+2\kappa+1)}{z^{2}-16(n+\kappa)(n+\kappa+1)},$
then $q_{n}(z)$ satisfies
$q_{n+1}+q_{n-1}=\frac{4}{z}\frac{(n+2\kappa)q_{n}+2-(-1)^{n}}{1-q_{n}^{2}},$
which is asymmetric dPII (3.7) with $\lambda=2\kappa$, $\rho=2$ and $\varphi=-1$, whilst $x_{n}(z)$ and $y_{n}(z)$ satisfy the discrete system
$x_{n+1}+x_{n}$ $=\frac{4}{z}\frac{(2n+2\kappa+1)y_{n}+3}{1-y_{n}^{2}},\qquad y_{n}+y_{n-1}=\frac{4}{z}\frac{2(n+\kappa)x_{n}+1}{1-x_{n}^{2}}.$
More generally, using the solution $\widehat{w}_{m,n}^{(\kappa)}$, for fixed $m$ and $n$, then we have the following Lemma.
Lemma 5.20.
Consider the rational functions
$Q_{2N}(z)=X_{N}(z)=\frac{\widehat{w}_{m,n}^{(N+\kappa)}(z)+1}{\widehat{w}_{m,n}^{(N+\kappa)}(z)-1},\qquad Q_{2N+1}(z)=Y_{N}(z)=\frac{\widehat{v}_{m,n}^{(N+\kappa+n+1/2)}(z)+1}{\widehat{v}_{m,n}^{(N+\kappa+n+1/2)}(z)-1},$
with $\widehat{w}_{m,n}^{(\kappa)}(z)$ and $\widehat{v}_{m,n}^{(\kappa)}(z)$ as given in Theorem 4.9, then $Q_{N}(z)$ satisfies
$Q_{N+1}+Q_{N-1}=\frac{4}{z}\frac{(N+m+n+2\kappa)Q_{N}+m-n(-1)^{N}}{1-Q_{N}^{2}},$
which is asymmetric dPII (3.7) with $\lambda=m+n+2\kappa$, $\rho=m$ and $\varphi=-n$, whilst $X_{N}(z)$ and $Y_{N}(z)$ satisfy the discrete system
$X_{N+1}+X_{N}$ $=\frac{4}{z}\frac{(2N+m+n+2\kappa+1)Y_{N}+m+n}{1-Y_{N}^{2}},$ (5.22a)
$Y_{N}+Y_{N-1}$ $=\frac{4}{z}\frac{(2N+m+n+2\kappa)X_{N}+m-n}{1-X_{N}^{2}}.$ (5.22b)
Proof.
From the transformations
$\mathcal{R}_{1}\big(\widehat{w}_{m,n}^{(\kappa)};m+\kappa,-n-\kappa,m+n\big)$ $=\big(\widehat{v}_{m,n}^{(\kappa+n+1/2)};m+n+\kappa+\tfrac{1}{2},-\kappa-\tfrac{1}{2},m-n\big),$
$\mathcal{R}_{1}\big(\widehat{v}_{m,n}^{(\kappa)};m+\kappa,n-\kappa,m-n\big)$ $=\big(\widehat{w}_{m,n}^{(\kappa-n+1/2)};m-n+\kappa+\tfrac{1}{2},-\kappa-\tfrac{1}{2},m+n\big),$
we obtain the sequence of solutions
$\widehat{w}_{m,n}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{v}_{m,n}^{(\kappa+n+1/2)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{m,n}^{(\kappa+1)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{v}_{m,n}^{(\kappa+n+3/2)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{m,n}^{(\kappa+2)}$
from which the result follows. We note that
$\mathcal{R}_{1}^{2}\Big(\widehat{w}_{m,n}^{(\kappa)}\Big)=\widehat{w}_{m,n}^{(\kappa+1)},\qquad\mathcal{R}_{1}^{2}\Big(\widehat{v}_{m,n}^{(\kappa)}\Big)=\widehat{v}_{m,n}^{(\kappa+1)},$
so that only the parameter changes. ∎
Alternatively, using the solution $\widehat{v}_{m,n}^{(\kappa)}$, for fixed $m$ and $n$, then we have the following Lemma.
Lemma 5.21.
Consider the rational functions
$Q_{2N}(z)=X_{N}(z)=\frac{\widehat{v}_{N+m,N+n}^{(\kappa)}(z)+1}{\widehat{v}_{N+m,N+n}^{(\kappa)}(z)-1},\qquad Q_{2N+1}(z)=Y_{N}(z)=\frac{\widehat{u}_{N+m,N+n}^{(2\kappa-2N-2n)}(z)+1}{\widehat{u}_{N+m,N+n}^{(2\kappa-2N-2n)}(z)-1},$
with $\widehat{v}_{m,n}^{(\kappa)}(z)$ and $\widehat{u}_{m,n}^{(\kappa)}(z)$ as given in Theorem 4.9, then $Q_{N}(z)$ satisfies
$Q_{N+1}+Q_{N-1}=\frac{4}{z}\frac{(N+m+n)Q_{N}+m-n+\kappa[1+(-1)^{N}]}{1-Q_{N}^{2}},$
which is asymmetric dPII (3.7) with $\lambda=m+n$, $\rho=m-n+\kappa$ and $\varphi=\kappa$, whilst $X_{N}(z)$ and $Y_{N}(z)$ satisfy the discrete system
$X_{N+1}+X_{N}$ $=\frac{4}{z}\frac{(2N+m+n+1)Y_{N}+m-n}{1-Y_{N}^{2}},$ (5.23a)
$Y_{N}+Y_{N-1}$ $=\frac{4}{z}\frac{(2N+m+n)X_{N}+2\kappa+m-n}{1-X_{N}^{2}}.$ (5.23b)
Proof.
From the transformations
$\mathcal{R}_{1}\big(\widehat{v}_{m,n}^{(\kappa)};m+\kappa,-n+\kappa,m-n\big)$ $=\big(\widehat{u}_{m,n}^{(2\kappa-2n)};m+\tfrac{1}{2},-n-\tfrac{1}{2},m-n+2\kappa\big),$
$\mathcal{R}_{1}\big(\widehat{u}_{m,n}^{(2\kappa)};m+\tfrac{1}{2},-n-\tfrac{1}{2},m+n+2\kappa\big)$ $=\big(\widehat{v}_{m+1,n+1}^{(n+\kappa)};m+n+\kappa+1,\kappa-1,m-n\big),$
we obtain the sequence of solutions
$\widehat{v}_{m,n}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{m,n}^{(2\kappa-2n)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{v}_{m+1,n+1}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{m+1,n+1}^{(2\kappa-2n-2)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{v}_{m+2,n+2}^{(\kappa)}$
from which the result follows. We note that
$\mathcal{R}_{1}^{2}\Big(\widehat{w}_{m,n}^{(\kappa)}\Big)=\widehat{w}_{m+1,n+1}^{(\kappa)},\qquad\mathcal{R}_{1}^{2}\Big(\widehat{u}_{m,n}^{(2\kappa-2n)}\Big)=\widehat{u}_{m+1,n+1}^{(2\kappa-2n-2)}.$
Remark 5.22.
From the transformations
$\mathcal{R}_{1}\big(\widehat{w}_{m,n}^{(\kappa)};m+\kappa,n+\kappa,m+n\big)$ $=\big(\widehat{u}_{m,n-1}^{(2\kappa+1)};m+\tfrac{1}{2},n-\tfrac{1}{2},m+n+2\kappa\big),$ (5.24a)
$\mathcal{R}_{1}\big(u_{m,n-1}^{(2\kappa+1)};m+\tfrac{1}{2},n-\tfrac{1}{2},m+n+2\kappa\big)$ $=\big(\widehat{w}_{m+1,n-1}^{(\kappa)};m+\kappa+1,n+\kappa-1,m+n\big),$ (5.24b)
we obtain the sequence
$\widehat{w}_{m,n}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{m,n-1}^{(2\kappa+1)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{m+1,n-1}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{m,n-2}^{(2\kappa+1)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{m+2,n-2}^{(\kappa)}.$
It might be expected that the sequence would terminate when the second index becomes zero. However it can be shown that
$\widehat{w}_{m,1}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{m,0}^{(2\kappa+1)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{m+1,0}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{m+1,0}^{(2\kappa)}$
and
$\widehat{u}_{m+1,0}^{(2\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{v}_{m+2,1}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{m+2,1}^{(2\kappa-2)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{v}_{m+3,2}^{(\kappa)}.$
This is illustrated in the following example.
Example 5.23.
Consider the rational solution
$\widehat{w}_{1,1}^{(\kappa)}(z)=-\frac{(z-4\kappa-2)(z+4\kappa+6)}{(z+4\kappa+2)(z-4\kappa-6)},$
which satisfies PV with $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(\kappa+1)^{2},\tfrac{1}{2}(\kappa+1)^{2},2\big)$, and so using (5.24) we obtain the sequence
$\widehat{w}_{1,1}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{1,0}^{(2\kappa+1)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{2,0}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{2,0}^{(2\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{v}_{3,1}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{3,1}^{(2\kappa-2)}.$
The first few solutions are
$q_{0}(z)$ $=\frac{\widehat{w}_{1,1}^{(\kappa)}(z)+1}{\widehat{w}_{1,1}^{(\kappa)}(z)-1}=\frac{4z}{z^{2}-4(2\kappa+1)(2\kappa+3)},$
$q_{1}(z)$ $=\frac{4(\kappa+1)}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z-4\kappa-6}{z-4\kappa-2},$
$q_{2}(z)$ $=\frac{6\big[z^{2}-8(\kappa+1)z+4(2\kappa+1)(2\kappa+3)\big]}{z^{3}-12(\kappa+1)z^{2}+12(2\kappa+1)(2\kappa+3)z-16(\kappa+1)(2\kappa+1)(2\kappa+3)}-\frac{2}{z-4\kappa-4},$
$q_{3}(z)$ $=\frac{4(\kappa+1)}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{3}-6(2\kappa+3)z^{2}+48(\kappa+1)(\kappa+2)z-32(\kappa+1)(\kappa+2)(2\kappa+3)}{z^{3}-6(2\kappa+1)z^{2}+48\kappa(\kappa+1)z-32\kappa(\kappa+1)(2\kappa+1)},$
and satisfy
$q_{n+1}+q_{n-1}=\frac{4}{z}\frac{nq_{n}+2+[1+(-1)^{n}]\kappa}{1-q_{n}^{2}},$ (5.25)
which is asymmetric dPII (3.7) with $\lambda=0$, $\rho=2+\kappa$ and $\varphi=\kappa$. In general
$q_{2n}=x_{n}=2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{n+1,n-2}^{(2\kappa-2n+3)}}{U_{n,n-1}^{(2\kappa-2n+3)}},\qquad q_{2n+1}=y_{n}=\frac{4(\kappa+1)}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{n+1,n-1}^{(2\kappa-2n+3)}}{U_{n+1,n-1}^{(2\kappa-2n+1)}},$
satisfy (5.25) and
$x_{n+1}+x_{n}$ $=\frac{4}{z}\,\frac{(2n+1)y_{n}+2}{1-y_{n}^{2}},\qquad y_{n}+y_{n-1}=\frac{4}{z}\,\frac{2nx_{n}+2\kappa+2}{1-x_{n}^{2}}.$
Symmetric case. In the symmetric case when $\varphi=0$ in (3.7), then
$q_{n}(z)=\frac{\widehat{w}_{m,0}^{(\kappa+n/2)}(z)+1}{\widehat{w}_{m,0}^{(\kappa+n/2)}(z)-1},\qquad p_{n}(z)=\frac{\widehat{w}_{0,m}^{(\kappa-n/2)}(z)+1}{\widehat{w}_{0,m}^{(\kappa-n/2)}(z)-1},$
where $\widehat{w}_{m,0}^{(\kappa)}(z)$ and $\widehat{w}_{0,m}^{(\kappa)}(z)$ are given by (4.20c) and (4.20d) respectively, which satisfy
$q_{n+1}+q_{n-1}$ $=\frac{4}{z}\frac{(n+m+2\kappa)q_{n}+m}{1-q_{n}^{2}},\qquad p_{n+1}+p_{n-1}=\frac{4}{z}\frac{(n-m-2\kappa)p_{n}+m}{1-p_{n}^{2}}.$
5.2.2 Second discrete equation
In the following example, we illustrate how using the Bäcklund transformation $\mathcal{R}_{2}$ (3.12) gives rise to a hierarchy of solutions of (3.17) in terms of generalised Umemura polynomials.
Example 5.24.
Applying the Bäcklund transformation $\mathcal{R}_{2}$ (3.12), with $(a,b,c)=(-\kappa-1,-\kappa-m,m+1)$ to the solution $\widehat{w}_{1,m}^{(\kappa)}(z)$ gives the sequence
$\widehat{w}_{1,m}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{v}_{2,m}^{(\kappa+m-1/2)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{3,m}^{(\kappa-1)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{v}_{4,m}^{(\kappa+m-3/2)}$
Hence if we define the rational functions
$q_{2n}(z)$ $=x_{n}(z)=\frac{\widehat{w}_{2n+1,m}^{(\kappa-n)}(z)+1}{\widehat{w}_{2n+1,m}^{(\kappa-n)}(z)-1}=2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{2n+1,m}^{(2\kappa-2n)}(z)}{U_{2n,m-1}^{(2\kappa-2n+2)}(z)},$
$q_{2n+1}(z)$ $=y_{n}(z)=\frac{\widehat{v}_{2n+2,m}^{(\kappa+m-n-1/2)}(z)+1}{\widehat{v}_{2n+2,m}^{(\kappa+m-n-1/2)}(z)-1}=2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{2n+2,m-1}^{(2\kappa-2n-2m+2)}(z)}{U_{2n+1,m}^{(2\kappa-2n-2m+2)}(z)},$
then $q_{n}(z)$ satisfies
$\frac{2n+3}{q_{n+1}+q_{n}}+\frac{2n+1}{q_{n}+q_{n-1}}=z-\frac{2\big[n+1-m(-1)^{n}\big]q_{n}+4\kappa+2m+2}{1-q_{n}^{2}},$
which is equation (3.17) with $\lambda=m$, $\varphi=-n$ and $\rho=m+n+2\kappa$, whilst $x_{n}(z)$ and $y_{n}(z)$ satisfy the discrete system
$\frac{4n+5}{x_{n+1}+y_{n}}+\frac{4n+3}{x_{n}+y_{n}}$ $=z-\frac{2(2n+m+1)y_{n}+4\kappa+2m+2}{1-y_{n}^{2}},$
$\frac{4n+3}{x_{n}+y_{n}}+\frac{4n+1}{x_{n}+y_{n-1}}$ $=z-\frac{2(2n-m+1)x_{n}+4\kappa+2m+2}{1-x_{n}^{2}}.$
More generally, using the solution $\widehat{w}_{m,n}^{(\kappa)}$, for fixed $m$ and $n$, then we have the following Lemma.
Lemma 5.25.
Consider the rational functions
$Q_{2N}(z)=X_{N}(z)=\frac{\widehat{w}_{2N+m,n}^{(\kappa-N)}(z)+1}{\widehat{w}_{2N+m,n}^{(\kappa-N)}(z)-1},\qquad Q_{2N+1}(z)=Y_{N}(z)=\frac{\widehat{v}_{2N+m+1,n}^{(\kappa+n-N-1/2)}(z)+1}{\widehat{v}_{2N+m+1,n}^{(\kappa+n-N-1/2)}(z)-1},$
with $\widehat{w}_{m,n}^{(\kappa)}(z)$ and $\widehat{v}_{m,n}^{(\kappa)}(z)$ as given in Theorem 4.9, then $Q_{N}(z)$ satisfies
$\frac{2N+2m+1}{Q_{N+1}+Q_{N}}+\frac{2N+2m-1}{Q_{N}+Q_{N-1}}=z-\frac{2\big[N+m-n(-1)^{N}\big]Q_{N}+2(m+n+2\kappa)}{1-Q_{N}^{2}},$
which is equation (3.17) with $\lambda=m$, $\varphi=-n$ and $\rho=m+n+2\kappa$, whilst $X_{N}(z)$ and $Y_{N}(z)$ satisfy the discrete system
$\frac{4N+2m+3}{X_{N+1}+Y_{N}}+\frac{4N+2m+1}{X_{N}+Y_{N}}$ $=z-\frac{2[(2N+m+n+1)Y_{N}+m+n+2\kappa]}{1-Y_{N}^{2}},$ (5.26a)
$\frac{4N+2m+1}{X_{N}+Y_{N}}+\frac{4N+2m-1}{X_{N}+Y_{N-1}}$ $=z-\frac{2[(2N+m-n)X_{N}+m+n+2\kappa]}{1-X_{N}^{2}}.$ (5.26b)
Proof.
From the transformations
$\mathcal{R}_{2}\big(\widehat{w}_{m,n}^{(\kappa)};-m-\kappa,-n-\kappa,m+n\big)$ $=\big(\widehat{v}_{m+1,n}^{(\kappa+n-1/2)};-m-n-\kappa-\tfrac{1}{2},-\kappa+\tfrac{1}{2},m-n+1\big),$
$\mathcal{R}_{2}\big(\widehat{v}_{m,n}^{(\kappa)};-m-\kappa,n-\kappa,m-n\big)$ $=\big(\widehat{w}_{m+1,n}^{(\kappa-n-1/2)};-m-\kappa-\tfrac{1}{2},n-\kappa+\tfrac{1}{2},m+n+1\big)$
we obtain the sequence of solutions
$\widehat{w}_{m,n}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{v}_{m+1,n}^{(\kappa+n-1/2)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{m+2,n}^{(\kappa-1)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{v}_{m+3,n}^{(\kappa+n-3/2)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{m+4,n}^{(\kappa-2)}$
from which the result follows. We note that
$\mathcal{R}_{2}^{2}\Big(\widehat{w}_{m,n}^{(\kappa)}\Big)=\widehat{w}_{m+2,n}^{(\kappa-1)},\qquad\mathcal{R}_{2}^{2}\Big(\widehat{v}_{m,n}^{(\kappa)}\Big)=\widehat{v}_{m+2,n}^{(\kappa-1)}.$
When $n=0$ there is no alternating term and so
$Q_{2N}(z)=\frac{\widehat{w}_{2N+m,0}^{(\kappa-N)}(z)+1}{\widehat{w}_{2N+m,0}^{(\kappa-N)}(z)-1},\qquad Q_{2N+1}(z)=\frac{\widehat{v}_{2N+m+1,0}^{(\kappa-N-1/2)}(z)+1}{\widehat{v}_{2N+m+1,0}^{(\kappa-N-1/2)}(z)-1},$
satisfy
$\frac{2N+2m+1}{Q_{N+1}+Q_{N}}+\frac{2N+2m-1}{Q_{N}+Q_{N-1}}=z-\frac{2(N+m)Q_{N}+2m+4\kappa}{1-Q_{N}^{2}}.$
It is also possible to obtain a sequence of solutions involving the solutions $\widehat{w}_{m,n}^{(\kappa)}$ and $\widehat{u}_{m,n}^{(2\kappa+1)}$ as illustrated in the following Lemma.
Lemma 5.26.
Consider the rational functions
$Q_{2N}(z)=X_{N}(z)=\frac{\widehat{w}_{N+m,N+n}^{(\kappa)}(z)+1}{\widehat{w}_{N+m,N+n}^{(\kappa)}(z)-1},\qquad Q_{2N+1}(z)=Y_{N}(z)=\frac{\widehat{u}_{N+m,N+n}^{(2\kappa+1)}(z)+1}{\widehat{u}_{N+m,N+n}^{(2\kappa+1)}(z)-1},$
with $\widehat{w}_{m,n}^{(\kappa)}(z)$ and $\widehat{u}_{m,n}^{(\kappa)}(z)$ as given in Theorem 4.9, then $Q_{N}(z)$ satisfies
$\frac{2N+2m+2n+2\kappa+1}{Q_{N+1}+Q_{N}}+\frac{2N+2m+2n+2\kappa-1}{Q_{N}+Q_{N-1}}=z-\frac{2\big[N+m+n+\kappa\big\{1+(-1)^{N}\big\}\big]Q_{N}+2(m-n)}{1-Q_{N}^{2}},$
which is equation (3.17) with $\lambda=m+n+\kappa$, $\varphi=\kappa$ and $\rho=m-n$, whilst $X_{N}(z)$ and $Y_{N}(z)$ satisfy the discrete system
$\frac{4N+2m+2n+2\kappa+3}{X_{N+1}+Y_{N}}+\frac{4N+2m+2n+2\kappa+1}{X_{N}+Y_{N}}$ $=z-\frac{2[(2N+m+n+1)Y_{N}+m-n]}{1-Y_{N}^{2}},$ (5.27a)
$\frac{4N+2m+2n+2\kappa+1}{X_{N}+Y_{N}}+\frac{4N+2m+2n+2\kappa-1}{X_{N}+Y_{N-1}}$ $=z-\frac{2[(2N+m+n+2\kappa)X_{N}+m-n]}{1-X_{N}^{2}}.$ (5.27b)
Proof.
From the transformations
$\mathcal{R}_{2}\big(\widehat{w}_{m,n}^{(\kappa)};-m-\kappa,n+\kappa,m+n\big)$ $=\big(\widehat{u}_{m,n}^{(2\kappa+1)};-m-\tfrac{1}{2},n+\tfrac{1}{2},m+n+2\kappa+1\big),$
$\mathcal{R}_{2}\big(\widehat{u}_{m,n}^{(2\kappa+1)};-m-\tfrac{1}{2},n+\tfrac{1}{2},m+n+2\kappa+1\big)$ $=\big(\widehat{w}_{m+1,n+1}^{(\kappa)};-m-\kappa-1,n+\kappa+1,m+n+2\big).$
we obtain the sequence of solutions
$\widehat{w}_{m,n}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{m,n}^{(2\kappa+1)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{m+1,n+1}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{m+1,n+1}^{(2\kappa+1)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{m+2,n+2}^{(\kappa)}$
from which the result follows. We note that
$\mathcal{R}_{2}^{2}\Big(\widehat{w}_{m,n}^{(\kappa)}\Big)=\widehat{w}_{m+1,n+1}^{(\kappa)},\qquad\mathcal{R}_{2}^{2}\Big(\widehat{u}_{m,n}^{(2\kappa+1)}\Big)=\widehat{u}_{m+1,n+1}^{(2\kappa+1)}.$
When $\kappa=0$ then there is no alternating term and so
$Q_{2N}(z)=\frac{\widehat{w}_{N+m,N+n}^{(0)}(z)+1}{\widehat{w}_{N+m,N+n}^{(0)}(z)-1},\qquad Q_{2N+1}(z)=\frac{\widehat{u}_{N+m,N+n}^{(1)}(z)+1}{\widehat{u}_{N+m,N+n}^{(1)}(z)-1},$
satisfy
$\frac{2N+2m+2n+1}{Q_{N+1}+Q_{N}}+\frac{2N+2m+2n-1}{Q_{N}+Q_{N-1}}=z-\frac{2\big[(N+m+n)Q_{N}+m-n\big]}{1-Q_{N}^{2}},$
which is equation (3.17) with $\lambda=m+n$, $\rho=m-n$ and $\varphi=0$. As previously, although there is no alternating term, the structure of the even and odd terms is different.
Symmetric case. In the symmetric case when $\varphi=0$ in (3.18), from the transformations
$\mathcal{R}_{2}\big(\widehat{w}_{m,0}^{(\kappa)};-m-\kappa,\kappa,m\big)$ $=\big(\widehat{w}_{m+1,0}^{(\kappa-1/2)};-m-\kappa-\tfrac{1}{2},-\kappa+\tfrac{1}{2},m+1\big),$
$\mathcal{R}_{2}\big(\widehat{w}_{0,n}^{(\kappa)};n,n+\tfrac{1}{2},n+\kappa\big)$ $=\big(\widehat{w}_{0,n+1}^{(\kappa-1/2)};\kappa-\tfrac{1}{2},n+\kappa+\tfrac{1}{2},n+1\big),$
recall (4.20c) and (4.20d), we obtain the respective sequences
$\begin{split}&\widehat{w}_{m,0}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{m+1,0}^{(\kappa-1/2)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{m+2,0}^{(\kappa-1)}\\ &\widehat{w}_{0,n}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{0,n+1}^{(\kappa-1/2)}\enskip\begin{array}[]{c}\mathcal{R}_{2}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{0,n+2}^{(\kappa-1)}\end{split}.$
If we define
$q_{n}(z)=\frac{\widehat{w}_{m+n,0}^{(\kappa-n/2)}(z)+1}{\widehat{w}_{m+n,0}^{(\kappa-n/2)}(z)-1},\qquad p_{n}(z)=\frac{\widehat{w}_{0,n+m}^{(\kappa-n/2)}(z)+1}{\widehat{w}_{0,n+m}^{(\kappa-n/2)}(z)-1},$
then ${q}_{n}(z)$ and $p_{n}(z)$ satisfy
$\frac{2n+2m+1}{q_{n+1}+q_{n}}+\frac{2n+2m-1}{q_{n}+q_{n-1}}$ $=z-\frac{2(n+m)q_{n}+2m+4\kappa}{1-q_{n}^{2}},$
$\frac{2n+2m+1}{p_{n+1}+p_{n}}+\frac{2n+2m-1}{p_{n}+p_{n-1}}$ $=z-\frac{2(n+m)p_{n}-2m-4\kappa}{1-p_{n}^{2}},$
which are the discrete equation (3.18) with $\lambda=m$, $\rho=2\kappa+m$, $\varphi=0$ and $\lambda=m$, $\rho=-2\kappa-m$, $\varphi=0$, respectively.
5.2.3 Third discrete equation: ternary discrete PI
In the following example, we illustrate how using the Bäcklund transformation $\mathcal{R}_{3}$ (3.23) gives rise to a hierarchy of solutions of ternary dPI (3.27) in terms of generalised Umemura polynomials.
Example 5.27.
Applying the Bäcklund transformation $\mathcal{R}_{3}$ (3.23), with $(a,b,c)=(\kappa+1,-\kappa-m,m+1)$ to the solution ${w}_{1,m}^{(\kappa)}$ gives the sequence of rational solutions
$\widehat{w}_{1,m}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{v}_{1,m}^{(\kappa+m+1/2)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{1,m}^{(2\kappa+1)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{2,m}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{v}_{2,m}^{(\kappa+m+1/2)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{2,m}^{(2\kappa+1)}$
so that
$\mathcal{R}_{3}^{3}\Big(\widehat{w}_{1,m}^{(\kappa)}\Big)=\widehat{w}_{2,m}^{(\kappa)},\qquad\mathcal{R}_{3}^{3}\Big(\widehat{v}_{1,m}^{(\kappa+m+1/2)}\Big)=\widehat{v}_{2,m}^{(\kappa+m+1/2)},\qquad\mathcal{R}_{3}^{3}\Big(\widehat{u}_{1,m}^{(2\kappa+1)}\Big)=\widehat{u}_{2,m}^{(2\kappa+1)}.$
Hence if we define the rational functions
$x_{3n}(z)={X}_{n}(z)$ $=\frac{1}{\widehat{w}_{n+1,m}^{(\kappa)}(z)-1}=-\frac{1}{2}+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{n+1,m}^{(2\kappa)}(z)}{U_{n,m-1}^{(2\kappa+2)}(z)},$
$x_{3n+1}(z)={Y}_{n}(z)$ $=\frac{1}{\widehat{v}_{n+1,m}^{(\kappa+m+1/2)}(z)-1}=-\frac{1}{2}+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{n+1,m-1}^{(2\kappa+2)}(z)}{U_{n,m}^{(2\kappa+2)}(z)},$
$x_{3n+2}(z)={Z}_{n}(z)$ $=\frac{1}{\widehat{u}_{n+1,m}^{(2\kappa+1)}(z)-1}=-\frac{1}{2}+\frac{n+m+2\kappa+2}{z}+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{U_{n+1,m}^{(2\kappa+2)}(z)}{U_{n+1,m}^{(2\kappa)}(z)},$
then $x_{n}(z)$ satisfies ternary dPI
$x_{n}(x_{n+1}+x_{n-1}+1)+\frac{a_{n}}{z}=0,$
with
$a_{n}=\tfrac{1}{3}(n+m+2\kappa)+1+\tfrac{1}{3}(\kappa-m)\cos\!\big(\tfrac{2}{3}\pi n\big)+\tfrac{1}{9}(3m+3\kappa+1)\sqrt{3}\sin\!\big(\tfrac{2}{3}\pi n\big),$
so
$a_{3n}=n+\kappa+1,\qquad a_{3n+1}=n+m+\kappa+\tfrac{3}{2},\qquad a_{3n+2}=n+\tfrac{3}{2},$
whilst ${X}_{n}(z)$, ${Y}_{n}(z)$ and ${Z}_{n}(z)$ satisfy
${X}_{n}\big({Y}_{n}+{Z}_{n-1}+1\big)+\frac{n+\kappa+1}{z}=0,$
${Y}_{n}\big({Z}_{n}+{X}_{n}+1\big)+\frac{n+m+\kappa+\tfrac{3}{2}}{z}=0,$
$Z_{n}\big({X}_{n+1}+{Y}_{n}+1\big)+\frac{n+\tfrac{3}{2}}{z}=0.$
More generally, using the solution $\widehat{w}_{m,n}^{(\kappa)}$, for fixed $m$ and $n$, then we have the following Lemma.
Lemma 5.28.
Consider the rational functions
${X}_{N}(z)=\frac{1}{\widehat{w}_{N+m,n}^{(\kappa)}(z)-1},\qquad{Y}_{N}(z)=\frac{1}{\widehat{v}_{N+m,n}^{(\kappa+n+1/2)}(z)-1},\qquad{Z}_{N}(z)=\frac{1}{\widehat{u}_{N+m,n}^{(2\kappa+1)}(z)-1},$ (5.28)
with $\widehat{w}_{m,n}^{(\kappa)}(z)$, $\widehat{v}_{m,n}^{(\kappa)}(z)$ and $\widehat{u}_{m,n}^{(\kappa)}(z)$ as given in Theorem 4.9, then ${X}_{N}(z)$, ${Y}_{N}(z)$ and ${Z}_{N}(z)$ satisfy
${X}_{N}\big({Y}_{N}+{Z}_{N-1}+1\big)+\frac{N+m+\kappa}{z}=0,$ (5.29a)
${Y}_{N}\big({Z}_{N}+{X}_{N}+1\big)+\frac{N+m+n+\kappa+\tfrac{1}{2}}{z}=0,$ (5.29b)
$Z_{N}\big({X}_{N+1}+{Y}_{N}+1\big)+\frac{N+m+\tfrac{1}{2}}{z}=0,$ (5.29c)
which is the system (3.28) with $\lambda=\tfrac{1}{3}(\kappa-n)$, $\varphi=n+\kappa+\tfrac{1}{3}$ and $\rho=m+\tfrac{1}{3}n+\tfrac{2}{3}\kappa$.
Proof.
From the transformations
$\mathcal{R}_{3}\big(\widehat{w}_{m,n}^{(\kappa)};m+\kappa,-n-\kappa,m+n\big)$ $=\big(\widehat{v}_{m,n}^{(\kappa+n+1/2)};m+n+\kappa+\tfrac{1}{2},\kappa+\tfrac{1}{2},m-n\big),$
$\mathcal{R}_{3}\big(\widehat{v}_{m,n}^{(\kappa)};m+\kappa,-n+\kappa,m-n\big)$ $=\big(\widehat{u}_{m,n}^{(2\kappa-2n)};m+n+\tfrac{1}{2},\tfrac{1}{2},m-n+2\kappa\big),$
$\mathcal{R}_{3}\big(\widehat{u}_{m,n}^{(2\kappa)};m+\tfrac{1}{2},n+\tfrac{1}{2},m+n+2\kappa\big)$ $=\big(\widehat{w}_{m+1,n}^{(\kappa-1/2)};m+\kappa+\tfrac{1}{2},-n-\kappa+\tfrac{1}{2},m+n+1\big),$
we obtain the sequence of solutions
$\widehat{w}_{m,n}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{v}_{m,n}^{(\kappa+n+1/2)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{m,n}^{(2\kappa+1)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{m+1,n}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{v}_{m+1,n}^{(\kappa+n+1/2)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{m+1,n}^{(2\kappa+1)}$
from which the result follows. We note that
$\mathcal{R}_{3}^{3}\Big(\widehat{w}_{m,n}^{(\kappa)}\Big)=\widehat{w}_{m+1,n}^{(\kappa)},\qquad\mathcal{R}_{3}^{3}\Big(\widehat{v}_{m,n}^{(\kappa)}\Big)=\widehat{v}_{m+1,n}^{(\kappa)},\qquad\mathcal{R}_{3}^{3}\Big(\widehat{u}_{m,n}^{(2\kappa+1)}\Big)=\widehat{u}_{m+1,n}^{(2\kappa+1)}.$
Remark 5.29.
It is interesting to note that ${X}_{N}(z)$, ${Y}_{N}(z)$ and $Z_{N}(z)$ arise from the three different sets of rational solutions of PV that are expressed in terms of the generalised Umemura polynomials.
As was the case for $\mathcal{R}_{2}$, it is also possible to obtain another sequence of solutions as illustrated in the following Lemma.
Lemma 5.30.
Consider the rational functions
${X}_{N}(z)=\frac{1}{\widehat{w}_{N+m,n}^{(\kappa)}(z)-1},\qquad{Y}_{N}(z)=\frac{1}{\widehat{u}_{N+m,n-1}^{(2\kappa+1)}(z)-1},\qquad{Z}_{N}(z)=\frac{1}{\widehat{v}_{N+m+1,n}^{(\kappa+n-1/2)}(z)-1},$ (5.30)
with $\widehat{w}_{m,n}^{(\kappa)}(z)$, $\widehat{v}_{m,n}^{(\kappa)}(z)$ and $\widehat{u}_{m,n}^{(\kappa)}(z)$ as given in Theorem 4.9, then ${X}_{N}(z)$, ${Y}_{N}(z)$ and ${Z}_{N}(z)$ satisfy
${X}_{N}\big({Y}_{N}+{Z}_{N-1}+1\big)+\frac{N+m+\kappa}{z}=0,$ (5.31a)
${Y}_{N}\big({Z}_{N}+{X}_{N}+1\big)+\frac{N+m+\tfrac{1}{2}}{z}=0,$ (5.31b)
${Z}_{N}\big({X}_{N+1}+{Y}_{N}+1\big)+\frac{N+m+n+\kappa+\tfrac{1}{2}}{z}=0,$ (5.31c)
which is the system (3.28) with $\lambda=\tfrac{1}{3}(\kappa-n)$, $\varphi=-n-\kappa+\tfrac{1}{3}$ and $\rho=m+\tfrac{1}{3}n+\tfrac{2}{3}\kappa$.
Proof.
From the transformations
$\mathcal{R}_{3}\big(\widehat{w}_{m,n}^{(\kappa)};m+\kappa,n+\kappa,m+n\big)$ $=\big(\widehat{u}_{m,n-1}^{(2\kappa+1)};m+\tfrac{1}{2},-n+\tfrac{1}{2},m+n+2\kappa\big),$
$\mathcal{R}_{3}\big(\widehat{u}_{m,n}^{(2\kappa)};m+\tfrac{1}{2},-n-\tfrac{1}{2},m+n+2\kappa\big)$ $=\big(\widehat{v}_{m+1,n+1}^{(\kappa+n)};m+n+\kappa+1,-\kappa+1,m-n\big),$
$\mathcal{R}_{3}\big(\widehat{v}_{m,n}^{(\kappa)};m+\kappa,n-\kappa,m+n\big)$ $=\big(\widehat{w}_{m,n}^{(\kappa-n+1/2)};m+\kappa+\tfrac{1}{2},-n+\kappa+\tfrac{1}{2},m+n\big),$
we obtain the sequence of solutions
$\widehat{w}_{m,n}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{m,n-1}^{(2\kappa+1)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{v}_{m,n}^{(\kappa+n-1/2)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{w}_{m+1,n}^{(\kappa)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\widehat{u}_{m+1,n-1}^{(2\kappa+1)}$
from which the result follows. We note that
$\mathcal{R}_{3}^{3}\Big(\widehat{w}_{m,n}^{(\kappa)}\Big)=\widehat{w}_{m+1,n}^{(\kappa)},\qquad\mathcal{R}_{3}^{3}\Big(\widehat{v}_{m,n}^{(\kappa)}\Big)=\widehat{v}_{m+1,n}^{(\kappa)},\qquad\mathcal{R}_{3}^{3}\Big(\widehat{u}_{m,n}^{(2\kappa+1)}\Big)=\widehat{u}_{m+1,n}^{(2\kappa+1)}.$
Remark 5.31.
There are no solutions in terms of generalised Umemura polynomials in the symmetric case when $\lambda=\varphi=0$.
5.2.4 Partial difference systems
In this subsection we derive some partial difference systems which have solutions in terms of generalised Umemura polynomials.
Lemma 5.32.
If we define
$\widehat{\mathcal{W}}_{m,n}^{(\kappa)}(z)=\frac{\widehat{w}_{m,n}^{(\kappa)}(z)+1}{\widehat{w}_{m,n}^{(\kappa)}(z)-1},\qquad\widehat{\mathcal{V}}_{m,n}^{(\kappa)}(z)=\frac{\widehat{v}_{m,n}^{(\kappa)}(z)+1}{\widehat{v}_{m,n}^{(\kappa)}(z)-1},\qquad\widehat{\mathcal{U}}_{m,n}^{(\kappa)}(z)=\frac{\widehat{u}_{m,n}^{(\kappa)}(z)+1}{\widehat{u}_{m,n}^{(\kappa)}(z)-1},$ (5.32)
with $\widehat{w}_{m,n}^{(\kappa)}(z)$, $\widehat{v}_{m,n}^{(\kappa)}(z)$ and $\widehat{u}_{m,n}^{(\kappa)}(z)$ as given in Theorem 4.9, then they satisfy the partial difference systems
$\widehat{\mathcal{W}}_{m,n}^{(\kappa+1)}+\widehat{\mathcal{W}}_{m,n}^{(\kappa)}$ $=\frac{4}{z}\frac{(m+n+2\kappa+1)\widehat{\mathcal{V}}_{m,n}^{(\kappa+n+1/2)}+m+n}{1-\Big(\widehat{\mathcal{V}}_{m,n}^{(\kappa+n+1/2)}\Big)^{\!2}},$ (5.33a)
$\widehat{\mathcal{V}}_{m,n}^{(\kappa+n+1/2)}+\widehat{\mathcal{V}}_{m,n}^{(\kappa+n-1/2)}$ $=\frac{4}{z}\frac{(m+n+2\kappa)\widehat{\mathcal{W}}_{m,n}^{(\kappa)}+m-n}{1-\Big(\widehat{\mathcal{W}}_{m,n}^{(\kappa)}\Big)^{\!2}},$ (5.33b)
and
$\widehat{\mathcal{V}}_{m,n}^{(\kappa+n)}+\widehat{\mathcal{V}}_{m+1,n+1}^{(\kappa+n)}$ $=\frac{4}{z}\frac{(m+n+1)\,\widehat{\mathcal{U}}_{m,n}^{(2\kappa)}+m-n}{1-\Big(\widehat{\mathcal{U}}_{m,n}^{(2\kappa)}\Big)^{\!2}},$ (5.34a)
$\widehat{\mathcal{U}}_{m,n}^{(2\kappa)}+\widehat{\mathcal{U}}_{m+1,n+1}^{(2\kappa-2)}$ $=\frac{4}{z}\frac{(m+n+2)\widehat{\mathcal{V}}_{m+1,n+1}^{(\kappa+n)}+m+n+2\kappa}{1-\Big(\widehat{\mathcal{V}}_{m+1,n+1}^{(\kappa+n)}\Big)^{\!2}}.$ (5.34b)
Proof.
These follow from setting $N=0$ in (5.22) and (5.23), respectively. ∎
Lemma 5.33.
If $\widehat{\mathcal{W}}_{m,n}^{(\kappa)}(z)$ and $\widehat{\mathcal{U}}_{m,n}^{(\kappa)}(z)$ are given by (5.32),
$\widehat{\mathcal{W}}_{m+1,n}^{(\kappa)}+\widehat{\mathcal{W}}_{m,n+1}^{(\kappa)}$ $=\frac{4}{z}\frac{(m-n)\,\widehat{\mathcal{U}}_{m,n}^{(2\kappa+1)}+m+n+1}{1-\Big(\widehat{\mathcal{U}}_{m,n}^{(2\kappa+1)}\Big)^{\!2}},$ (5.35a)
$\widehat{\mathcal{U}}_{m-1,n}^{(2\kappa+1)}+\widehat{\mathcal{U}}_{m,n-1}^{(2\kappa+1)}$ $=\frac{4}{z}\frac{(m-n)\widehat{\mathcal{W}}_{m,n}^{(\kappa)}+m+n+2\kappa}{1-\Big(\widehat{\mathcal{W}}_{m,n}^{(\kappa)}\Big)^{\!2}}.$ (5.35b)
Proof.
From the transformations (5.24) we obtain the sequences
$\begin{array}[]{ccccc}\widehat{w}_{m,n+1}^{(\kappa)}&\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip&\widehat{u}_{m,n}^{(2\kappa+1)}&\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip&\widehat{w}_{m+1,n}^{(\kappa)}\\ \widehat{u}_{m-1,n}^{(2\kappa+1)}&\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip&\widehat{w}_{m,n}^{(\kappa)}&\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip&\widehat{u}_{m,n-1}^{(2\kappa+1)}\end{array}$
from which the result follows. We note that
$\mathcal{R}_{3}^{3}\Big(\widehat{w}_{m,n}^{(\kappa)}\Big)=\widehat{w}_{m+1,n}^{(\kappa)},\qquad\mathcal{R}_{3}^{3}\Big(\widehat{v}_{m,n}^{(\kappa)}\Big)=\widehat{v}_{m+1,n}^{(\kappa)},\qquad\mathcal{R}_{3}^{3}\Big(\widehat{u}_{m,n}^{(2\kappa+1)}\Big)=\widehat{u}_{m+1,n}^{(2\kappa+1)}.$
Lemma 5.34.
If $\widehat{\mathcal{W}}_{m,n}^{(\kappa)}(z)$, $\widehat{\mathcal{V}}_{m,n}^{(\kappa)}(z)$ and $\widehat{\mathcal{U}}_{m,n}^{(\kappa)}(z)$ are given by (5.32), then they satisfy the partial difference systems
$\frac{2m+1}{\widehat{\mathcal{W}}_{m+1,n}^{(\kappa-1)}+\widehat{\mathcal{V}}_{m,n}^{(\kappa+n-1/2)}}+\frac{2m-1}{\widehat{\mathcal{W}}_{m-1,n}^{(\kappa)}+\widehat{\mathcal{V}}_{m,n}^{(\kappa+n-1/2)}}$ $=z-\frac{2\Big[(m+n)\widehat{\mathcal{V}}_{m,n}^{(\kappa+n-1/2)}+m+n+2\kappa-1\Big]}{1-\Big(\widehat{\mathcal{V}}_{m,n}^{(\kappa+n-1/2)}\Big)^{\!2}},$
$\frac{2m+1}{\widehat{\mathcal{W}}_{m,n}^{(\kappa)}+\widehat{\mathcal{V}}_{m+1,n}^{(\kappa+n-1/2)}}+\frac{2m-1}{\widehat{\mathcal{W}}_{m,n}^{(\kappa)}+\widehat{\mathcal{V}}_{m-1,n}^{(\kappa+n+1/2)}}$ $=z-\frac{2\Big[(m-n)\widehat{\mathcal{W}}_{m,n}^{(\kappa)}+m+n+2\kappa\Big]}{1-\Big(\widehat{\mathcal{W}}_{m,n}^{(\kappa)}\Big)^{\!2}},$
and
$\frac{2m+2n+2\kappa+3}{\widehat{\mathcal{W}}_{m+1,n+1}^{(\kappa)}+\widehat{\mathcal{U}}_{m,n}^{(2\kappa+1)}}+\frac{2m+2n+2\kappa+1}{\widehat{\mathcal{W}}_{m,n}^{(\kappa)}+\widehat{\mathcal{U}}_{m,n}^{(2\kappa+1)}}$ $=z-\frac{2\Big[(m+n+1)\widehat{\mathcal{U}}_{m,n}^{(2\kappa+1)}+m-n\Big]}{1-\Big(\widehat{\mathcal{U}}_{m,n}^{(2\kappa+1)}\Big)^{\!2}},$
$\frac{2m+2n+2\kappa+1}{\widehat{\mathcal{W}}_{m,n}^{(\kappa)}+\widehat{\mathcal{U}}_{m,n}^{(2\kappa+1)}}+\frac{2m+2n+2\kappa-1}{\widehat{\mathcal{W}}_{m,n}^{(\kappa)}+\widehat{\mathcal{U}}_{m-1,n-1}^{(2\kappa+1)}}$ $=z-\frac{2\Big[(m+n+2\kappa)\widehat{\mathcal{W}}_{m,n}^{(\kappa)}+m-n\Big]}{1-\Big(\widehat{\mathcal{W}}_{m,n}^{(\kappa)}\Big)^{\!2}}.$
Proof.
These follow from setting $N=0$ in (5.26) and (5.27), respectively. ∎
Lemma 5.35.
If we define
$\widehat{\mathcal{X}}_{m,n}^{(\kappa)}(z)=\frac{1}{\widehat{w}_{m,n}^{(\kappa)}(z)-1},\qquad\widehat{\mathcal{Y}}_{m,n}^{(\kappa)}(z)=\frac{1}{\widehat{v}_{m,n}^{(\kappa)}(z)-1},\qquad\widehat{\mathcal{Z}}_{m,n}^{(\kappa)}(z)=\frac{1}{\widehat{u}_{m,n}^{(\kappa)}(z)-1},$ (5.36)
with $\widehat{w}_{m,n}^{(\kappa)}(z)$, $\widehat{v}_{m,n}^{(\kappa)}(z)$ and $\widehat{u}_{m,n}^{(\kappa)}(z)$ as given in Theorem 4.9, then they satisfy the partial difference system
$\widehat{\mathcal{X}}_{m,n}^{(\kappa)}\Big(\widehat{\mathcal{Y}}_{m,n}^{(\kappa+n+1/2)}+\widehat{\mathcal{Z}}_{m-1,n}^{(2\kappa+1)}+1\Big)+\frac{m+\kappa}{z}=0,$ (5.37a)
$\widehat{\mathcal{Y}}_{m,n}^{(\kappa+n+1/2)}\Big(\widehat{\mathcal{Z}}_{m,n}^{(2\kappa+1)}+\widehat{\mathcal{X}}_{m,n}^{(\kappa)}+1\Big)+\frac{m+n+\kappa+\tfrac{1}{2}}{z}=0,$ (5.37b)
$\widehat{\mathcal{Z}}_{m,n}^{(2\kappa+1)}\Big(\widehat{\mathcal{X}}_{m+1,n}^{(\kappa)}+\widehat{\mathcal{Y}}_{m,n}^{(\kappa+n+1/2)}+1\Big)+\frac{m+\tfrac{1}{2}}{z}=0.$ (5.37c)
Proof.
This follows from setting $N=0$ in (5.28) and (5.29). ∎
5.3 Non-unique solutions of the discrete equations
Kitaev, Law and McLeod [40, Theorem 1.2] prove that if $\gamma\not\in\mathbb{Z}$ in case (i), then the rational solution of PV is unique and if $\gamma\in\mathbb{Z}$ then there are at most two rational solutions. Clarkson and Dunning [12] gave examples of non-unique rational solutions which arise for solutions expressed in terms of generalised Laguerre polynomials discussed in §5.1. For example, the rational solutions
$w_{1,1}^{(1)}(z)$ $=-\frac{(z-3)(z^{2}-8z+20)}{(z-2)(z-6)},\qquad u_{1,2}^{(1)}(z)=\frac{(z^{2}+4z+6)(z^{3}+9z^{2}+36z+60)}{z^{4}+12z^{3}+54z^{2}+96z+72},$ (5.38)
both satisfy PV (1.2) for the parameters $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2},-\tfrac{25}{2},1\big)$. Hence the solutions
$q(z)$ $=\frac{w_{1,1}^{(1)}(z)+1}{w_{1,1}^{(1)}(z)-1}=\frac{z^{3}-12z^{2}+52z-72}{(z-4)(z^{2}-6z+12)},$ (5.39a)
$p(z)$ $=\frac{u_{1,2}^{(1)}(z)+1}{u_{1,2}^{(1)}(z)-1}=\frac{z^{5}+14z^{4}+90z^{3}+312z^{2}+552z+432}{(z^{2}+6z+12)(z^{3}+6z^{2}+18z+24)},$ (5.39b)
both satisfy equation (4.10) for the parameters $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2},-\tfrac{25}{2},1\big)$. In this subsection, we use such pairs of solutions to derive examples of distinct hierarchies of rational solutions which satisfy the same discrete equation.
5.3.1 First discrete equation: asymmetric dPII
Applying the Bäcklund transformation $\mathcal{R}_{1}$ (3.3), with (a,b,c)=(1,5,1)(a,b,c)=(1,-5,1), to the solutions (5.38) gives the sequences of rational solutions
$w_{1,1}^{(1)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\frac{1}{{v}_{1,1}^{(4)}}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip w_{2,1}^{(1)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\frac{1}{{v}_{2,1}^{(4)}}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip w_{3,1}^{(1)}$
$u_{1,2}^{(1)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip{v}_{2,1}^{(4)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{2,2}^{(1)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip v_{2,2}^{(4)}\enskip\begin{array}[]{c}\mathcal{R}_{1}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{3,2}^{(1)}$
Hence we obtain the hierarchies
$q_{2n}(z)$ $=\frac{1+v_{n,1}^{(4)}(z)}{1-v_{n,1}^{(4)}(z)}=1-\frac{8}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{n,1}^{(1)}(z)}{T_{n-1,2}^{(1)}(z)},$ (5.40a)
$q_{2n+1}(z)$ $=\frac{w_{n+1,1}^{(1)}(z)+1}{w_{n+1,1}^{(1)}(z)-1}=1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{n,1}^{(2)}(z)}{T_{n,2}^{(0)}(z)},$ (5.40b)
and
$p_{2n}(z)$ $=\frac{v_{2,n}^{(-4)}(z)+1}{v_{2,n}^{(-4)}(z)-1}=-1-\frac{8}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{1,n+1}^{(-2n-5)}(z)}{T_{2,n}^{(-2n-5)}(z)},$ (5.41a)
$p_{2n+1}(z)$ $=\frac{u_{n+1,2}^{(1)}(z)+1}{u_{n+1,2}^{(1)}(z)-1}=1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{2,n+1}^{(-2n-6)}(z)}{T_{1,n+1}^{(-2n-6)}(z)},$ (5.41b)
where $q_{2n}$ and $p_{2n}$ both satisfy equation (4.10) for the parameters $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(n+3)^{2},-\tfrac{1}{2}(n+2)^{2},-4\big)$ whilst $q_{2n+1}$ and $p_{2n+1}$ both satisfy equation (4.10) for the parameters $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(n+1)^{2},-\tfrac{1}{2}(n+5)^{2},1\big)$. We note that there is a difference in the structure of the solutions in the hierarchies (5.40) and (5.41).
Both $q_{n}$ and $p_{n}$ satisfy the discrete equation
$q_{n+1}+q_{n-1}=\frac{4}{z}\frac{(n+5)q_{n}-\tfrac{3}{2}+\frac{5}{2}(-1)^{n}}{1-q_{n}^{2}},$ (5.42)
which is asymmetric dPII (3.7) with $\lambda=5$, $\rho=-\tfrac{3}{2}$ and $\varphi=\tfrac{5}{2}$. Further, if either $(x_{n},y_{n})=(q_{2n},q_{2n+1})$ or $(x_{n},y_{n})=(p_{2n},p_{2n+1})$ then we obtain solutions to the discrete system
$x_{n+1}+x_{n}$ $=\frac{4}{z}\frac{(2n+6)y_{n}-4}{1-y_{n}^{2}},\qquad y_{n}+y_{n-1}=\frac{4}{z}\frac{(2n+5)x_{n}+1}{1-x_{n}^{2}}.$
The first few solutions in the hierarchies (5.40) and (5.41) are given by
$q_{0}$ $=1-\frac{8}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln(z-3),$
$q_{2}$ $=1-\frac{8}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{2}-8z+12}{z^{2}-8z+20},$
$q_{4}$ $=1-\frac{8}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{3}-15z^{2}+60z-60}{z^{4}-20z^{3}+150z^{2}-480z+600},$
$q_{6}$ $=1-\frac{8}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{4}-24z^{3}+180z^{2}-480z+360}{z^{6}-36z^{5}+522z^{4}-3840z^{3}+15120z^{2}-30240z+25200},$
$q_{1}$ $=1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z-4}{z^{2}-6z+12},$
$q_{3}$ $=1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{2}-10z+20}{z^{4}-16z^{3}+96z^{2}-240z+240},$
$q_{5}$ $=1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{3}-18z^{2}+90z-120}{z^{6}-30z^{5}+360z^{4}-2160z^{3}+6840z^{2}-10800z+7200},$
$q_{7}$ $=1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{4}-28z^{3}+252z^{2}-840z+840}{z^{8}-48z^{7}+960z^{6}-10320z^{5}+64800z^{4}-241920z^{3}+524160z^{2}-604800z+302400},$
and
$p_{0}$ $=-1-\frac{8}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln(z^{2}+4z+6),$
$p_{2}$ $=-1-\frac{8}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{4}+12z^{3}+54z^{2}+96z+72}{z^{3}+9z^{2}+36z+60},$
$p_{4}$ $=-1-\frac{8}{z}+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{6}+24z^{5}+228z^{4}+1056z^{3}+2520z^{2}+2880z+1440}{z^{6}+24z^{5}+252z^{4}+1440z^{3}+4680z^{2}+8640z+7200},$
$p_{1}$ $=1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{3}+6z^{2}+18z+24}{z^{2}+6z+12},$
$p_{3}$ $=1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{6}+18z^{5}+144z^{4}+624z^{3}+1512z^{2}+2160z+1440}{z^{4}+16z^{3}+96z^{2}+240z+240},$
$p_{5}$ $=1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{9}+36z^{8}+576z^{7}+5280z^{6}+30240z^{5}+112320z^{4}+270720z^{3}+414720z^{2}+388800z+172800}{z^{6}+30z^{5}+360z^{4}+2160z^{3}+6840z^{2}+10800z+7200},$
with $q_{-1}(z)=p_{-1}(z)=1$.
5.3.2 Second discrete equation
The rational solutions
$v_{0,n}^{(m+n+1)}(z)=(-1)^{n}\frac{m+1}{n+1}\,\frac{T_{0,n}^{(m-n+1)}(z)}{T_{0,n+1}^{(m-n-1)}(z)},\qquad\frac{1}{v_{m,0}^{(-m-n-1)}}(z)=-\frac{m+1}{n+1}\,\frac{T_{m,1}^{(-m-n-3)}(z)}{T_{m-1,1}^{(-m-n-3)}(z)},$
both satisfy PV (1.2) for the parameters $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(n+1)^{2},-\tfrac{1}{2}(m+1)^{2},m+n+1\big)$. In the case when m=n1m=n-1, applying the Bäcklund transformation $\mathcal{R}_{2}$ (3.12) gives the hierarchy of solutions
$q_{2n}(z)$ $=\frac{1+v_{n,0}^{(-2n)}(z)}{1-v_{n,0}^{(-2n)}(z)}=1+\frac{4n}{z}-2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln T_{n-1,1}^{(-2n-1)}(z),$ (5.43a)
$q_{2n+1}(z)$ $=\frac{1+\widetilde{v}_{n,0}^{(-2n)}(z)}{1-\widetilde{v}_{n,0}^{(-2n)}(z)}=-1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln T_{n,1}^{(-2n-3)}(z),$ (5.43b)
and
$p_{2n}(z)$ $=\frac{v_{0,n-1}^{(2n)}(z)+1}{v_{0,n-1}^{(2n)}(z)-1}=-1+\frac{4n}{z}-2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln T_{0,n-1}^{(1)}(z)$ (5.44a)
$p_{2n+1}(z)$ $=\frac{\widetilde{v}_{0,n-1}^{(2n)}(z)-1}{\widetilde{v}_{0,n-1}^{(2n)}(z)+1}=1+2\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln T_{0,n}^{(1)}(z),$ (5.44b)
where
${v}_{n,0}^{(\mu)}(z)$ $=\frac{n+\mu}{n+1}\frac{T_{n-1,1}^{(\mu-2)}(z)}{T_{n,1}^{(\mu-2)}(z)}=\frac{m+\mu}{m+1}\,\frac{L^{(\mu-1)}_{m}(z)}{L^{(\mu-1)}_{m+1}(z)},$
$\widetilde{v}_{n,0}^{(\mu)}(z)$ $=\frac{T_{n,1}^{(\mu-2)}(z)}{T_{n-1,1}^{(\mu-2)}(z)}=\frac{L^{(\mu-1)}_{n+1}(z)}{L^{(\mu-1)}_{n}(z)},$
${v}_{0,n}^{(\mu)}(z)$ $=(-1)^{n}\frac{\mu-n}{n+1}\,\frac{T_{0,n}^{(\mu-2n)}(z)}{T_{0,n+1}^{(\mu-2n-2)}(z)}={\frac{n-\mu}{n+1}\,\frac{L_{n}^{(-\mu-1)}(-z)}{L_{n+1}^{(-\mu-1)}(-z)}},$
$\widetilde{v}_{0,n}^{(\mu)}(z)$ $=(-1)^{n+1}\frac{T_{0,n+1}^{(\mu-2n-2)}(z)}{T_{0,n}^{(\mu-2n)}(z)}=\frac{L_{n+1}^{(-1-\mu)}(-z)}{L_{n}^{(-1-\mu)}(-z)}.$
The solutions $q_{2n}$ and $p_{2n}$ both satisfy equation (4.10) for the parameters $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}n^{2},-\tfrac{1}{2}(n+1)^{2},2n\big)$ whilst $q_{2n+1}$ and $p_{2n+1}$ both satisfy equation (4.10) for the parameters $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}n^{2},-\tfrac{1}{2}(n+1)^{2},2n+2\big)$. Again there is a difference in the structure of the solutions in the hierarchies (5.43) and (5.44).
Both $q_{n}$ and $p_{n}$ satisfy the discrete equation
$\frac{2n+2}{q_{n+1}+q_{n}}+\frac{2n}{q_{n}+q_{n-1}}=z-\frac{\big[2n+1+(-1)^{n}\big]q_{n}-2}{1-q_{n}^{2}},$
which is equation (3.17) with $\lambda=\tfrac{1}{2}$, $\rho=-2$ and $\varphi=\tfrac{1}{2}$. Further, if either $(x_{n},y_{n})=(q_{2n},q_{2n+1})$ or $(x_{n},y_{n})=(p_{2n},p_{2n+1})$ then we obtain solutions of the discrete system
$\frac{2n+2}{x_{n+1}+y_{n}}+\frac{2n+1}{x_{n}+y_{n}}$ $=\tfrac{1}{2}z-\frac{(2n+1)y_{n}-1}{1-y_{n}^{2}},$
$\frac{2n+1}{x_{n}+y_{n}}+\frac{2n}{x_{n}+y_{n-1}}$ $=\tfrac{1}{2}z-\frac{(2n+1)x_{n}-1}{1-x_{n}^{2}}.$
The first few solutions in the hierarchies (5.43) and (5.44) are given by
$\begin{split}q_{2}&=1+\frac{4}{z}-\frac{2}{z+1},\\ q_{4}&=1+\frac{8}{z}-\frac{4(2+z)}{z^{2}+4z+6},\\ q_{6}&=1+\frac{12}{z}-\frac{6(z^{2}+6z+12)}{z^{3}+9z^{2}+36z+60},\\ q_{8}&=1+\frac{16}{z}-\frac{8(z^{3}+12z^{2}+60z+120)}{z^{4}+16z^{3}+120z^{2}+480z+840},\end{split}\qquad\begin{split}q_{1}&=-1+\frac{2}{z+1},\\ q_{3}&=-1+\frac{4(z+2)}{z^{2}+4z+6},\\ q_{5}&=-1+\frac{6(z^{2}+6z+12)}{z^{3}+9z^{2}+36z+60},\\ q_{7}&=-1+\frac{8(z^{3}+12z^{2}+60z+120)}{z^{4}+16z^{3}+120z^{2}+480z+840},\end{split}$
and
$\begin{split}p_{2}&=-1+\frac{4}{z},\\ p_{4}&=-1+\frac{8}{z}-\frac{2}{z-3},\\ p_{6}&=-1+\frac{12}{z}-\frac{4(z-4)}{z^{2}-8z+20},\\ p_{8}&=-1+\frac{16}{z}-\frac{6(z^{2}-10z+30)}{z^{3}-15z^{2}+90z-210},\end{split}\qquad\begin{split}p_{1}&=1,\\ p_{3}&=1+\frac{2}{z-3},\\ p_{5}&=1+\frac{4(z-4)}{z^{2}-8z+20},\\ p_{7}&=1+\frac{6(z^{2}-10z+30)}{z^{3}-15z^{2}+90z-210}.\end{split}$
5.3.3 Third discrete equation: ternary dPI
The rational solutions
$v_{3,1}^{(-1)}(z)=-\frac{(z-2)(z-6)}{(z-3)(z^{2}-8z+20)},\qquad v_{2,2}^{(-1)}(z)=-\frac{z^{4}+12z^{3}+54z^{2}+96z+72}{(z^{2}+4z+6)(z^{3}+9z^{2}+36z+60)},$ (5.45)
both satisfy PV (1.2) for the parameters $(\alpha,\beta,\gamma)=\big(\tfrac{25}{2},-\tfrac{1}{2},-1\big)$ and so
$x(z)$ $=\frac{1}{v_{3,1}^{(-1)}(z)-1}=-\frac{(z-3)(z^{2}-8z+20)}{(z-4)(z^{2}-6z+12)},$ (5.46a)
$y(z)$ $=\frac{1}{v_{2,2}^{(-1)}(z)-1}=-\displaystyle\frac{(z^{2}+4z+6)(z^{3}+9z^{2}+36z+60)}{(z^{2}+6z+12)(z^{3}+6z^{2}+18z+24)},$ (5.46b)
both satisfy equation (4.11) for the parameters $(\alpha,\beta,\gamma)=\big(\tfrac{25}{2},-\tfrac{1}{2},-1\big)$. We note that
$v_{3,1}^{(-1)}(z)={1/w_{1,1}^{(1)}(z)},\qquad v_{2,2}^{(-1)}(z)={1/u_{1,2}^{(1)}(z)}.$
Applying the Bäcklund transformation $\mathcal{R}_{3}$ (3.23), with (a,b,c)=(5,1,1)(a,b,c)=(5,1,-1), to the solutions (5.45) gives the sequences of rational solutions
${v}_{3,1}^{(-1)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{2,3}^{(6)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\frac{1}{{w}_{3,2}^{(-5)}}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip{v}_{3,2}^{(0)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{3,3}^{(7)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\frac{1}{{w}_{3,3}^{(-6)}}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip{v}_{3,3}^{(1)}$
${v}_{2,2}^{(-1)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\frac{1}{{w}_{3,2}^{(-6)}}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{3,2}^{(5)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip{v}_{2,3}^{(0)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip\frac{1}{{w}_{3,3}^{(-7)}}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip u_{4,2}^{(6)}\enskip\begin{array}[]{c}\mathcal{R}_{3}\\[-8.0pt] -\!\!\!-\!\!\!-\!\!\!\longrightarrow\end{array}\enskip{v}_{2,4}^{(1)}$
Hence we obtain the solution hierarchies
$x_{3n}(z)=-1+\frac{n+3}{z}-\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln T_{0,n}^{(2-n)}(z)=-1+\frac{n+3}{z}-\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln L_{n}^{(-n-3)}(-z),$ (5.47a)
$x_{3n+1}(z)=-1+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{0,n+1}^{(1-n)}(z)}{T_{0,n}^{(3-n)}(z)}=-1+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{L_{n+1}^{(-n-2)}(-z)}{L_{n}^{(-n-2)}(-z)},$ (5.47b)
$x_{3n+2}(z)=\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln T_{0,n+1}^{(2-n)}(z)=-1+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{L_{n+1}^{(-n-2)}(-z)}{L_{n}^{(-n-2)}(-z)},$ (5.47c)
and
$y_{3n}(z)=\frac{n+3}{z}-\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln T_{1,1}^{(-n-4)}(z)=\frac{n+3}{z}-\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln L_{2}^{(-n-3)}(z),$ (5.48a)
$y_{3n+1}(z)=-1+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{T_{1,1}^{(-n-5)}(z)}{T_{2,1}^{(-n-5)}(z)}=-1+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{L_{2}^{(-n-4)}(z)}{L_{3}^{(-n-4)}(z)},$ (5.48b)
$y_{3n+2}(z)=-1+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln T_{2,1}^{(-n-6)}(z)=-1+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln L_{3}^{(-n-5)}(z),$ (5.48c)
where $x_{n}$ and $y_{n}$ both satisfy equation (4.11) for the parameters $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(n+1)^{2},-\tfrac{9}{2},n+3\big)$, $x_{3n+1}$ and $y_{3n+1}$ for the parameters $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(n+4)^{2},-\tfrac{1}{2},n-2\big)$ and $x_{3n+2}$ and $y_{3n+2}$ for the parameters $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(n+1)^{2},-\tfrac{9}{2},n+5\big)$. We remark that whilst $x_{n}$ is only defined for $n\geq-1$, $y_{n}$ is defined for all $n\in\mathbb{Z}$. Also the solutions $x_{n}$ only involve the generalised Laguerre polynomials $T_{0,n}^{(\mu)}(z)=(-1)^{\lceil n/2\rceil}L_{n}^{(-\mu-2n-1)}(-z)$, whereas the solutions $y_{n}$ only involve the generalised Laguerre polynomials $T_{1,1}^{(\mu)}(z)=L_{2}^{(\mu+1)}(z)$ and $T_{2,1}^{(\mu)}(z)=L_{3}^{(\mu+1)}(z)$. Both $x_{n}$ and $y_{n}$ satisfy the ternary dPI equation
$x_{n}(x_{n+1}+x_{n-1}+1)+\frac{a_{n}}{z}=0,$
and the equation
$\frac{a_{n+1}}{x_{n+2}+x_{n}+1}+\frac{a_{n-1}}{x_{n}+x_{n-2}+1}=z+\frac{a_{n}}{x_{n}},$
with
$a_{n}=\tfrac{1}{3}n+\tfrac{5}{3}-\tfrac{2}{3}\cos(\tfrac{2}{3}\pi n)+\tfrac{10}{9}\sqrt{3}\,\sin(\tfrac{2}{3}\pi n).$
Further if $(X_{n},Y_{n},Z_{n})=(x_{3n},x_{3n+1},x_{3n+2})$ or $(X_{n},Y_{n},Z_{n})=(y_{3n},y_{3n+1},y_{3n+2})$, then we obtain solutions of the discrete system
$X_{n}\big(Y_{n}+Z_{n-1}+1\big)+\frac{n+1}{z}=0,\quad Y_{n}\big(Z_{n}+X_{n}+1\big)+\frac{n+4}{z}=0,\quad Z_{n}\big(X_{n+1}+Y_{n}+1\big)+\frac{n+1}{z}=0.$
The first few solutions in the hierarchies (5.47) and (5.48) are given by
$\begin{split}x_{0}&=-1+\frac{3}{z},\\ x_{3}&=-1+\frac{4}{z}-\frac{1}{z-3},\\ x_{6}&=-1+\frac{5}{z}-\frac{2(z-3)}{z^{2}-6z+12},\end{split}\qquad\begin{split}x_{1}&=-1+\frac{1}{z-3},\\ x_{4}&=-1+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{2}-6z+12}{z-4},\\ x_{7}&=-1+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{3}-9z^{2}+36z-60}{z^{2}-8z+20},\end{split}\qquad\begin{split}x_{2}&=\frac{1}{z-4},\\ x_{5}&=\frac{2(z-4)}{z^{2}-8z+20},\\ x_{8}&=\frac{3(z^{2}-8z+20)}{z^{3}-12z^{2}+60z-120},\end{split}$
and
$\begin{split}y_{0}&=\frac{3}{z}-\frac{2(z+1)}{z^{2}+2z+2},\\ y_{3}&=\frac{4}{z}-\frac{2(z+2)}{z^{2}+4z+6},\\ y_{6}&=\frac{5}{z}-\frac{2(z+3)}{z^{2}+6z+12},\end{split}\qquad\begin{split}y_{1}&=-1+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{2}+4z+6}{z^{3}+3z^{2}+6z+6},\\ y_{4}&=-1+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{2}+6z+12}{z^{3}+6z^{2}+18z+24},\\ y_{7}&=-1+\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}\ln\frac{z^{2}+8z+20}{z^{3}+9z^{2}+36z+60},\end{split}\qquad\begin{split}y_{2}&=-1+\frac{3(z^{2}+4z+6)}{z^{3}+6z^{2}+18z+25},\\ y_{5}&=-1+\frac{3(z^{2}+6z+12)}{z^{3}+9z^{2}+36z+60},\\ y_{8}&=-1+\frac{3(z^{2}+8z+20)}{z^{3}+12z^{2}+60z+120}.\end{split}$
6 Discussion
In this paper we have derived discrete equations from the Bäcklund transformations of PV (1.2) and obtained hierarchies of rational solutions of the discrete equations. Whilst the asymmetric dPII (3.7), the discrete equation (3.17) and the ternary dPI (3.27) were derived by Tokihiro, Grammaticos and Ramani [69] from Bäcklund transformations of PV (1.2), our approach gives the explicit relationship between solutions of the discrete equations and solutions of PV. This enabled us to derive the hierarchies of rational solutions of the discrete equations.
Exact solutions of the discrete equations (3.7), (3.17) and (3.27) in terms of Kummer functions and modified Bessel functions (which are a special case of Kummer functions, cf. [54, §13.6(iii)]), using the explicit relationship between these equations and PV (1.2) given here, is currently under investigation, and we do not pursue this further here.
Acknowledgements
We thank Andy Hone, Edoardo Peroni and Alexander Stokes for stimulating discussions. BM is grateful to EPSRC for the support of Doctoral Training Partnerships grant EP/W524050/1.
ORCID iDs
Peter A. Clarkson 0000-0002-8777-5284
Clare Dunning 0000-0003-0535-9891
Appendix A Properties of the generalised Laguerre polynomials and the generalised Umemura polynomials
For convenience, in this appendix we list various differential-difference equations and discrete equations satisfied by the generalised Laguerre polynomials $T_{m,n}^{(\mu)}(z)$ and the generalised Umemura polynomials $U_{m,n}^{(\kappa)}(z)$.
A.1 Properties of the generalised Laguerre polynomials
Clarkson and Dunning [12, §3] prove the following differential-difference equations for the generalised Laguerre polynomials
${\rm D}_{z}\big(T_{m,n-1}^{(\mu+1)}{\,\scriptscriptstyle\bullet\,}T_{m,n}^{(\mu)}\big)=T_{m+1,n-1}^{(\mu)}\,T_{m-1,n}^{(\mu+1)},$ (A.1a)
${\rm D}_{z}\big(T_{m,n-1}^{(\mu+1)}{\,\scriptscriptstyle\bullet\,}T_{m+1,n}^{(\mu-1)}\big)=T_{m+1,n-1}^{(\mu)}\,T_{m,n}^{(\mu)},$ (A.1b)
${\rm D}_{z}\big(T_{m,n-1}^{(\mu+1)}{\,\scriptscriptstyle\bullet\,}T_{m,n}^{(\mu-1)}\big)=T_{m+1,n-1}^{(\mu)}\,T_{m-1,n}^{(\mu)},$ (A.1c)
${\rm D}_{z}\big(T_{m+1,n}^{(\mu)}{\,\scriptscriptstyle\bullet\,}T_{m,n}^{(\mu+1)}\big)=T_{m+1,n-1}^{(\mu+1)}\,T_{m,n+1}^{(\mu)},$ (A.1d)
${\rm D}_{z}\big(T_{m,n}^{(\mu)}{\,\scriptscriptstyle\bullet\,}T_{m,n}^{(\mu+1)}\big)=T_{m+1,n-1}^{(\mu+1)}\,T_{m-1,n+1}^{(\mu)},$ (A.1e)
${\rm D}_{z}\big(T_{m+1,n}^{(\mu)}{\,\scriptscriptstyle\bullet\,}T_{m,n}^{(\mu)}\big)=T_{m+1,n-1}^{(\mu+1)}\,T_{m,n+1}^{(\mu-1)},$ (A.1f)
where ${\rm D}_{z}(f{\,\scriptscriptstyle\bullet\,}g)$ is the Hirota operator (4.15), and the discrete equations
$T_{m,n+1}^{(\mu-1)}\,T_{m,n-1}^{(\mu+1)}$ $=T_{m+1,n}^{(\mu-1)}T_{m-1,n}^{(\mu+1)}-\big(T_{m,n}^{(\mu)}\big)^{\!2},$ (A.2a)
${T_{m+1,n}^{(\mu)}\,T_{m,n}^{(\mu)}}$ $={T_{m+1,n}^{(\mu-1)}\,T_{m,n}^{(\mu+1)}}-T_{m,n+1}^{(\mu-1)}\,T_{m+1,n-1}^{(\mu+1)},$ (A.2b)
$T_{m,n}^{(\mu-1)}\,T_{m,n-1}^{(\mu+1)}$ $=T_{m,n}^{(\mu)}\,T_{m,n-1}^{(\mu)}-T_{m-1,n}^{(\mu)}\,T_{m+1,n-1}^{(\mu)}.$ (A.2c)
The following differential-difference equations arise from the definitions of $w_{m,n}^{(\mu)}$, $v_{m,n}^{(\mu)}$ and $u_{m,n}^{(\mu)}$ in Theorem 4.4 and Remark 4.5
$z{\rm D}_{z}\big(T_{m-1,n}^{(\mu)}{\,\scriptscriptstyle\bullet\,}T_{m+1,n-1}^{(\mu)}\big)$ $=(m+1)T_{m,n-1}^{(\mu)}\,T_{m,n}^{(\mu)}+(z-m-2n-\mu)T_{m-1,n}^{(\mu)}T_{m+1,n-1}^{(\mu)},$ (A.3a)
$z{\rm D}_{z}\big(T_{m-1,n-1}^{(\mu+1)}{\,\scriptscriptstyle\bullet\,}T_{m,n}^{(\mu-1)}\big)$ $=(m+n+\mu)T_{m,n-1}^{(\mu+1)}\,T_{m-1,n}^{(\mu-1)}-(m+n)T_{m,n}^{(\mu-1)}\,T_{m-1,n-1}^{(\mu+1)},$ (A.3b)
$z{\rm D}_{z}\big(T_{m,n-1}^{(\mu+1)}{\,\scriptscriptstyle\bullet\,}T_{m-1,n+1}^{(\mu-1)}\big)$ $=nT_{m-1,n}^{(\mu+1)}\,T_{m,n}^{(\mu-1)}-(z-n-\mu)T_{m,n-1}^{(\mu+1)}\,T_{m-1,n+1}^{(\mu-1)},$ (A.3c)
and so from (A.1b) and (A.3b)
$zT_{m,n-1}^{(\mu)}\,T_{m-1,n}^{(\mu)}=(m+n+\mu)T_{m,n-1}^{(\mu+1)}\,T_{m-1,n}^{(\mu-1)}-(m+n)T_{m,n}^{(\mu-1)}\,T_{m-1,n-1}^{(\mu+1)}.$ (A.4)
The generalised Laguerre polynomials $T_{m,n}^{(\mu)}(z)$ satisfy the differential-difference equations
$z{\rm D}_{z}^{2}\Big(T_{m,n}^{(\mu)}{\,\scriptscriptstyle\bullet\,}T_{m,n}^{(\mu)}\Big)+2T_{m,n}^{(\mu)}\frac{\mathrm{d}{T_{m,n}^{(\mu)}}}{{\mathrm{d}{z}}}$ $=2(\mu+m+n+1)T_{m+1,n-1}^{(\mu+1)}T_{m-1,n+1}^{(\mu-1)},$ (A.5a)
$z{\rm D}_{z}^{2}\Big(T_{m,n}^{(\mu)}{\,\scriptscriptstyle\bullet\,}T_{m,n}^{(\mu)}\Big)+2T_{m,n}^{(\mu)}\frac{\mathrm{d}{T_{m,n}^{(\mu)}}}{{\mathrm{d}{z}}}$ $=2(\mu+m+n+1)\left\{T_{m,n}^{(\mu+1)}T_{m,n}^{(\mu-1)}-\left(T_{m,n}^{(\mu)}\right)^{\!2}\right\},$ (A.5b)
where
${\rm D}_{z}^{2}\big(f{\,\scriptscriptstyle\bullet\,}g\big)=g\frac{\mathrm{d}^{2}{f}}{{\mathrm{d}{z}}^{2}}-2\frac{\mathrm{d}{f}}{{\mathrm{d}{z}}}\frac{\mathrm{d}{g}}{{\mathrm{d}{z}}}+f\frac{\mathrm{d}{g}}{{\mathrm{d}{z}}},$
from which we obtain the identity
$T_{m+1,n-1}^{(\mu+1)}T_{m-1,n+1}^{(\mu-1)}=T_{m,n}^{(\mu+1)}T_{m,n}^{(\mu-1)}-\left(T_{m,n}^{(\mu)}\right)^{\!2}.$ (A.6)
A.2 Properties of generalised Umemura polynomials
Masuda, Ohta and Kajiwara [45], see also [44], prove the following discrete equations for the generalised Umemura polynomials
$U_{m+1,n}^{(\kappa-1)}\,U_{m-1,n+1}^{(\kappa+1)}$ $=U_{m+1,n+1}^{(\kappa-1)}\,U_{m-1,n}^{(\kappa+1)}+2(2m+1)\,U_{m,n+1}^{(\kappa-1)}\,U_{m,n}^{(\kappa+1)},$ (A.7a)
$U_{m,n+1}^{(\kappa-1)}\,U_{m+1,n-1}^{(\kappa+1)}$ $=U_{m+1,n+1}^{(\kappa-1)}\,U_{m,n-1}^{(\kappa+1)}-2(2n+1)\,U_{m+1,n}^{(\kappa-1)}\,U_{m,n}^{(\kappa+1)},$ (A.7b)
$U_{m-1,n}^{(\kappa)}\,U_{m+1,n-1}^{(\kappa)}$ $=z\,U_{m,n}^{(\kappa-2)}\,U_{m,n-1}^{(\kappa+2)}-2(2m+\kappa)\,U_{m,n}^{(\kappa)}\,U_{m,n-1}^{(\kappa)},$ (A.7c)
$U_{m,n-1}^{(\kappa)}\,U_{m-1,n+1}^{(\kappa)}$ $=z\,U_{m,n}^{(\kappa-2)}\,U_{m-1,n}^{(\kappa+2)}+2(2n+\kappa)\,U_{m,n}^{(\kappa)}\,U_{m-1,n}^{(\kappa)},$ (A.7d)
$U_{m+1,n}^{(\kappa-2)}\,U_{m-1,n-1}^{(\kappa+2)}$ $=z\,U_{m,n}^{(\kappa-2)}\,U_{m,n-1}^{(\kappa+2)}-2(\kappa-1)\,U_{m,n}^{(\kappa)}\,U_{m,n-1}^{(\kappa)},$ (A.7e)
$U_{m,n+1}^{(\kappa-2)}\,U_{m-1,n-1}^{(\kappa+2)}$ $=z\,U_{m,n}^{(\kappa-2)}\,U_{m-1,n}^{(\kappa+2)}+2(\kappa-1)\,U_{m,n}^{(\kappa)}\,U_{m-1,n}^{(\kappa)},$ (A.7f)
$U_{m-1,n+1}^{(\kappa-1)}\,U_{m,n-1}^{(\kappa+1)}$ $=z\,U_{m-1,n}^{(\kappa-1)}\,U_{m,n}^{(\kappa+1)}+2(2m+\kappa-1)\,U_{m,n}^{(\kappa-1)}\,U_{m-1,n}^{(\kappa+1)},$ (A.7g)
$U_{m+1,n-1}^{(\kappa-1)}\,U_{m-1,n}^{(\kappa+1)}$ $=z\,U_{m,n-1}^{(\kappa-1)}\,U_{m,n}^{(\kappa+1)}-2(2n+\kappa-1)\,U_{m,n}^{(\kappa-1)}\,U_{m,n-1}^{(\kappa+1)},$ (A.7h)
$U_{m+1,n}^{(\kappa-1)}\,U_{m-1,n-1}^{(\kappa+1)}$ $=z\,U_{m,n-1}^{(\kappa-1)}\,U_{m,n}^{(\kappa+1)}-2(2m+2n+\kappa)\,U_{m,n}^{(\kappa-1)}\,U_{m,n-1}^{(\kappa+1)},$ (A.7i)
$U_{m,n+1}^{(\kappa-1)}\,U_{m-1,n-1}^{(\kappa+1)}$ $=z\,U_{m-1,n}^{(\kappa-1)}\,U_{m,n}^{(\kappa+1)}+2(2m+2n+\kappa)\,U_{m,n}^{(\kappa-1)}\,U_{m-1,n}^{(\kappa+1)},$ (A.7j)
where ${\rm D}_{z}(f{\,\scriptscriptstyle\bullet\,}g)$ is the Hirota operator (4.15), see equation (4.16) and Proposition 4.1 in [45]. From equations (A.12) and (A.13) in [44, Appendix A] one can derive the following differential-difference equations for the generalised Umemura polynomials
$2{\rm D}_{z}\big(U_{m,n-1}^{(\kappa)}{\,\scriptscriptstyle\bullet\,}U_{m-1,n}^{(\kappa)}\big)$ $=U_{m,n-1}^{(\kappa)}\,U_{m-1,n}^{(\kappa)}-U_{m,n}^{(\kappa)}\,U_{m-1,n-1}^{(\kappa)},$ (A.8a)
$2{\rm D}_{z}\big(U_{m,n-1}^{(\kappa)}{\,\scriptscriptstyle\bullet\,}U_{m-1,n}^{(\kappa)}\big)$ $=U_{m,n}^{(\kappa-2)}\,U_{m-1,n-1}^{(\kappa+2)}-U_{m,n-1}^{(\kappa)}\,U_{m-1,n}^{(\kappa)},$ (A.8b)
$2{\rm D}_{z}\big(U_{m,n}^{(\kappa-1)}{\,\scriptscriptstyle\bullet\,}U_{m-1,n-1}^{(\kappa+1)}\big)$ $=U_{m,n}^{(\kappa-1)}\,U_{m-1,n-1}^{(\kappa+1)}-U_{m,n-1}^{(\kappa+1)}\,U_{m-1,n}^{(\kappa-1)},$ (A.8c)
$2{\rm D}_{z}\big(U_{m,n}^{(\kappa-1)}{\,\scriptscriptstyle\bullet\,}U_{m-1,n-1}^{(\kappa+1)}\big)$ $=U_{m,n-1}^{(\kappa-1)}\,U_{m-1,n}^{(\kappa+1)}-U_{m,n}^{(\kappa-1)}\,U_{m-1,n-1}^{(\kappa+1)}.$ (A.8d)
By adding and subtracting these equations we obtain the differential-difference equations
$4{\rm D}_{z}\big(U_{m,n-1}^{(\kappa)}{\,\scriptscriptstyle\bullet\,}U_{m-1,n}^{(\kappa)}\big)$ $=U_{m,n}^{(\kappa-2)}\,U_{m-1,n-1}^{(\kappa+2)}-U_{m,n}^{(\kappa)}\,U_{m-1,n-1}^{(\kappa)},$ (A.9a)
$4{\rm D}_{z}\big(U_{m,n}^{(\kappa-1)}{\,\scriptscriptstyle\bullet\,}U_{m-1,n-1}^{(\kappa+1)}\big)$ $=U_{m,n-1}^{(\kappa-1)}\,U_{m-1,n}^{(\kappa+1)}-U_{m,n-1}^{(\kappa+1)}\,U_{m-1,n}^{(\kappa-1)},$ (A.9b)
and the discrete equations
$2U_{m,n-1}^{(\kappa)}\,U_{m-1,n}^{(\kappa)}$ $=U_{m,n}^{(\kappa)}\,U_{m-1,n-1}^{(\kappa)}+U_{m,n}^{(\kappa-2)}\,U_{m-1,n-1}^{(\kappa+2)},$ (A.10a)
$2U_{m,n}^{(\kappa-1)}\,U_{m-1,n-1}^{(\kappa+1)}$ $=U_{m,n-1}^{(\kappa+1)}\,U_{m-1,n}^{(\kappa-1)}+U_{m,n-1}^{(\kappa-1)}\,U_{m-1,n}^{(\kappa+1)}.$ (A.10b)
Further differential-difference equations are given by
${\rm D}_{z}\big(U_{m+1,n}^{(\kappa-1)}{\,\scriptscriptstyle\bullet\,}U_{m-1,n}^{(\kappa+1)}\big)$ $=(2m+1)U_{m,n}^{(\kappa-1)}\,U_{m,n}^{(\kappa+1)},$ (A.11a)
${\rm D}_{z}\big(U_{m,n+1}^{(\kappa-1)}{\,\scriptscriptstyle\bullet\,}U_{m,n-1}^{(\kappa+1)}\big)$ $=(2n+1)U_{m,n}^{(\kappa-1)}\,U_{m,n}^{(\kappa+1)}.$ (A.11b)
The following differential-difference equations arise from the definitions of $\widehat{w}_{m,n}^{(\kappa)}$, $\widehat{v}_{m,n}^{(\kappa)}$ and $\widehat{u}_{m,n}^{(\kappa)}$ in Theorem 4.9 and Remark 4.10
$z{\rm D}_{z}\big(U_{m-1,n}^{(2\kappa)}{\,\scriptscriptstyle\bullet\,}U_{m,n-1}^{(2\kappa+2)}\big)$ $=(n+\kappa)U_{m-1,n}^{(2\kappa)}U_{m,n-1}^{(2\kappa+2)}-(m+\kappa)U_{m,n-1}^{(2\kappa)}\,U_{m-1,n}^{(2\kappa+2)},$ (A.12a)
$z{\rm D}_{z}\big(U_{m,n-1}^{(2\kappa)}{\,\scriptscriptstyle\bullet\,}U_{m-1,n}^{(2\kappa+2)}\big)$ $=(m+\kappa)U_{m,n-1}^{(2\kappa)}U_{m-1,n}^{(2\kappa+2)}-(n+\kappa)U_{m-1,n}^{(2\kappa)}\,U_{m,n-1}^{(2\kappa+2)},$ (A.12b)
$z{\rm D}_{z}\big(U_{m-1,n-1}^{(2\kappa+1)}{\,\scriptscriptstyle\bullet\,}U_{m,n}^{(2\kappa+1)}\big)$ $=\kappa U_{m-1,n-1}^{(2\kappa+1)}\,U_{m,n}^{(2\kappa+1)}-(m+n+\kappa)U_{m-1,n-1}^{(2\kappa+3)}\,U_{m,n}^{(2\kappa-1)},$ (A.12c)
$2z{\rm D}_{z}\big(U_{m+1,n}^{(\kappa-1)}{\,\scriptscriptstyle\bullet\,}U_{m-1,n}^{(\kappa+1)}\big)$ $=(2m+1)\left\{U_{m-1,n}^{(\kappa+1)}\,U_{m+1,n}^{(\kappa-1)}+U_{m,n-1}^{(\kappa+1)}\,U_{m,n+1}^{(\kappa-1)}\right\},$ (A.12d)
$2z{\rm D}_{z}\big(U_{m,n+1}^{(\kappa-1)}{\,\scriptscriptstyle\bullet\,}U_{m,n-1}^{(\kappa+1)}\big)$ $=(2n+1)\left\{U_{m-1,n}^{(\kappa+1)}\,U_{m+1,n}^{(\kappa-1)}+U_{m,n-1}^{(\kappa+1)}\,U_{m,n+1}^{(\kappa-1)}\right\}.$ (A.12e)
Consequently from (A.11a) and (A.12d), or from (A.11b) and (A.12e), we obtain
$2zU_{m,n}^{(\kappa-1)}\,U_{m,n}^{(\kappa+1)}=U_{m-1,n}^{(\kappa+1)}\,U_{m+1,n}^{(\kappa-1)}+U_{m,n-1}^{(\kappa+1)}\,U_{m,n+1}^{(\kappa-1)}.$ (A.13)
The generalised Umemura polynomials $U_{m,n}^{(\kappa)}(z)$ satisfy the differential-difference equations
$U_{m+1,n}^{(\kappa)}\,U_{m-1,n}^{(\kappa)}$ $=4z{\rm D}_{z}^{2}\Big(U_{m,n}^{(\kappa)}{\,\scriptscriptstyle\bullet\,}U_{m,n}^{(\kappa)}\Big)+8U_{m,n}^{(\kappa)}\frac{\mathrm{d}{U_{m,n}^{(\kappa)}}}{{\mathrm{d}{z}}}+(z-6m-2n-2\kappa-2)\left(U_{m,n}^{(\kappa)}\right)^{\!2},$ (A.14a)
$U_{m,n+1}^{(\kappa)}\,U_{m,n-1}^{(\kappa)}$ $=4z{\rm D}_{z}^{2}\Big(U_{m,n}^{(\kappa)}{\,\scriptscriptstyle\bullet\,}U_{m,n}^{(\kappa)}\Big)+8U_{m,n}^{(\kappa)}\frac{\mathrm{d}{U_{m,n}^{(\kappa)}}}{{\mathrm{d}{z}}}+(z+2m+6n+2\kappa+2)\left(U_{m,n}^{(\kappa)}\right)^{\!2},$ (A.14b)
and
$U_{m+1,n}^{(\kappa-2)}\,U_{m-1,n}^{(\kappa+2)}$ $=4z{\rm D}_{z}^{2}\Big(U_{m,n}^{(\kappa)}{\,\scriptscriptstyle\bullet\,}U_{m,n}^{(\kappa)}\Big)+8U_{m,n}^{(\kappa)}\frac{\mathrm{d}{U_{m,n}^{(\kappa)}}}{{\mathrm{d}{z}}}+(z+2m-2n-2\kappa+2)\left(U_{m,n}^{(\kappa)}\right)^{\!2},$ (A.15a)
$U_{m,n+1}^{(\kappa-2)}\,U_{m,n-1}^{(\kappa+2)}$ $=4z{\rm D}_{z}^{2}\Big(U_{m,n}^{(\kappa)}{\,\scriptscriptstyle\bullet\,}U_{m,n}^{(\kappa)}\Big)+8U_{m,n}^{(\kappa)}\frac{\mathrm{d}{U_{m,n}^{(\kappa)}}}{{\mathrm{d}{z}}}+(z+2m-2n+2\kappa-2)\left(U_{m,n}^{(\kappa)}\right)^{\!2},$ (A.15b)
with $U_{-1,-1}^{((\kappa))}(z)=U_{-1,0}^{((\kappa))}(z)=U_{0,-1}^{((\kappa))}(z)=U_{0,0}^{((\kappa))}(z)=1$. Eliminating the derivative terms in (A.14) and (A.15) gives
$U_{m,n+1}^{(\kappa)}\,U_{m,n-1}^{(\kappa)}-U_{m+1,n}^{(\kappa)}\,U_{m-1,n}^{(\kappa)}$ $=4(2m+2n+\kappa+1)\left(U_{m,n}^{(\kappa)}\right)^{\!2},$ (A.16a)
$U_{m,n+1}^{(\kappa-2)}\,U_{m,n-1}^{(\kappa+2)}-U_{m+1,n}^{(\kappa-2)}\,U_{m-1,n}^{(\kappa+2)}$ $=4(\kappa-1)\left(U_{m,n}^{(\kappa)}\right)^{\!2}.$ (A.16b)
Also from (A.14) and (A.15) we obtain the difference equations
$U_{m,n+1}^{(\kappa-2)}\,U_{m,n-1}^{(\kappa+2)}-U_{m+1,n}^{(\kappa)}\,U_{m-1,n}^{(\kappa)}$ $=4(2m+\kappa)\left(U_{m,n}^{(\kappa)}\right)^{\!2},$ (A.17a)
$U_{m,n+1}^{(\kappa)}\,U_{m,n-1}^{(\kappa)}-U_{m+1,n}^{(\kappa-2)}\,U_{m-1,n}^{(\kappa+2)}$ $=4(2n+\kappa)\left(U_{m,n}^{(\kappa)}\right)^{\!2},$ (A.17b)
$U_{m+1,n}^{(\kappa-2)}\,U_{m-1,n}^{(\kappa+2)}-U_{m+1,n}^{(\kappa)}\,U_{m-1,n}^{(\kappa)}$ $=4(2m+1)\left(U_{m,n}^{(\kappa)}\right)^{\!2},$ (A.17c)
$U_{m,n+1}^{(\kappa)}\,U_{m,n-1}^{(\kappa)}-U_{m,n+1}^{(\kappa-2)}\,U_{m,n-1}^{(\kappa+2)}$ $=4(2n+1)\left(U_{m,n}^{(\kappa)}\right)^{\!2}.$ (A.17d)